### v3

# Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem

Abstract
Software signing is the strongest method for ensuring the integrity and authenticity of components in a software supply chain. Although software signing has been recommended by regulation and industry consortia, practical adoption of software signing is generally limited. Previous studies have recognized tooling as a key factor in influencing software signing adoption and quality. However, most research has primarily focused on its user interface aspects, with little attention given to other usability considerations, such as case-specific user challenges and factors influencing tool adoption and changes.
To understand how software tools influence the practice and adoption of software signing, we study the formative usability of Sigstore, a modern and widely adopted software signing tool. We interviewed thirteen (13) experienced security practitioners to study the factors that influence the selection of a tool, the problems associated with the use of such tools, how practitioners’ software signing tools have evolved, and what drives this migration. To summarize our findings: (1) We highlight the various factors practitioner’s consider before adopting a software signing tool; (2) We highlight the problems and advantages associated with the current tooling choices of practitioners; and (3) We describe the evolution of tooling adoption of our sample population. Our findings provide the software signing tool development community with valuable insights to improve their design of software signing tools.
1 Introduction
The reuse of software components in modern software products results in the formation of complex software supply chains, leading to an increased attack platform for these products [39, 79, 67]. A key security strategy to mitigate this expanding attack surface is ensuring the authenticity of actors and integrity of artifacts by establishing provenance [54, 68]. Software signing is the strongest guarantee for establishing provenance [69] and has been extensively advocated by academia, industry consortia, and government regulations [81, 16, 14]. To effectively implement this process, software signing tools–such as OpenPGP [55], Sigstore [53], and Notary [57]–have been proposed and are currently used in practice to sign software artifacts.
Although tooling has significantly advanced various technological fields [71, 41, 3, 47], its impact on software signing must be carefully examined to maximize its benefits. While prior "Why Johnny Can’t Encrypt" inspired studies have explored the usability of signing tools, their focus has been limited primarily to user interface design [89, 73, 63]. Furthermore, their evaluation techniques rely primarily on summative usability assessments (ease of use), which have been criticized for being difficult to measure and for providing limited insights to inform tool enhancements [13, 48]. The applicability of these studies is also constrained by their narrow focus on message (and email) signing and their reliance on novice volunteer participants, which limits the real-world relevance of their findings for software signing.
Our work provides the first examination of the usability of tooling as a key factor influencing the adoption and practice of software signing. We conducted interviews with 13 experienced security practitioners to explore current software signing tools, with a focus on Sigstore as a case study. We apply a formative usability framework  [17], which emphasizes the iterative refinement of the studied tooling option, to develop our study instruments. Our study explores usability in two categories: users’ experience with the specified tooling option and the usability of the tooling options with respect to other software signing tools. Our study thus explores three key themes: the evolution and motivations behind software signing tool adoption, limitations and challenges encountered with currently adopted tooling, and the specific implementations of adopted tools.
Our results highlight the factors practitioners consider when adopting specific software signing tools, using a case study of a key software signing tool. This serves as formative feedback for tool designers, offering insights into the drivers of tool usage. We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools. Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of modern Software signing tools.
In summary, we make the following contributions:
  1. 1.
    We report on the difficulties and advantages of using Sigstore, a next-generation software signing tool (section 5.1.1, section 5.1.2).
  2. 2.
    We discuss how practitioners’ software signing tools evolve over time and the factors that influence practitioners’ decisions to switch tooling options (section 5.2).
2 Background & Related Works
In this section, we discuss the foundations of software signing (section 2.1), usability studies in software signing (section 2.2), and the Sigstore ecosystem where we situate our work (section 2.3).
2.1 Software Signing
2.1.1 Signing Overview and Workflow
Software signing is a process that ensures the authorship and integrity of a software artifact by linking a cryptographic signature created by the software maintainer with their software artifact. Fundamentally, software signing is based on the public-key cryptography model  [69], and a traditional signing workflow follows a three-stage process of creation, distribution, and verification. We describe a typical signing workflow in Appendix C.
While the fundamental software signing process remains consistent across implementations, different signing tools and policies are shaped by the target platform, software artifacts, and root of trust. The choice of signing tools and the policies under which developers operate often depend on the operating system and distribution vendor. For example, Microsoft Authenticode [30] and Apple Xcode [33] provide platform-specific signing mechanisms, while package registries like PyPi [45] and Maven [1] enforce signing policies through dedicated tooling plugins. Developers also have a variety of software signing tools available to them, such as GNU Privacy Guard (GPG) [58], Signify [85], Sigstore [53], and Notary [57]. Each of these tools present different advantages and challenges, shaping how developers integrate signing into their workflows.
2.1.2 History and State-of-Art Tooling
Traditional Signing Tools: Many traditional signing tools, such as Git Commit Signing and GNU Privacy Guard are based on PGP. PGP (Pretty Good Privacy) is a proprietary encryption software originally developed by Phil Zimmerman in 1991. It is a public key cryptography-based protocol whose capabilities include encrypting files, creating and managing public and private keys, and creating digital signatures. Its open-source standard is OpenPGP [23, 55]. However, two main struggles associated with PGP-based software signing involve key management and establishing trust in public keys.
PGP relies on a "web of trust" model, where a signer’s public key is considered trustworthy based on endorsements from other trusted signers [23]. Each created public key is stored in a "key certificate" that specifies the public key, the name and/or email address associated with the public key, its creation date, and others who have signed the key. The signer is also responsible for secure management of their private key, preventing exposure, as the combination of their private key and passphrase enables the signing of artifacts [45].
State-of-art Signing Tools: Although software signing is widely recognized as an effective technique for mitigating software supply chain attacks, several obstacles hinder its adoption. Research highlights the prevalence of unsigned artifacts in software registries, emphasizing the need for improved signing practices [53, 90, 46]. This has led to the emergence of the next generation of software signing. Newman et al. introduced a new software signing platform in 2021 called Sigstore that addresses key adoption barriers, particularly key management, by leveraging keyless/ephemeral software signing techniques [53].
Keyless signing mitigates security risks associated with private key compromise. Instead of relying on a long-term private key, the signer uses a short-lived key that exists only for the duration of the signing process. This approach enhances security by preventing private key leaks while improving accessibility and usability [2]. As a result, many software registries are transitioning to this model to streamline verification and enforce stronger software integrity measures [15, 34]. While Sigstore’s development aims to solve the usability issues associated with traditional signing methodology, firsthand perspectives from practitioners using Sigstore will provide valuable insight into its actual usability and convenience.
2.2 Usability Studies in Software Signing
2.2.1 Usable Tools: Definition and Theory
Usable Tools: In basic terms, a software engineering tool assists developers in their software production process [27]. For software signing specifically, these tools are used to provide a method for which users can verify the integrity and authenticity of the code to be published, and developers use these tools to verify signatures of the software artifacts they may incorporate into their own projects. Deciding which software engineering tools to use is a critical question developers face when many options are available. A specific tool can affect a developer’s workflow efficiency, productivity [44], and the resulting quality of the work that is produced. This allows members of the computing industry to remain competitive by selecting the most suitable tools and environments that are conducive for the efficient production of software [27]. Additionally, tooling is important because it enables engineers to use software development practices they have not in the past [11].
Usability Evaluation: Studies assessing implementation strategies of products are typically framed as usability evaluations [49]. Usability evaluations are generally categorized into two forms: summative (validation or ease of use) and formative (diagnostic usability) [48]. Summative usability refers to the extent to which a system or product can be used to achieve a specified goal effectively, efficiently, and satisfactorily within a given context [35, 65]. In contrast, formative usability focuses on product assessments aimed at identifying areas for improvement [82]. Although summative usability evaluations are more commonly used, they are not designed to provide actionable feedback for enhancing the product [48]. Summative evaluations primarily measure metrics such as effectiveness, efficiency, and user satisfaction [35]. However, such metrics have been widely criticized for being subjective and difficult to quantify [48, 13].
In contrast, formative usability evaluations aim to diagnose user challenges and provide iterative feedback to enhance product functionality [82, 48]. These evaluations emphasize real-world usability issues and are more suited for refining a product’s design. While summative evaluations are more common, their limitations highlight the need for formative assessments when improving software signing tools.
2.2.2 Empirical Studies
Signing Tooling Usability Studies: The usability of digital signing tools has been widely studied, though often within the context of encryption rather than broader software signing applications. Whitten and Tygar’s seminal study, Why Johnny Can’t Encrypt [89], found that PGP 5.0’s user interface was too complex for non-expert users, making encryption and signing inaccessible. Sheng et al. [73] later observed that PGP 9’s interface had not significantly improved, particularly in key verification, encryption transparency, and digital signing usability. Other studies [10, 64, 60] have highlighted persistent usability issues, including the complexity of the public-key model.
Further research has explored the differences between manual and automated signing tools. Fahl et al. [19] and Ruoti et al. [62] found no significant usability differences between these approaches. Meanwhile, Atwater et al. [5] demonstrated that users preferred automated solutions based on the PGP standard. While these studies provide insights into encryption tools, their relevance to broader software signing workflows is limited.
These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing. First, they primarily focus on evaluating the usability of signing tooling in the context of email or message encryption, limiting their application to software signing. Additionally, they limit their scope to user interface design, overlooking critical factors such as tooling adaptability and macroenvironment (e.g., tooling user communities). Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences. Their methodology predominantly employs summative usability assessments, which focus on ease of use [13]. These studies typically rely on novice volunteers to evaluate usability. In addition to the flaws of the summative evaluation approaches, the reliance on novice participants limits the applicability of their findings. Evaluating usability through inexperienced users does not yield specialized feedback that can effectively guide tool enhancements [42, 50].
2.2.3 Related Software Signing Research
The use of software signing as a security method to mitigate Software Supply Chain security has been studied in the literature. These studies has focussed on the adoption and practice of Software signing. Schorlemmer et al. [69] examine factors influencing software signing adoption through a framework of signing events. They identify tooling as a major factor affecting the quality of software signatures but focus primarily on the role of dedicated tooling within software registries, rather than its broader impact on engineering workflows. In our parent study, Kalu et al. [37], we provided valuable insights into the implementation and adoption frameworks of software signing, with a focus on broad software signing challenges faced in commercial settings. These challenges included issues of usability, compatibility, and integration. However, while we highlighted these broad challenges, our findings lacked the specificity needed to identify which tools are most impacted by these issues and under what conditions of use, limiting practitioners’ contextual understanding of these problems.
These prior studies acknowledge that improved tooling can drive the adoption of software signing. However, they fail to clearly define what constitutes an "improved tool" from the perspective of practitioners. While they emphasize broad usability factors, they overlook the specific technical and workflow-related requirements that practitioners consider when selecting tools for real-world use.
2.3 Sigstore
Due to the emergence of next-generation software signing and the growing adoption of Sigstore as a signing platform, we focus on the Sigstore tool specifically in our research.
Importance: Sigstore, launched in 2021, was developed to address critical challenges in software signing, including complex key management, usability barriers, and software supply chain security risks. Traditional signing tools, such as OpenPGP and enterprise certificate authorities, rely on long-term cryptographic key storage, which introduces risks of key compromise, revocation difficulties, and usability friction. In contrast, Sigstore simplifies the signing process by providing keyless, ephemeral authentication mechanisms. Within 13 months of its public release, Sigstore recorded 46 million artifact signatures  [29]. It has since gained significant adoption and support from open-source ecosystems  [84, 25, 31, 6, 86] and various organizations  [38, 76, 75, 77]. Given its growing adoption and integration in major software ecosystems, Sigstore represents a state-of-the-art software signing approach, making it a compelling case study for examining usability and security trade-offs.
Design Aspects Affecting Usability: One key challenge that is associated with traditional signing methods is key management, which is addressed by Sigstore. In conventional certificate authority-based signing, such as Windows Authenticode, a software artifact’s validity is linked to the certificate’s status—whether it is still valid, expired, or revoked [43, 70]. A revoked certificate can indicate compromise, but it may not be immediately clear when the compromise occurred [40]. This uncertainty allows some revoked certificates to be used for signing malware while still appearing valid to verification systems. Sigstore mitigates this risk by issuing short-lived certificates. These certificates significantly reduce the chances of key and certificate compromise since they automatically expire after a short duration, limiting their potential misuse.
Complexity: Sigstore uses a modular design of several components to implement software signing. As described by [53], Sigstore consists of eight key components: signers, verifiers, the artifact log, an OpenID Connect (OIDC) provider, a certificate authority, monitors, transparency logs, and package repositories. While Sigstore is designed to function as a cohesive system, these components can also be implemented individually based on specific use cases. The signing process starts with the signer using OpenID Connect (OIDC) to obtain a signing certificate. Unlike traditional methods requiring long-term key management, Sigstore associates a signer’s identity with short-lived tokens issued by services such as Google, GitHub, and Microsoft. Next, a software signing certificate is issued to the signer based on their verified identity. In this step, Fulcio serves as the certificate authority (CA), ensuring that the OIDC token was correctly issued to the requesting signer before granting the certificate. To monitor signing behavior, a verifier makes use of Fulcio’s log and the transparency log (Rekor) to verify for identity of the creator of the signatures, the certificate validity, and signature validity.
3 Knowledge Gaps & Research Questions
The existing literature on software signing, as outlined in  section 2, has examined its practice and adoption across different contexts, including private and open-source ecosystems. These studies highlight the critical role that the choice of a software signing tool plays in ensuring the quality of the software signing process.
Although certain aspects of software signing tools have been explored, such as the usability of the user interface design and functional exclusivity (dedication of tooling resources), the broader impact and role of software signing tools remain largely understudied. The findings of these works underscore that tooling is a multi-faceted and multi-layered factor. A general usability evaluation of these tooling options would significantly enhance their adaptability and facilitate improvements. Also, the existing research on software signing tooling primarily relies on experiments with novice users and quasi-experiments to investigate isolated aspects of tooling. Practitioner perspectives on the nuanced role of tooling in shaping the use and practice of software signing remain limited, particularly regarding how the usability of these tooling contributes to their adoption.
Usability Framework:
Figure 1: Usability Framework from Cresswell et al. [17]. Their framework assesses usability based on four factors: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors. Our framework applies their framework to two concerns of software signing tool usage: (1) the tool itself and (2) previous signing experiences.
To examine the usability of software signing tooling, an examination of experienced practitioner experiences is necessary. As noted in section 2.2, the summative usability approach mostly is mostly hindered by its use of subjective measurement metrics. To effectively, answer usability in terms of how software signing tooling are used in practice, a formative usability framework is warranted. A formative framework addresses the experiences of regular users of a tool, rather than controlled experience of experimental subjects  [48].
Our formative usability evaluation framework, Figure 1, is based on Cresswell et al.’s [17] formative framework. Their framework identifies four key factors that influence the overall usability of a product or technology: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors. Usability is determined by the interplay of these factors.
Our formative evaluation applies of their four-factor framework to two key areas: users’ experiences with the specified tooling option and the usability of the tooling option compared to previous software signing tools. We designed our study instrument to address these broad usability themes, emphasizing the identified usability factors.
Research Questions: Our research addresses the following questions in light of the positions outlined above within the context of Sigstore:
  • RQ1:
    What drawbacks and benefits of software signing tools influence their adoption in practice?
  • RQ2:
    What factors influence practitioners’ decisions to switch software signing tools over time?
4 Methodology
Figure 2: Study Methodology. We developed our interview protocol by reviewing Sigstore grey literature and the academic literature using our adopted formative usability framework. Finally, we analyzed our results using a thematic analysis informed by our usability evaluation framework (Figure 1).
We undertook a qualitative case study of the Sigstore software signing ecosystem through interviews of engineers. We applied a formative usability framework to identify the key factors that impact the adoption of Sigstore in practice.
Our methodology is illustrated in Figure 2. This study was approved by our Institutional Review Board (IRB).
4.1 Research Design and Rationale
4.1.1 Rationale
The case study is a research method used to investigate a phenomenon within its real-life context  [61, 91, 18, 24]. We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice, rather than a broad but superficial cross-analysis  [59]. In selecting Sigstore, we adhered to the criteria outlined by Yin [91] and Crowe et al. [18]. We considered the intrinsic uniqueness of the unit (high), access to data (adequate), and risks of participation (low).
Our research objectives explore the impact of Software Tooling as a factor in the practice of software signing from the viewpoint of practitioners, with a particular focus on how the usability of software signing tools influences their adoption. We opted for semi-structured interviews [66] to effectively examine various aspects of tooling, including usability challenges, perceived benefits, and comparative experiences with other signing tools. This approach allows us to gather descriptive, open-ended, opinionated, and detailed responses, providing deeper insights into how usability considerations shape adoption decisions.
4.1.2 Instrument Design & Development
To develop our interview instrument, we relied on the usability framework (Figure 1) and various signing literature. We outline how these resources informed our instrument design:
  1. 1.
    Grey Literature: To derive our questions, we primarily examined Sigstore blogs [74], and Sigstore usage reports [8, 86, 31, 38, 76, 75, 77, 15]. These sources included reports on how various entities implemented Sigstore in their engineering processes, news about major updates, and other related Sigstore developments.
  2. 2.
    Background & Literature Review (section 2): We employed a snowball approach to examine works from top security conferences on software signing, the design of software signing primitives [53, 52], and empirical studies on software signing (all cited in (section 2)). Our literature review informed us of existing knowledge gaps and provided insights into the design of modern software signing tools. This helped us avoid duplicating questions already addressed in the literature.
  3. 3.
    We designed our questions to align with our formative usability framework (section 3). We developed these questions to focus on identifying the factors responsible for the adoption of Sigstore in practice. Categories D3 and D4 focused on participants’ experiences with their current tools, aligning with the framework’s emphasis on assessing present tooling experiences. Meanwhile, categories D1, D2, and D5 captured participants’ experiences with previous software signing tools, offering insights into their transitions and comparisons over time.
Next, we refined our interview protocols through a series of practice and pilot interviews. We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
Our full interview instrument consisted of three main topics (B-D). As suggested by the table, topics B and C do not fit into the same analysis framework as topic D. Thus, following the guidance of Voils et al. [88], we analyzed those topics in another work [37]. In this work, we focus on the data for Topic D. Topic D consisted of five main question categories, conducted in a semi-structured manner. A summary of the questions in our interview protocol is shown in Table 1.
Table 1: Summary of interview protocol. Topic B and C are analyzed in [37]. We analyze only topic D in this paper. The decision to analyze these parts of the protocol separately was made in accordance with best practices. We present our full protocol in  Appendix A
Topic(# Questions) Sample Questions
A. Demographics (4) What is your role in your team?
B. Perceived software supply chain risks (4) Can you describe any specific software supply chain attacks (e.g., , incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
C. Mitigating risks with signing (8) How does the team use software signing to protect its source code (what parts of the process is signing required)?
D. Signing tool adoption (5)
Question Categories in Topic D
D1. What factors did the team consider before adopting this tool/method (Sigstore) over others?
D2. What was the team’s previous signing practice/tool before the introduction of Sigstore?
D3. How does your team implement Sigstore (which components of Sigstore does the team mostly use)?
D4. Have you encountered any challenge(s) using this tool of choice?
D5. Have you/your team considered switching Sigstore for another tool?
4.2 Data Collection
Recruitment: Our study investigates signing tools using a case study of the Sigstore ecosystem. As such, our target participant pool is centered on Sigstore users. Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is implemented and the challenges and benefits of Sigstore.
To effectively answer these questions related to decision-making, we required a target population of high-ranking or experienced practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [7]. We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [83, 32] Steering Committee (ITSC). This initial invitation yielded six (6) participants. We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (five(5)) and authors contacts (seven) to recruit an additional twelve(12) participants. We interviewed 18 subjects, but only responses from 13 participants fit our analysis framework [^1] We interviewed 18 participants but ultimately excluded data from five for this study. Two of them used software signing tools other than Sigstore, and three gave limited information about their software signing tool..
Participant Demographics:
Table 2: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We highlight type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products to the one stated. We also highlight which participants were recruited from the In-toto Steering committee (I) and which were from snowball sampling (S).
ID Role Experience Software Type Rec.
P1 Research leader 5 years Internal POC software I
P2 Senior mgmt. 15 years SAAS security tool*O S
P3 Senior mgmt. 13 years SAAS security tool*O I
P4 Technical leader 20 years Open-source tooling S
P5 Engineer 2 years Internal security tooling S
P6 Technical leader 27 years Internal security tooling*O S
P7 Manager 6 years Security tooling*O S
P8 Engineer 2.5 years SAAS security* I
P9 Technical leader 4 years Consultancy S
P10 Research leader 13 years POC security Software* S
P11 Senior mgmt. 15 years Internal security tooling I
P12 Senior mgmt. 15 years SAAS S
P13 Manager 11 years Security tooling S
We summarize our interview participants’ demographic information in Table 2. Our interview sampling approach yielded 13 experienced security practitioners who indicated that they are responsible for initiating or implementing their organization’s security controls and (or) have compliance-related responsibilities. This involvement gives them the relevant context to assess their organization’s strategies for adopting software signing tools.   Interviews: We conducted our interviews over Zoom. Each interview lasted a median of 50 minutes. The lead author conducted these interviews. We offered each subject a \$100 gift card as an incentive, in recognition of their experience level.
4.3 Data Analysis
We transcribed our interview recordings using www.rev.com’s human transcription service and anonymized all potentially identifiable information.
To analyze our interviews, we employed a thematic analysis approach [9] alongside our modified usability framework (section 3, Figure 1) within an interpretivist paradigm [4].
An interpretivist approach allows each participant’s response to be treated as a unique case, considering its specific context and surrounding variables. This is particularly valuable in our study, as participants belong to different teams with diverse implementation contexts and varying uses of Sigstore.
Our analysis began with a data familiarization stage [80] and proceeded through three phases, carried out by two analysts  [^2] One of the analysts is the lead author.. In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [21]. In the final phase, we derived themes from our inductively generated codes and then applied our usability framework to further refine the extracted themes. We describe these steps in detail next.
Memoing & Codebook Creation: First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [56] and Campbell et al. [12]. This approach ensured the reliability of our analysis process while optimizing data resources. [^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [56]. We randomly selected approximately 33% (six transcripts) to develop our codebook. This slightly exceeded O’Connor & Joffe’s recommended percentage to allow the analysts a better familiarization with the nature of our collected data, and analysis format.
After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos.
We then reviewed the coded memos and finalized an initial codebook with 36 code categories. To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [51] and Campbell et al. [12]. Specifically, we randomly selected an unanalyzed transcript, which we coded independently. Following Feng et al. ’s [20] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook. We obtained an 89% agreement score and resolved all disagreements through discussions and code refinements. This high agreement indicates the stability of our coding scheme, allowing us to reliably singly code the remaining transcripts [12, 56].
Codebook Revision: Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round. In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst. The same analytical process was applied to the final six transcripts, leading to the addition of four new code categories, which were subsequently reviewed by the other analyst. Notably, of the 10 codes added during this phase of codebook refinement, only three provided new insights into the data. By mutual agreement between both analysts, these were categorized as minor codes.
To ensure our sample size and analysis was adequate, we calculated code saturation for each category in our codebook. Following the recommendations and methods of Guest et al. [26] and Maxam & Davis [51], we measured the cumulative number of new codes introduced in each interview. The results are presented in Figure 6. As shown, saturation was achieved after interview 11 , with no new codes emerging beyond this point. Each interview had a median of 15 unique codes (for topic D alone), as illustrated in Figure 6.
Thematic Analysis & Usability Framework Analysis: Next, following the method proposed by Braun & Clarke [9], we derived initial themes from our code categories. We focused on identifying themes related to the D part of our protocol, as highlighted earlier in Table 1. We then conducted another round of thematic analysis, incorporating the initial themes alongside our formative usability framework (section 3). In this phase, we mapped our initial themes onto the factors defined within our framework.
Survey Attempt: In addition to our interviews, we designed and conducted a survey to gather insights from members of the Sigstore community who might not have had the time for an interview. However, our survey received only a few responses, highlighting the challenges of conducting surveys within highly specialized communities [78]. We elaborate on our survey results and details in Appendix D.
4.4 Study Limitation & Threats to Validity
To discuss the limitations of our work, we follow Verdecchia et al.’s [87] recommendation to reflect on the greatest threats to our research. In addition, we discuss how our experiences and perspectives may have influenced our research.
Construct Validity: Cresswell’s usability framework was developed for health information technology. We adapted it to software signing tooling, but may have introduced flaws.
Internal Validity: Threats to internal validity reduce the reliability of conclusions of cause-and-effect relationships. An additional internal threat exists in the potential subjectivity of the development of our codebook. To mitigate this subjectivity, we utilized multiple raters and measured their agreement.
External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore. We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature. A secondary threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects. A final threat to the validity of our research is that it reflects the state of Sigstore at only one point in time. Our interviews were conducted between November 1, 2023, and February 9, 2024 and reflect the experiences of Sigstore users at that time.
Statement of Positionality: We acknowledge that one of our paper’s authors is a Sigstore tool contributor. This author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
5 Results
5.1 RQ1: Problems & Strengths of Sigstore
Table 3: Summary of Practitioner-Reported Advantages of Using Sigstore
Topics & Associated Examples Subjects
Technological Factors
Ease of Use 8
Signing Workflow & Verification P1, P2, P7, P8, P10-13
Setting up with automated CI/CD actions P9
No key distribution problems P1
Use of Short-lived Keys & Certificate 3 –P2, P3, P11
Signer ID Management 4
Use of OIDC(Keyless) to authenticate signers P3, P5, P10, P12
Compatibility with Several New Technologies 4
Integrability with SLSA build P12
Integrability with several container registries/technologies P9, P10
Integrability with several cloud-native applications P11
Precence of a Transparency Log 3
Transparency logs increase security P5, P10
Evaluation of signing adoption using logs P8
Bundling Signatures With Provenance Attestations 2 – P3, P4
Reliability of Service 1 – P7
Macroenvironmental Factors
Free/Open-Source 2 – P7, P13
The first aim of our research work was to ascertain from the the experiences of our subjects the problem and strengths they have experienced while using Sigstore. We categorized each identified theme based on our usability framework factors: Technological Factors – situations where observed conditions were linked to the behavior of Sigstore technology itself, Human/Social Factors – participants’ observations related to gaps in support resources, previous experiences (with other tools) Organizational Factors – aspects tied to users’ organizations or their organizational capacities, and Macroenvironmental Factors – external influences affecting adoption and use.
5.1.1 Sigstore’s Strength
We present a summarized list of practitioner-reported advantages of using Sigstore in Table 3. First, we extracted the benefits our participants experienced while using Sigstore from their responses. These reflections capture their firsthand experiences with Sigstore.
We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service. When mapped to our modified usability framework, these factors fell under two categories: Technological and Macroenvironmental factors from our framework.
Technological Factors
Ease of use: Ease of Use was a predominant reason why many participants chose to use Sigstore, making it the most frequently cited advantage over other Software signing tools. Participants commonly highlighted the simplicity of following the Sigstore workflow for signing and verifying signatures. P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
Identity management for Signers: Sigstore’s integration of OIDC identity management was a key strength reported by our participants. In traditional software signing, signer identity is often an ad-hoc feature that may or may not be required by the signing authority. In Sigstore, however, this feature—implemented via OIDC—is integrated with the Fulcio certificate authority as an identity management step before creating a signature. Participants greatly valued the elimination of the key generation and distribution phase, as it reduced the time spent distributing keys to key servers and searching for keys, streamlining the signing process. P12 (SAAS, SSC security tool) Highlights this point, “The other strength that it has is I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key … that’s super useful because I could say, "Oh, this is signed with Mike’s public GitHub identity." So unless my GitHub identity has been compromised, that’s much better.”
Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage. The mentioned technologies included security tooling such as SLSA build attestations, as well as containerization technologies and cloud computing applications. In most cases, the resources required to set up connections to these technologies were minimal or nonexistent. P9 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
Use of Short-lived Keys/Certificates: Sigstore’s enforcement of short-lived keys and certificates as a security feature was recognized by our participants as a major strength of the tooling. Traditional software signing tools require users to manually set expiration times for cryptographic elements such as signatures, keys, and certificates at the time of generation. While this approach provides customization flexibility, it also introduces the risk of long-lived cryptographic materials remaining valid for extended periods, increasing the likelihood of compromise. Participants appreciated Sigstore’s automated management of key and certificate lifetimes, which mitigates this risk. P2 (SAAS, SSC security tool) states, “So (with) PGP, you still have to figure out the key distribution problem. Because Sigstore uses short-lived keys that solves a lot of those issues around key storage and key distribution…”
Presence of a Transparency Log: The inclusion of a transparency log in the Sigstore ecosystem, serving as a tamper-resistant secondary record of changes to an artifact’s metadata, was seen as highly beneficial by participants. This feature provided an additional layer of trust, allowing users to verify signed artifacts with greater confidence. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
Other technological factors noted by participants are the reliability of Sigstore’s public instance was a valuable aspect of Sigstore.
Macroenvironmental factors
Free/Open-Source:
Sigstore’s open-source nature was also highlighted by participants as a key advantage. They noted that the no-cost solution reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
5.1.2 Sigstore’s Weaknessses
We summarize the difficulties reported by participants in using Sigstore in  Table 4. Most of the weaknesses reported often belong to multi-factors. e.g., In the problems associated with Enterprise applicabiliirty of Sigstore, Some participants reported this as a consequence of regulations place on organizations which are both macroenvironmental (regulations) and Organizational.
Table 4: Reported Difficulties Encountered by Participants in Using Sigstore. We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.
Topics & Associated Examples Subjects
Enterprise Adoption Limitations 7
Rate Limiting Problems – T P3, P7, P10, P11
Lack of dedicated Support & Maintenance – T & P P11, P2
Not Suited for Regulated Organizations – M & O P3
Latency Concerns – T P6
Transparency Log Issues 6
Not Suitable for private Setup – T P2, P3, P6, P10, P13
Use in Air Gap Conditions – T P2, P3, P8
Efforts to Monitor Logs – P P2
Private Sigstore Instance Setup 5
Documentation – P P8, P9, P6
Limited Community Support – M P11
Infrastructure Requirements & Maintenance Costs – T P5, P6
Other Documentations and Usage Information IssuesP 3 — P1, P10, P13
Integration to Other Systems 3
Attestation Storage – T P1
Gitlab & Jenkins – T P8
Other Unsupported technologies – T P12
Offline CapabilitiesT 2 – P3, P4
Fulcio Issues 1 – P3
Timestamping Issues – T
Fulcio-OIDC Workflow – T
Software LibrariesT 1 – P7
Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications. The issues reported by participants include rate limiting—where Sigstore restricts the number of signatures that can be created per unit time—along with latency concerns, where the service’s turnaround time for a large volume of signatures is problematic. Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore ’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “I think the major challenge was that we were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise. We would constantly be getting rate limited by the service.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable. And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
Transparency Log Issues: While the transparency log bundled with Sigstore was highlighted as a major advantage by participants, it was also a significant concern. Their concerns stemmed from the public nature of the log, where sensitive artifact metadata from company assets could be exposed in a publicly accessible record. Another common issue raised was the log’s usability in air-gapped environments—systems that are physically isolated from unsecured networks for security reasons. Additionally, participants noted the manual effort required to continuously monitor and manage the log. Most of these participants were in the process of experimenting with setting up the transparency log at the time of the interviews. Transparency log issues were mostly technological.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P6 (security tooling, digital technology): “We’ve got some teams piloting Rekor, for instance, for different traceability usages and so on…I’m all for the transparency log, but we just have to be careful who it’s transparent to at what point in time.”
Setting up Private Sigstore Instance : Sigstore’s customizability was a feature frequently utilized and discussed by participants; however, it was also associated with certain challenges. The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments. Issues here were primarily related to participants’ need for support resources, which where macroenvironmental and social in nature. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P11 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to help people deploy Fulcio, like on prem, or in their own infrastructure.”
Other Documentation Issues: Overall, the documentation for Sigstore, along with practical usage information and support, was found to be insufficient. Participants noted that updates to the documentation lagged behind changes in the Sigstore product and that there was a lack of clarity regarding the use of different Sigstore components. P13 (security tooling, internet service) highlights this, “ I would say the documentation lacks a bit. Just I think if you look at the documentation, it was updated quite a long time ago, like the ReadMe file. And so if we could have these different options saying that, okay, if you are starting over, this is how you can do rekor and Fulcio.”
Integration to Other Systems: Participants also expressed a desire for a wider range of compatible technologies beyond what Sigstore currently supports. Commonly mentioned integrations included other CI/CD platforms such as Jenkins and GitLab, as well as attestation storage databases. These limitations were often attributed to Sigstore being an emerging technology, as highlighted by P12 (SAAS, SSC security tool), “The weakness is not everything supports it. There’s a lot of weaknesses just around not just it’s being a newer technology”
Fulcio Issues: Some Sigstore issues were also tied to its certificate authority’s (Fulcio) time stamping capabilities, and implementation of OIDC for keyless signing. Fulcio acts as an intermediary using OIDC identities to bind to a short-lived certificate, this was criticized by P3. P3 also commented on the initial lack of time stamping abilities for the initial versions of Fulcio. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “I think we do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. And then I think the other big thing that we did, because Sigstore originally didn’t support it, was time-stamping. Because we’re using Fulcio, we want to ensure that the signature was created when the certificate was valid. And so we use external time-stamping authority in order to be able to do that.”
Offline Capabilities: Another reported limitation of Sigstore is the absence of offline verification capabilities using offline keys. This issue is related to the transparency log’s inability to function in air-gapped environments. However, in this case, participants specifically desired the ability to use Sigstore’s signing and verification features in offline scenarios.
Software Libraries: The programming libraries of Sigstore were also mentioned as a challenge, particularly in cases where users wanted to integrate Sigstore’s signing capabilities directly into an application.
5.1.3 Effect of Identified Problems on Sigstore components used by Subjects
Table 5: Sigstore Functionalities Used by Subjects
Tool # of Subjects
Cosign (Signing & Verification CLI) 9
Certificate Authority (Fulcio) 9
Keyless Signing & Identity Manager (OIDC) 9
Transparency Log (Rekor) 8
Gitsign (Commit Signing) 4
Component Customization & Local Sigstore Deployment 2
We assessed the effect of participants’ reported problems with Sigstore on their usage of each component. Table 5 summarizes our subjects’ reported usage of Sigstore components and functionalities.
Key components like Cosign, Fulcio, and keyless signing saw high adoption, while local Sigstore deployment had low usage due to numerous reported issues. Several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table 4, section 5.1.2).
The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues. Of the eight participants who used it, three were still in the pilot phase. Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other crucial elements was notably absent. In particular, there was very little mention of log witnessing [28], or monitoring [22] for either Fulcio or Rekor. This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior. We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
5.2 RQ2: How and Why Practitioners Change Software Signing Tools
Table 6: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
Topics & Associated Examples Subjects
Human/Social Factors
Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P8, P10
GPG Issues 6
Low adoption rates P1, P10
Key management issues & Other usability concerns P1, P6, P9, P12, P13
Steep learning curve P9, P12
Compatibility with newer technology P12
Notary Issues 2
Non-demand from customers P9
Compatibility with other tools P9
Lack of regular updates P9
Key & Identity Management P5
Proprietary Tool Issues 1
Difficult to setup P11
Technological Factors
Available Sigstore Functionalities 3 – P1, P5, P10
A transparency log, etc P5, P10
Compatibility to other Tools P1
Macroenvironmental Factors
Regulation & Standards 4 – P5, P6, P11, P13
Large User Community 1 – P8
Inherent Trust of Creators 1
Trust of CNCF products P3
Figure 3: Practitioners’ Evolution of Software Signing Tools. Most reported their organizations moved from no former tools, a form of GPG signing or notary to Sigstore. Most participants depicted with the flow from none to a sigstore implementation report that; they either initially used no tools, or have always used Sigstore.
We also tackle the usability of software signing tool especially as it relates to decisions that prompts the switch of practitioners in using different tools. In the context of our study, we asked participants what factors prior to adopting Sigstore influence practitioner’s to switch to Sigstore. The factors here are mostly Macroenvironmental (relating to regulations, user communities), Human (Previous experiences), and Technological (available capabilities of sigstore)
First, we present a Sankey flow diagram ( Figure 3) illustrating the evolution of signing tool usage among our participants’ teams up to the present. A significant portion of participants (6), as depicted transitioned from using no tools to adopting Sigstore —either they had always used Sigstore since the inception of their organization or continued using their initial signing tools alongside Sigstore. The most notable transition was from GPG-based signing implementations (3) to Sigstore, followed by a shift from Notary (2). Other reported transitions included Skopeo (1), unspecified internal tools (1), and a proprietary tool (1).
The reasons most of our participants chose to switch to Sigstore stemmed from a combination of two key factors: issues related to a subpar user experience with previous tools and advantages associated with Sigstore. We summarize these factors in Table 6.
Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore. Some participants (e.g., P6) stated that their organizations have a policy of contributing to open-source communities they intend to adopt. Others noted that their organizations were relatively new compared to Sigstore—meaning that Sigstore had already existed before these organizations were established, and some of their members had contributed to Sigstore before joining their respective organizations. Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project. P10 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community. There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
Available Sigstore Functionalities: Sigstore’s advertised functionalities also influenced its adoption. While several participants were drawn to the transparency log (as highlighted in section 5.1.1), its integrability with existing technologies they were using at the time also played a role in their decision. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “… Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, as well as they also maintain a transparency log as well …” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P1 (POC software, digital technology): “Factors I considered were, …, the existence of tools to use it.”
Regulations and Standards: The impact of regulations and standards on adopting Sigstore was significant among our participants. Some participants who previously used other tools (2) switched to Sigstore due to security frameworks and standards recommending it as a best practice. Additionally, the remaining participants (2) noted that regulations played a partial role in their decision to adopt Sigstore.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “…standards actually really helped to convince everyone that we should be doing this … But once a standard is put in place, once the requirements are set, then it just sped up the process of adopting these new technologies like Sigstore …”
User Community & Trust Of Sigstore Creators: The large user community of Sigstore was also a key reason for its adoption among participants. This aligns with feedback from some participants (P1, P10) who cited GPG’s smaller user base as a drawback. Additionally, some participants expressed inherent trust in the CNCF (Cloud Native Computing Foundation) security tooling community, which they had grown to rely on over time. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “We do defer often at the higher level projects to different foundations. So we would be more likely to trust something from the CNCF, the Cloud Native Computing Foundation or Sigstore rather than small little independent developer projects or things like that.”
GPG & Notary Issues: Issues associated with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore. GPG-based signing implementations have historically been criticized for key management challenges and usability concerns. Our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies.
Notary-related concerns primarily revolved around customer demands, compatibility with other technologies, and infrequent updates. Key and identity management challenges were common to both GPG and Notary. Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
6 Discussion
In this section, we highlight two reflections from our data.
The role of community in software signing tool adoption: A substantial number of participants reported that they adopted Sigstore because of their involvement as contributors to the Sigstore tooling community (section 5.2). This reflects a key macroenvironmental factor from our usability frameworks that influenced users’ decisions to adopt these tools. Similarly, several reported challenges were linked to the lack of community support in setting up private Sigstore instances and accessing usage information (section 5.1.2).
These findings highlight the critical role of community, as emphasized in prior research [72]. New adopters often depend on communities they trust, contribute to, or receive support from, which significantly shapes their decision to adopt such tools. Toolmakers should prioritize fostering and maintaining a strong community around their tools, as this can greatly influence adoption and success.
Privacy Concerns: Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern. This issue also affects the use of OIDC for keyless signing. While this challenge is not unique to Sigstore [52] and is a recurring issue in software signing and modern cryptography, it underscores the need for continued innovation to address privacy-preserving mechanisms that balance transparency with user privacy, ensuring broader adoption and trust in cryptographic tools like Sigstore. While works like Speranza[52] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
7 Conclusion
Software signing allows software provenance to be guaranteed through cryptographic primitives. This property would address many software supply chain attacks. However, for decades, the widespread adoption of software signing has eluded the software industry. In this study, we examined a recent change in software signing practices by focusing on Sigstore, the leading next-generation software signing platform. We interviewed industry experts to understand Sigstore from a usability perspective. We identified the factors they weigh and the challenges they face. Our results shed light on usability issues for software signing in particular, and will inform more usable software tools writ broad.
8 Acknowledgments
We thank the study participants for contributing their time to this work. We acknowledge support from Google, Cisco, and NSF #2229703. We also thank Hayden Blauzvern and all reviewers who provided feedback on this work.
Appendix A Interview Protocol
Table 1 gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
We include Section D of the protocol for completeness, although it was not analyzed in this study.
Table 7: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions. We only analyze parts A and D in this study
Themes Sub-themes Questions
A: Demographic A-1: What best describes your role in your team? (Security engineer, Infrastructure, software engineer, etc)
A-2: What is your seniority level? How many years of experience?
A-3: What is the team size?
A-4: $\dagger$ What are the team’s major software products/artifacts?
A-5: What type of Organization/company?
B: Software Supply Chain
Failure Experienced
B-1: Describe briefly the team’s process from project conception to product release and maintenance – This is to understand the unique context of each practitioner’s software production case.
B-2: $\star$ What do you consider a Software Supply chain attack/incident to be?
B-3: Can you describe any specific software supply chain risks (or incidences with third-party dependencies, code contributors, open source, etc.) that your team has encountered during your software development process? How were these addressed? (Alternatively – How did this affect the decision to implement software signing?
[…]
B-5: What is your team’s greatest source of software supply chain security risk? Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from Section B): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
Third-party dependencies C-2: What are the team’s peculiar selection strategies for Third-party dependencies? How does the presence of a signature influence this decision? How is the authenticity of this signature verified? Any other methods/practices to ensure the trustworthiness of the third-party dependencies before integrating them into your projects? Possible Follow-up – Can you describe these methods/practices?
C-3: What influences the team to discontinue the use of a package? How does the Signature on the dependency influence this decision?
C-4: How does the team manage its third-party dependencies’ security? How do you maintain awareness of potential vulnerabilities or threats related to third-party components?
First-party source code C-5: Is software signing a requirement for team developers (insider threats) and open-source contributors (if any)? how does the team use software signing to protect its source code (what parts of the process is signing required e.g. Commit signing)?
Build Process C-6: (How) Does the team utilize signing in its build process?
Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism? (This question is derived from  [36]).
D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
D-2: What factors did the team consider before adopting this tooling over others?
D-3: What was the team’s previous signing practice before the introduction of current tooling?
D-4: How does your team implement this tooling (which components of tool does the team use)?
D-5: Have you encountered any challenge(s) using this tool of choice? How have you coped with this limitation(s)? Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
Appendix B Codebooks
Table 8 describes our codebook, with code definition, and example quote tagged with that code.
Table 8: Excerpts from the final codebook we used for our analysis.
Code Definition Sample Quote
Sigstore Criticisms Criticisms and weaknesses associated with implementing Sigstore for signing. S3: "We do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider. "
Sigstore Improvements Suggested improvements for Sigstore. S9: "I would say probably integrations within different platforms, so say you use GitLab or GitHub or Jenkins or something like that. Try to leverage some type of integration to make that, I guess, less friction for the developers to implement, something like that."
Sigstore Strengths Strengths of Sigstore and reasons why subjects prefer using Sigstore. S7: "The convenience and the good user experience is probably what keeps us using Sigstore "
Sigstore Alternatives Considered Alternatives of Sigstore that were or are being considered by the subject (for subjects whose primary signing implementation is Sigstore). S5: "We are also looking at other new alternatives that are coming up, like OpenPubkey. Docker just released a new standard called OpenPubKey, which is trying to solve this issue where they will not require you to have infrastructures and stuff."
Sigstore Components Used Components of Sigstore (Cosign, Fulcio, etc.) being used by the subject. S2: "Mostly fulcio is what’s integrated into our product."
Notary Criticism Criticisms and weaknesses Associated with Notary – promoting users to stop using them. "S5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
Appendix C Traditional vs Next-Gen Software Signing Workflow
We show a typical traditional package signing workflow in Figure 4 compared to a next-generation workflow in Figure 5.
Figure 4: Typical workflow for software signing and verifying signatures. The component author packages (A) and signs (B) their software. The signed package (C) and public key (D) are published. To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
Figure 5: Typical workflow for next-generation software signing. The component author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
Traditional Signing: The signature creation process begins when the maintainer’s artifact is ready for submission. The signer generates a key pair that provides them with a private key for signing their software artifact and a public key for a verifier to validate the artifact’s signature. Once the artifact is signed, it must be submitted to a package registry (such as PyPi, npm, or Maven) or the intended party. During this phase, the signer is responsible for making their public key available and discoverable so future users can verify the signer’s identity. In the verification phase, the verifier retrieves the signer’s public key, uses it to decrypt the signature file, and checks for equivalency with the software artifact’s hashed digest.
Next-generation Signing:: Keyless/Ephemeral software signing mitigates the security risk of private key compromise. Keyless software signing differs from traditional public-key methodology in the fact that the signer does not rely on a long-term private key to sign their messages. Instead of the signer being responsible for the secure maintenance of a private key, in keyless signing, the signer uses a short-lived key that only lasts long enough to sign their artifact [2].
Appendix D Survey Attempt
In addition to our interviews, we developed a survey distributed to Sigstore users at the 2024 SigstoreCon conference and through post-conference communications with the Sigstore user community. This aimed to gather perspectives from those who might not have had time for an interview. We exclude this data from our work due to several factors that compromise the quality of the results; these factors include respondents who completed the survey in an unrealistically short amount of time and those who failed to complete the entire survey.
Here, we summarize the answers to an excerpt of the questions responded to in our survey:
What signing tool did your team use prior to the introduction of Sigstore?: Among 7 participants who utilize Sigstore, 3/7 responded that they use PGP/GPG-based signing and 4/7 did not use any toolings prior to using Sigstore.
Which of the following factors have influenced your team’s use of signing? (select all): Among 7 participants who utilize Sigstore, 4/7 selected "Organizational policies," 4/7 with "Best Practices/Other standards and recommendations" and 3/7 selected "Awareness of prominent SSC attack."
Which Sigstore components does your team use? (select all): Among 7 participants who utilize Sigstore, 3 responded to this question, and all 3 participants selected Cosign, Fulcio, and Rekor.
In which of the following ways have you implemented Sigstore?: Among 7 participants who utilize Sigstore, 3 responded to this question, with 2 responding that they use a private deployment of Sigstore and 1 responding that they use a public instance of Sigstore.
Do you use or require Sigstore (signing) at any of the following stages of your product?: Among 7 participants who utilize Sigstore, 3 responded to this question. All 3 participants shared the "Build/Deployment/final product" stage as part of their answer.
How do you verify Sigstore signatures? (select all): Among 7 participants who utilize Sigstore, 3 responded to this question. All 3 participants shared selecting certificate Verification as part of their answer. 2/3 selected that they check Rekor logs.
Appendix E Code Saturation
Figure 6: Saturation curve. Interviews are plotted in the order in which they were conducted. Each interview covered many topics in detail (orange line, blue line). However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
We present a saturation analysis of our studies in  Figure 6. The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview. We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 11. Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4. This is also reflected in the orange trendline (unique number of codes).
Appendix F Code Evolution Diagram
Figure 7 describes the evolution of our codebook over six rounds of revisions.
Figure 7: Codebook Evolution: The codebook underwent six rounds of revisions, with changes highlighted by different colors: round 2 (pink), round 3 (blue), round 4 (green), round 5 (yellow), and round 6 (purple). The final codes from round 6 were used in this paper, and categorized thematically. We combined framework and thematic analyses to develop the codebook. First, we induced codes from each transcript and analyzed them thematically. During the framework analysis, we mapped codes to the four stages/blocks of the software supply chain factory model. Round 1: The codebook was defined based on emerging themes. Round 2: No changes. Round 3: Blue-highlighted codes (5) were added after memoing and coding the second set of transcripts. Round 4: Green-highlighted codes (2) were added. Round 5: No changes. Round 6: Analysts A1 and A2 determined some codes should be further divided. Purple-highlighted codes (4) were added and the codebook was finalized. We focused on the topics of relating to sigstore(signing tools) including, Notary, Vault, Spiffe, GPG, and others.

### v4

@@
-Why Johnny Signs with Sigstore: Examining Tooling as a Factor in Software Signing Adoption in the Sigstore Ecosystem
+leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false
 Abstract
-Software signing is the strongest method for ensuring the integrity and authenticity of components in a software supply chain.
-Although software signing has been recommended by regulation and industry consortia, practical adoption of software signing is generally limited.
-Previous studies have recognized tooling as a key factor in influencing software signing adoption and quality.
-However, most research has primarily focused on its user interface aspects, with little attention given to other usability considerations, such as case-specific user challenges and factors influencing tool adoption and changes.
-To understand how software tools influence the practice and adoption of software signing, we study the formative usability of Sigstore, a modern and widely adopted software signing tool.
-We interviewed thirteen (13) experienced security practitioners to study the factors that influence the selection of a tool, the problems associated with the use of such tools, how practitioners’ software signing tools have evolved, and what drives this migration.
-To summarize our findings: (1) We highlight the various factors practitioner’s consider before adopting a software signing tool; (2) We highlight the problems and advantages associated with the current tooling choices of practitioners; and (3) We describe the evolution of tooling adoption of our sample population.
-Our findings provide the software signing tool development community with valuable insights to improve their design of software signing tools.
-1 Introduction
-The reuse of software components in modern software products results in the formation of complex software supply chains, leading to an increased attack platform for these products [39, 79, 67].
-A key security strategy to mitigate this expanding attack surface is ensuring the authenticity of actors and integrity of artifacts by establishing provenance [54, 68].
-Software signing is the strongest guarantee for establishing provenance [69] and has been extensively advocated by academia, industry consortia, and government regulations [81, 16, 14].
-To effectively implement this process, software signing tools–such as OpenPGP [55], Sigstore [53], and Notary [57]–have been proposed and are currently used in practice to sign software artifacts.
-Although tooling has significantly advanced various technological fields [71, 41, 3, 47], its impact on software signing must be carefully examined to maximize its benefits.
-While prior "Why Johnny Can’t Encrypt" inspired studies have explored the usability of signing tools, their focus has been limited primarily to user interface design [89, 73, 63].
-Furthermore, their evaluation techniques rely primarily on summative usability assessments (ease of use), which have been criticized for being difficult to measure and for providing limited insights to inform tool enhancements [13, 48].
-The applicability of these studies is also constrained by their narrow focus on message (and email) signing and their reliance on novice volunteer participants, which limits the real-world relevance of their findings for software signing.
-Our work provides the first examination of the usability of tooling as a key factor influencing the adoption and practice of software signing.
-We conducted interviews with 13 experienced security practitioners to explore current software signing tools, with a focus on Sigstore as a case study.
-We apply a formative usability framework  [17], which emphasizes the iterative refinement of the studied tooling option, to develop our study instruments.
-Our study explores usability in two categories: users’ experience with the specified tooling option and the usability of the tooling options with respect to other software signing tools.
-Our study thus explores three key themes: the evolution and motivations behind software signing tool adoption, limitations and challenges encountered with currently adopted tooling, and the specific implementations of adopted tools.
-Our results highlight the factors practitioners consider when adopting specific software signing tools, using a case study of a key software signing tool.
-This serves as formative feedback for tool designers, offering insights into the drivers of tool usage.
+Software signing is the most robust method for ensuring the integrity and authenticity of components in a software supply chain.
+However, traditional signing tools place key management and signer identification burdens on practitioners, leading to both security vulnerabilities and usability challenges.
+Next-generation signing tools such as Sigstore have automated some of these concerns, but little is known about their usability and adoption dynamics.
+This knowledge gap hampers the integration of signing into the software engineering process.
+To fill this gap, we conducted a usability study of Sigstore, a pioneering and widely adopted exemplar in this space.
+Through 18 interviews, we explored (1) the factors practitioners consider when selecting a signing tool, (2) the problems and advantages associated with the tooling choices of practitioners, and (3) practitioners’ signing-tool usage has evolved over time.
+Our findings illuminate the usability factors of next-generation signing tools and yield recommendations for toolmakers, including: (1) enhance integration flexibility through officially supported plugins and APIs, and (2) balance transparency with privacy by offering configurable logging options for enterprise use.
+I Introduction
+The reuse of software components in modern products creates complex supply chains and expands the attack surface for these systems [1, 2, 3].
+A core strategy for mitigating this risk is to establish provenance—verifying actor authenticity and artifact integrity—through software signing [4, 5].
+Software signing is the strongest guarantee of provenance [6] and is widely advocated by academia, industry consortia, and government regulations [7, 8, 9].
+High-profile supply-chain attacks—such as the SolarWinds breach—have further underscored the need for stronger, more automated signing infrastructure.
+In response, tools have evolved into two classes: traditional solutions (e.g., OpenPGP [10]), which delegate key management and signer identification to practitioners, and next-generation, identity-based tools (e.g., Sigstore [11]), which further automate the process via short-lived keys and external identity providers.
+Tooling has significantly advanced various technological fields [12, 13, 14, 15] and has often played a key role in the adoption of these technologies [16, 17].
+There is a continual need to evaluate the adoption practices and usability of such tools to inform practitioners, potential users, researchers, and toolmakers of their effectiveness in different contexts [18].
+Traditional software signing tools have been studied in this way, most notably in several “Why Johnny Can’t Encrypt” inspired studies[19, 20, 21].
+Although next-generation signing tools promise greater automation, there remains a lack of comprehensive usability and adoption analyses for these solutions.
+Such studies could assess how automation affects adoption suitability, elicit practitioners’ desired improvements, and guide future tool design and research.
+Our work offers the first examination of next-generation software signing tools’ usability as a critical factor influencing their adoption and practice.
+Building on and improving previous methodologies — which relied on novice users and summative usability assessments criticized for their lack of actionable feedback[18, 22] — we conducted interviews with 18 experienced security practitioners to study current software signing tools, focusing on Sigstore as a case study.
+We analyzed our data with a formative usability framework [23], focusing on two usability dimensions: practitioners’ experiences with Sigstore and its usability relative to other signing tools.
+From these interviews, we distilled three key themes: the evolution and motivations behind tool adoption, the limitations and challenges of currently adopted tools, and the specific implementation practices shaping real-world signing workflows.
+Our results highlight the usability factors practitioners consider when adopting next-generation software signing tools.
+This serves as formative feedback for tool designers and researchers, offering insights into the drivers of tool usage.
 We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools.
-Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of modern Software signing tools.
+Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of next-generation Software signing tools.
 In summary, we make the following contributions:
   1. 1.
-    We report on the difficulties and advantages of using Sigstore, a next-generation software signing tool (section 5.1.1, section 5.1.2).
+    We report on the difficulties and advantages of using Sigstore, a next-generation software signing tool.
   2. 2.
-    We discuss how practitioners’ software signing tools evolve over time and the factors that influence practitioners’ decisions to switch tooling options (section 5.2).
-2 Background & Related Works
-In this section, we discuss the foundations of software signing (section 2.1), usability studies in software signing (section 2.2), and the Sigstore ecosystem where we situate our work (section 2.3).
-2.1 Software Signing
-2.1.1 Signing Overview and Workflow
-Software signing is a process that ensures the authorship and integrity of a software artifact by linking a cryptographic signature created by the software maintainer with their software artifact.
-Fundamentally, software signing is based on the public-key cryptography model  [69], and a traditional signing workflow follows a three-stage process of creation, distribution, and verification.
-We describe a typical signing workflow in Appendix C.
-While the fundamental software signing process remains consistent across implementations, different signing tools and policies are shaped by the target platform, software artifacts, and root of trust.
-The choice of signing tools and the policies under which developers operate often depend on the operating system and distribution vendor.
-For example, Microsoft Authenticode [30] and Apple Xcode [33] provide platform-specific signing mechanisms, while package registries like PyPi [45] and Maven [1] enforce signing policies through dedicated tooling plugins.
-Developers also have a variety of software signing tools available to them, such as GNU Privacy Guard (GPG) [58], Signify [85], Sigstore [53], and Notary [57].
-Each of these tools present different advantages and challenges, shaping how developers integrate signing into their workflows.
-2.1.2 History and State-of-Art Tooling
-Traditional Signing Tools: Many traditional signing tools, such as Git Commit Signing and GNU Privacy Guard are based on PGP.
-PGP (Pretty Good Privacy) is a proprietary encryption software originally developed by Phil Zimmerman in 1991.
-It is a public key cryptography-based protocol whose capabilities include encrypting files, creating and managing public and private keys, and creating digital signatures.
-Its open-source standard is OpenPGP [23, 55].
-However, two main struggles associated with PGP-based software signing involve key management and establishing trust in public keys.
-PGP relies on a "web of trust" model, where a signer’s public key is considered trustworthy based on endorsements from other trusted signers [23].
-Each created public key is stored in a "key certificate" that specifies the public key, the name and/or email address associated with the public key, its creation date, and others who have signed the key.
-The signer is also responsible for secure management of their private key, preventing exposure, as the combination of their private key and passphrase enables the signing of artifacts [45].
-State-of-art Signing Tools: Although software signing is widely recognized as an effective technique for mitigating software supply chain attacks, several obstacles hinder its adoption.
-Research highlights the prevalence of unsigned artifacts in software registries, emphasizing the need for improved signing practices [53, 90, 46].
-This has led to the emergence of the next generation of software signing.
-Newman et al.
-introduced a new software signing platform in 2021 called Sigstore that addresses key adoption barriers, particularly key management, by leveraging keyless/ephemeral software signing techniques [53].
-Keyless signing mitigates security risks associated with private key compromise.
-Instead of relying on a long-term private key, the signer uses a short-lived key that exists only for the duration of the signing process.
-This approach enhances security by preventing private key leaks while improving accessibility and usability [2].
-As a result, many software registries are transitioning to this model to streamline verification and enforce stronger software integrity measures [15, 34].
-While Sigstore’s development aims to solve the usability issues associated with traditional signing methodology, firsthand perspectives from practitioners using Sigstore will provide valuable insight into its actual usability and convenience.
-2.2 Usability Studies in Software Signing
-2.2.1 Usable Tools: Definition and Theory
-Usable Tools: In basic terms, a software engineering tool assists developers in their software production process [27].
-For software signing specifically, these tools are used to provide a method for which users can verify the integrity and authenticity of the code to be published, and developers use these tools to verify signatures of the software artifacts they may incorporate into their own projects.
-Deciding which software engineering tools to use is a critical question developers face when many options are available.
-A specific tool can affect a developer’s workflow efficiency, productivity [44], and the resulting quality of the work that is produced.
-This allows members of the computing industry to remain competitive by selecting the most suitable tools and environments that are conducive for the efficient production of software [27].
-Additionally, tooling is important because it enables engineers to use software development practices they have not in the past [11].
-Usability Evaluation: Studies assessing implementation strategies of products are typically framed as usability evaluations [49].
-Usability evaluations are generally categorized into two forms: summative (validation or ease of use) and formative (diagnostic usability) [48].
-Summative usability refers to the extent to which a system or product can be used to achieve a specified goal effectively, efficiently, and satisfactorily within a given context [35, 65].
-In contrast, formative usability focuses on product assessments aimed at identifying areas for improvement [82].
-Although summative usability evaluations are more commonly used, they are not designed to provide actionable feedback for enhancing the product [48].
-Summative evaluations primarily measure metrics such as effectiveness, efficiency, and user satisfaction [35].
-However, such metrics have been widely criticized for being subjective and difficult to quantify [48, 13].
-In contrast, formative usability evaluations aim to diagnose user challenges and provide iterative feedback to enhance product functionality [82, 48].
-These evaluations emphasize real-world usability issues and are more suited for refining a product’s design.
-While summative evaluations are more common, their limitations highlight the need for formative assessments when improving software signing tools.
-2.2.2 Empirical Studies
-Signing Tooling Usability Studies: The usability of digital signing tools has been widely studied, though often within the context of encryption rather than broader software signing applications.
-Whitten and Tygar’s seminal study, Why Johnny Can’t Encrypt [89], found that PGP 5.0’s user interface was too complex for non-expert users, making encryption and signing inaccessible.
-Sheng et al.
-[73] later observed that PGP 9’s interface had not significantly improved, particularly in key verification, encryption transparency, and digital signing usability.
-Other studies [10, 64, 60] have highlighted persistent usability issues, including the complexity of the public-key model.
-Further research has explored the differences between manual and automated signing tools.
-Fahl et al.
-[19] and Ruoti et al.
-[62] found no significant usability differences between these approaches.
-Meanwhile, Atwater et al.
-[5] demonstrated that users preferred automated solutions based on the PGP standard.
-While these studies provide insights into encryption tools, their relevance to broader software signing workflows is limited.
+    We discuss how practitioners’ choices of software signing tools change over time and the factors influencing those decisions.
+Significance of Study: Our work targets Sigstore—a pioneer next-generation signing tool that has seen rapid uptake across major open-source ecosystems[24, 25, 26], cloud-native projects, and corporations[27, 28, 29, 30, 31], involving a large number of engineers worldwide.
+By unpacking the usability factors that drive or hinder Sigstore’s adoption, we not only inform the ongoing refinement of its automation workflows (e.g., certificate issuance and key management) but also provide a template for evaluating and improving similar Next-generation(identity-based) signing solutions.
+Also, since Sigstore exemplifies a broader class of tools that embed security into automated build and deployment processes, our findings can be potentially generalized to guide the design and implementation of other automated software-supply-chain technologies, ultimately advancing the usability—and thus the security—of automated software engineering practices.
+II Background & Related Works
+In this section, we review software signing technologies (section II-A) and usability assessments (section II-B).
+II-A Software Signing
+Software signing ensures the authorship and integrity of a software artifact by linking a maintainer’s cryptographic signature to their software artifact.
+The implementation of software signing is shaped by several factors: the software artifacts to be signed (e.g., driver signing[32]), the root-of-trust design (e.g., public key infrastructure/PKI [33]), and ecosystem policies (e.g., PyPi[34] and Maven[35]).
+II-A1 Software Signing Tools
+Software signing tools can be classified into two groups: traditional and next-generation (identity-based) [5].
+Traditional signing tools such as GPG [36] follow a conventional public-key cryptography approach, relying on long-lived asymmetric key pairs and delegating key-management responsibilities to the signer.
+Trust is established through endorsements by other signers.
+Next-generation (identity-based) signing tools automate key management by issuing short-lived, ephemeral certificates tied to a signer’s identity via external OpenID Connect (OIDC) or OAuth providers.
+This “keyless signing” approach — exemplified by Sigstore [11], OpenPubKey [37], and SignServer [38] — eliminates long-term key handling and embeds transparency through append-only logs, improving usability and security.
+We discuss the architecture and workflow for next-generation signing tools in section IV-B using the case of Sigstore.
+II-A2 Effects of Signing Tools on Software Signing Adoption
+Previous studies recognize that better tooling influences software-signing adoption.
+For example, Schorlemmer et al.
+[6] performed a measurement study of several software package registries, identifying tooling as a key determinant of signature quality.
+Similarly, Kalu et al.
+[39] explore implementation and adoption frameworks in commercial settings, highlighting tool usability, compatibility, and integration challenges.
+While these works emphasize broad usability factors, they offer little detailed guidance on what makes a signing tool usable, overlooking the fine-grained technical and workflow requirements that drive real-world tool adoption.
+II-B Usability Studies in Software Signing
+II-B1 Usability – Definition, Theory & Evaluation Frameworks
+Tools enable engineers to adopt new software development practices [40].
+A usable tool is one whose functionality effectively supports its intended purpose—achieving product goals [18, 41].
+Studies assessing implementation strategies of tools are typically framed as usability evaluations [42].
+To evaluate a tool’s usability, studies typically follow a usability evaluation framework [42].
+There are two kinds: summative, which measures a tool’s effectiveness [43, 44], and formative, which provides iterative feedback for improvement [18, 45].
+Given the aims of this work, we took a formative approach.
+We considered several formative frameworks [46, 47, 48, 49] and found that Cresswell et al.’s four-factor framework [23] helped explain our data the best.
+That framework is summarized in Figure 1.
+Figure 1: Cresswell’s Usability Framework [23].
+It assesses usability based on four factors.
+Technological (T) concerns the technical characteristics of the system (e.g., performance, ease of use) and data management (availability, integrity).
+Social/Human (P) relates to users and their interactions with the tool, e.g., user satisfaction, engagement, prior experience, and correct use of features.
+Organizational Context (O) are factors internal to participants’ organization, e.g., training and leadership support.
+Macroenvironmental (M) are external influences, e.g., regulations, economics, and vendor relations.
+II-B2 Empirical Studies
+The usability of traditional signing tools—often studied in the context of encryption—remains a significant challenge.
+Whitten and Tygar’s seminal study, “Why Johnny Can’t Encrypt” [19], found PGP 5.0’s interface too complex for non-experts, a problem that persisted in PGP 9 [20], especially around key verification, transparency, and signing.
+Subsequent work [50, 51, 52] has underscored the public-key model’s inherent complexity.
+Research comparing manual versus automated signing tools has yielded mixed results: some works observed no clear usability advantage [53, 54], whereas Atwater et al.
+showed a user preference for automated PGP-based solutions [55].
+Although informative, these studies focus on traditional encryption scenarios (e.g., email and messaging), limiting their applicability to broader software-signing workflows.
 These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing.
-First, they primarily focus on evaluating the usability of signing tooling in the context of email or message encryption, limiting their application to software signing.
-Additionally, they limit their scope to user interface design, overlooking critical factors such as tooling adaptability and macroenvironment (e.g., tooling user communities).
+First, they primarily focus on evaluating the usability of signing tooling in the context of signing communication (e.g., email, messages), limiting their application to signing software (i.e., securing the software supply chain).
 Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences.
-Their methodology predominantly employs summative usability assessments, which focus on ease of use [13].
-These studies typically rely on novice volunteers to evaluate usability.
-In addition to the flaws of the summative evaluation approaches, the reliance on novice participants limits the applicability of their findings.
-Evaluating usability through inexperienced users does not yield specialized feedback that can effectively guide tool enhancements [42, 50].
-2.2.3 Related Software Signing Research
-The use of software signing as a security method to mitigate Software Supply Chain security has been studied in the literature.
-These studies has focussed on the adoption and practice of Software signing.
-Schorlemmer et al.
-[69] examine factors influencing software signing adoption through a framework of signing events.
-They identify tooling as a major factor affecting the quality of software signatures but focus primarily on the role of dedicated tooling within software registries, rather than its broader impact on engineering workflows.
-In our parent study, Kalu et al.
-[37], we provided valuable insights into the implementation and adoption frameworks of software signing, with a focus on broad software signing challenges faced in commercial settings.
-These challenges included issues of usability, compatibility, and integration.
-However, while we highlighted these broad challenges, our findings lacked the specificity needed to identify which tools are most impacted by these issues and under what conditions of use, limiting practitioners’ contextual understanding of these problems.
-These prior studies acknowledge that improved tooling can drive the adoption of software signing.
-However, they fail to clearly define what constitutes an "improved tool" from the perspective of practitioners.
-While they emphasize broad usability factors, they overlook the specific technical and workflow-related requirements that practitioners consider when selecting tools for real-world use.
-2.3 Sigstore
-Due to the emergence of next-generation software signing and the growing adoption of Sigstore as a signing platform, we focus on the Sigstore tool specifically in our research.
-Importance: Sigstore, launched in 2021, was developed to address critical challenges in software signing, including complex key management, usability barriers, and software supply chain security risks.
-Traditional signing tools, such as OpenPGP and enterprise certificate authorities, rely on long-term cryptographic key storage, which introduces risks of key compromise, revocation difficulties, and usability friction.
-In contrast, Sigstore simplifies the signing process by providing keyless, ephemeral authentication mechanisms.
-Within 13 months of its public release, Sigstore recorded 46 million artifact signatures  [29].
-It has since gained significant adoption and support from open-source ecosystems  [84, 25, 31, 6, 86] and various organizations  [38, 76, 75, 77].
-Given its growing adoption and integration in major software ecosystems, Sigstore represents a state-of-the-art software signing approach, making it a compelling case study for examining usability and security trade-offs.
-Design Aspects Affecting Usability: One key challenge that is associated with traditional signing methods is key management, which is addressed by Sigstore.
-In conventional certificate authority-based signing, such as Windows Authenticode, a software artifact’s validity is linked to the certificate’s status—whether it is still valid, expired, or revoked [43, 70].
-A revoked certificate can indicate compromise, but it may not be immediately clear when the compromise occurred [40].
-This uncertainty allows some revoked certificates to be used for signing malware while still appearing valid to verification systems.
-Sigstore mitigates this risk by issuing short-lived certificates.
-These certificates significantly reduce the chances of key and certificate compromise since they automatically expire after a short duration, limiting their potential misuse.
-Complexity: Sigstore uses a modular design of several components to implement software signing.
-As described by [53], Sigstore consists of eight key components: signers, verifiers, the artifact log, an OpenID Connect (OIDC) provider, a certificate authority, monitors, transparency logs, and package repositories.
-While Sigstore is designed to function as a cohesive system, these components can also be implemented individually based on specific use cases.
-The signing process starts with the signer using OpenID Connect (OIDC) to obtain a signing certificate.
-Unlike traditional methods requiring long-term key management, Sigstore associates a signer’s identity with short-lived tokens issued by services such as Google, GitHub, and Microsoft.
-Next, a software signing certificate is issued to the signer based on their verified identity.
-In this step, Fulcio serves as the certificate authority (CA), ensuring that the OIDC token was correctly issued to the requesting signer before granting the certificate.
-To monitor signing behavior, a verifier makes use of Fulcio’s log and the transparency log (Rekor) to verify for identity of the creator of the signatures, the certificate validity, and signature validity.
-3 Knowledge Gaps & Research Questions
-The existing literature on software signing, as outlined in  section 2, has examined its practice and adoption across different contexts, including private and open-source ecosystems.
-These studies highlight the critical role that the choice of a software signing tool plays in ensuring the quality of the software signing process.
-Although certain aspects of software signing tools have been explored, such as the usability of the user interface design and functional exclusivity (dedication of tooling resources), the broader impact and role of software signing tools remain largely understudied.
-The findings of these works underscore that tooling is a multi-faceted and multi-layered factor.
-A general usability evaluation of these tooling options would significantly enhance their adaptability and facilitate improvements.
-Also, the existing research on software signing tooling primarily relies on experiments with novice users and quasi-experiments to investigate isolated aspects of tooling.
-Practitioner perspectives on the nuanced role of tooling in shaping the use and practice of software signing remain limited, particularly regarding how the usability of these tooling contributes to their adoption.
-Usability Framework:
-Figure 1: Usability Framework from Cresswell et al.
-[17].
-Their framework assesses usability based on four factors: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors.
-Our framework applies their framework to two concerns of software signing tool usage: (1) the tool itself and (2) previous signing experiences.
-To examine the usability of software signing tooling, an examination of experienced practitioner experiences is necessary.
-As noted in section 2.2, the summative usability approach mostly is mostly hindered by its use of subjective measurement metrics.
-To effectively, answer usability in terms of how software signing tooling are used in practice, a formative usability framework is warranted.
-A formative framework addresses the experiences of regular users of a tool, rather than controlled experience of experimental subjects  [48].
-Our formative usability evaluation framework, Figure 1, is based on Cresswell et al.’s [17] formative framework.
-Their framework identifies four key factors that influence the overall usability of a product or technology: Technological (T), Social/Human (People), Organizational (O), and Macroenvironmental (M) factors.
-Usability is determined by the interplay of these factors.
-Our formative evaluation applies of their four-factor framework to two key areas: users’ experiences with the specified tooling option and the usability of the tooling option compared to previous software signing tools.
-We designed our study instrument to address these broad usability themes, emphasizing the identified usability factors.
-Research Questions: Our research addresses the following questions in light of the positions outlined above within the context of Sigstore:
+Their methodology predominantly employs summative usability assessments that rely on novice volunteers, which limits the applicability of their findings for guiding tool enhancements [56, 57].
+III Knowledge Gaps & Research Questions
+Existing literature on software signing (see section II) has examined practice and adoption across private and open-source ecosystems, highlighting the importance of tool choice in signature quality.
+Prior usability studies have focused exclusively on traditional signing tools and have been limited to evaluating user-interface design.
+However, next-generation signing tools remain understudied, creating a need for:
+  • A holistic evaluation of how software signing tools influences software-signing adoption and practice, and how usability affects this process.
+  • Practitioner perspectives on the nuanced role of tooling in shaping software-signing use and practice.
+Current research (see section II-B2) is limited because it primarily relies on experiments with novice users, while studies in section II-A2 use quasi-experiments or general practitioner interviews without a specific focus on signing tools.
+In light of these gaps, a comprehensive usability evaluation of next-generation signing tools would offer valuable insights into their role in adoption, improve their design, and enhance their adaptability.
+Research Questions: Building on identified gaps in next-generation tooling usability and its impact on adoption, we frame our research around Sigstore—our chosen next-generation signing tool case study—and ask:
   • RQ1:
-    What drawbacks and benefits of software signing tools influence their adoption in practice?
+    What strengths and weaknesses of Sigstore influence its adoption in practice?
   • RQ2:
-    What factors influence practitioners’ decisions to switch software signing tools over time?
-4 Methodology
-Figure 2: Study Methodology.
-We developed our interview protocol by reviewing Sigstore grey literature and the academic literature using our adopted formative usability framework.
-Finally, we analyzed our results using a thematic analysis informed by our usability evaluation framework (Figure 1).
-We undertook a qualitative case study of the Sigstore software signing ecosystem through interviews of engineers.
-We applied a formative usability framework to identify the key factors that impact the adoption of Sigstore in practice.
-Our methodology is illustrated in Figure 2. This study was approved by our Institutional Review Board (IRB).
-4.1 Research Design and Rationale
-4.1.1 Rationale
+    What factors influence practitioners’ decisions to change the software signing tools they use?
+IV Methodology
+We selected a case study method to answer our RQs.
+Our methodology is illustrated in Figure 2.
+This section provides our design rationale (section IV-A), defines the case study context (section IV-B), discusses instrument design (section IV-C), data collection (section IV-D), and analysis (section IV-E).
+Our Institutional Review Board (IRB) approved this work.
+IV-A Rationale for Case Study
 The case study is a research method used to investigate a phenomenon within its real-life context  [58, 59, 60, 61].
@@
 [60].
-We considered the intrinsic uniqueness of the unit (high), access to data (adequate), and risks of participation (low).
-Our research objectives explore the impact of Software Tooling as a factor in the practice of software signing from the viewpoint of practitioners, with a particular focus on how the usability of software signing tools influences their adoption.
-We opted for semi-structured interviews [66] to effectively examine various aspects of tooling, including usability challenges, perceived benefits, and comparative experiences with other signing tools.
-This approach allows us to gather descriptive, open-ended, opinionated, and detailed responses, providing deeper insights into how usability considerations shape adoption decisions.
-4.1.2 Instrument Design & Development
-To develop our interview instrument, we relied on the usability framework (Figure 1) and various signing literature. We outline how these resources informed our instrument design:
+Specifically, we considered the intrinsic value and uniqueness of the case study unit (high), access to data (adequate), and risks of participation (low).
+In section VI we discuss how our findings may generalize.
+Figure 2: Study Methodology.
+We developed our interview protocol by reviewing Sigstore grey literature and the academic literature.
+Finally, we analyzed our results using a thematic analysis informed by Cresswell’s formative usability evaluation framework (Figure 1).
+IV-B Case Study Context – Sigstore
+Following guidelines outlined by Runeson & Höst [58], we describe the importance of our selected case study context and the technical details of the system.
+Use in Practice: Sigstore launched in 2021.
+It provides keyless, ephemeral keys and signer authentication mechanisms to simplify the signing process used by traditional tools.
+Within 13 months of its public release, Sigstore recorded 46 million artifact signatures [25].
+It has significant adoption and support from open-source ecosystems (e.g., PYPI [26], Maven [24], NPM [63]) and many companies (e.g., Autodesk [31], Verizon [29], Yahoo[64]).
+Thus, Sigstore is a major next-generation software signing platform, making it a suitable context for examining usability in modern signing tools.
+Components & Workflow: We detail Sigstore’s architecture and signing workflow, to familiarize the reader with this example of next-generation software signing tools, and to interpret participants’ statements referring to these components.
+Sigstore’s primary components are [11]:
+  • Cosign: CLI and Go library to create, verify, and bundle signatures and attestations.
+  • OIDC Identity Provider: OpenID Connect services (e.g., GitHub, Google, Microsoft) that authenticate signers and issue short-lived identity tokens.
+  • Fulcio (Certificate Authority): Issues ephemeral signing certificates after verifying the signer’s OIDC token.
+  • Rekor (Transparency Log): Append-only ledger that records each signing event and certificate issuance.
+  • Monitors and Verifiers: Services or tools that validate the signer’s identity, certificate status, and signature integrity.
+Figure 3 shows the Sigstore software signing workflow.
+First, the signer authenticates with an OIDC provider to obtain a short-lived token.
+Next, Fulcio verifies this token and issues a one-time signing certificate.
+The signer then uses Cosign to apply that certificate to the software artifact.
+Finally, Rekor logs the certificate and signature, and verifiers retrieve entries from both Fulcio’s log and Rekor to confirm the signer’s identity and ensure the certificate and signature remain valid.
+Figure 3: Sigstore Signing Workflow.
+The software author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider.
+The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
+See section -K for legible figure
+Commonality with other Next-Gen Tools: Sigstore, like other next-generation signing tools such as OpenPubKey [37], AWS Signer [65], and SignServer [38], shares core functionality in signer identity management and automated key handling.
+However, unlike Sigstore, most of these tools do not include a built-in certificate authority (e.g., OpenPubKey) or a transparency log (e.g., OpenPubKey, SignServer).
+Instead, such services must be added via external integrations.
+While our focus on Sigstore allows for an in-depth case analysis, many of its design principles—such as identity-based authentication and ephemeral keys—are shared by its peers, suggesting potential generalizability across next-generation signing tools
+IV-C Instrument Design & Development
+Following guidance from Saldaña [66], we used semi-structured interviews to examine various aspects of tooling usability, e.g., implementation challenges, perceived benefits, and comparison to other signing tools.
+This approach let us pose consistent questions across all subjects, with the ability to probe about unique aspects of each subject’s circumstances.
+To develop our interview instrument, we did not bind our questions to any prior usability framework.
+Sigstore is a nascent technology and in such contexts an exploratory design is recommended [59, 58].
+Only after this exploratory phase did we map our findings onto a formal usability framework, ensuring it reflected practitioners’ lived experiences rather than constraining data collection to preconceived categories.
+Therefore, to construct the interview protocol, we:
   1. 1.
-    Grey Literature: To derive our questions, we primarily examined Sigstore blogs [74], and Sigstore usage reports [8, 86, 31, 38, 76, 75, 77, 15].
-These sources included reports on how various entities implemented Sigstore in their engineering processes, news about major updates, and other related Sigstore developments.
+    Reviewed grey literature—Sigstore blogs [67] and usage reports [68, 27, 69, 28, 29, 30, 31, 70]——to ground our questions in real-world implementations and updates.
   2. 2.
-    Background & Literature Review (section 2): We employed a snowball approach to examine works from top security conferences on software signing, the design of software signing primitives [53, 52], and empirical studies on software signing (all cited in (section 2)).
-Our literature review informed us of existing knowledge gaps and provided insights into the design of modern software signing tools.
-This helped us avoid duplicating questions already addressed in the literature.
-  3. 3.
-    We designed our questions to align with our formative usability framework (section 3).
-We developed these questions to focus on identifying the factors responsible for the adoption of Sigstore in practice.
-Categories D3 and D4 focused on participants’ experiences with their current tools, aligning with the framework’s emphasis on assessing present tooling experiences.
-Meanwhile, categories D1, D2, and D5 captured participants’ experiences with previous software signing tools, offering insights into their transitions and comparisons over time.
-Next, we refined our interview protocols through a series of practice and pilot interviews.
+    Used a snowball review of top security–conference works on signing primitives [11, 71] and empirical studies (see section II), to identify gaps and avoid redundant questions.
+Next, we refined our interview protocols through a series of practice and pilot interviews[72].
 We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
+For example, we moved the question about each participant’s team’s major product to the demographics section to provide context for their responses.
 Our full interview instrument consisted of three main topics (B-D).
-As suggested by the table, topics B and C do not fit into the same analysis framework as topic D.
+Topics B and C did not fit into the same analysis framework as Topic D.
 Thus, following the guidance of Voils et al.
-[88], we analyzed those topics in another work [37].
+[73], we analyzed those topics in another work [BLINDED].
+[^1] That work is blinded for review.
+We attest no duplication of material.
 In this work, we focus on the data for Topic D.
-Topic D consisted of five main question categories, conducted in a semi-structured manner.
-A summary of the questions in our interview protocol is shown in Table 1.
-Table 1: Summary of interview protocol.
-Topic B and C are analyzed in [37].
-We analyze only topic D in this paper.
-The decision to analyze these parts of the protocol separately was made in accordance with best practices.
-We present our full protocol in  Appendix A
+A summary of the questions in our interview protocol is shown in Table I.
+Topic D consisted of five main question categories.
+Questions D3–D4 probe current tooling experiences.
+Questions D1–D2 explore previous tools, and D5 asks about transition factors.
+See section -B for the full protocol.
+TABLE I: Summary of interview protocol. Data from Topics B and C are analyzed in [BLINDED]. We analyze Topic D in this paper. We present our full protocol in section -B
 Topic(# Questions) Sample Questions
@@
 D5. Have you/your team considered switching Sigstore for another tool?
-4.2 Data Collection
-Recruitment: Our study investigates signing tools using a case study of the Sigstore ecosystem.
+IV-D Data Collection
+Population: Our study investigates signing tools using a case study of the Sigstore ecosystem.
 As such, our target participant pool is centered on Sigstore users.
-Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is implemented and the challenges and benefits of Sigstore.
+Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is used in signing, and the challenges and benefits of Sigstore.
 To effectively answer these questions related to decision-making, we required a target population of high-ranking or experienced practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
@@
 We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [75, 76] Steering Committee (ITSC).
-This initial invitation yielded six (6) participants.
-We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (five(5)) and authors contacts (seven) to recruit an additional twelve(12) participants.
-We interviewed 18 subjects, but only responses from 13 participants fit our analysis framework [^1] We interviewed 18 participants but ultimately excluded data from five for this study.
-Two of them used software signing tools other than Sigstore, and three gave limited information about their software signing tool..
-Participant Demographics:
-Table 2: Subject Demographics.
-For anonymity, we used generic job roles, not specific titles.
-“Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff.
-We highlight type of software subject’s immediate team produces.
-* refers to cases where subject’s team built other types of software in addition to the main stated.
-O refers to cases where subject belonged to multiple teams with other products to the one stated.
-We also highlight which participants were recruited from the In-toto Steering committee (I) and which were from snowball sampling (S).
-ID Role Experience Software Type Rec.
-P1 Research leader 5 years Internal POC software I
-P2 Senior mgmt. 15 years SAAS security tool*O S
-P3 Senior mgmt. 13 years SAAS security tool*O I
-P4 Technical leader 20 years Open-source tooling S
-P5 Engineer 2 years Internal security tooling S
-P6 Technical leader 27 years Internal security tooling*O S
-P7 Manager 6 years Security tooling*O S
-P8 Engineer 2.5 years SAAS security* I
-P9 Technical leader 4 years Consultancy S
-P10 Research leader 13 years POC security Software* S
-P11 Senior mgmt. 15 years Internal security tooling I
-P12 Senior mgmt. 15 years SAAS S
-P13 Manager 11 years Security tooling S
-We summarize our interview participants’ demographic information in Table 2.
-Our interview sampling approach yielded 13 experienced security practitioners who indicated that they are responsible for initiating or implementing their organization’s security controls and (or) have compliance-related responsibilities.
-This involvement gives them the relevant context to assess their organization’s strategies for adopting software signing tools.
-Interviews: We conducted our interviews over Zoom.
-Each interview lasted a median of 50 minutes.
-The lead author conducted these interviews.
-We offered each subject a \$100 gift card as an incentive, in recognition of their experience level.
-4.3 Data Analysis
+This initial invitation yielded 6 participants.
+We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (5) and authors’ contacts (7) to recruit an additional 12 participants.
+We interviewed 18 subjects, but only 17 participants could share enough details for analysis.
+Participant Demographics: Our subjects were experienced security practitioners responsible for initiating or implementing their organization’s security controls or compliance.
+This gives them the relevant context to assess their organization’s strategies for adopting software signing tools.
+Subjects came from 13 distinct organizations, all companies.
+Relevant demographics are in Table II.
+Thirteen participants reported their organizations use Sigstore.
+Two use internally developed tools and two rely on Notary v1 and PGP signing.
+Those four participants enrich our findings, as their experiences shed light on usability factors that hinder adoption.
+TABLE II: Demographics.
+For Role, “Senior management” are senior managers/directors/executives; “Technical leader” are senior/lead/partner/principal engineers; and “Engineer” and “Manager” are junior staff.
+ID Role Experience Software Type
+P1 Research leader 5 years Internal POC software
+P2 Senior mgmt. 15 years SAAS security tool
+P3 Senior mgmt. 13 years SAAS security tool
+P4 Technical leader 20 years Open-source tooling
+P5 Engineer 2 years Internal security tooling
+P6 Technical leader 27 years Internal security tooling
+P7 Manager 6 years Security tooling
+P8 Technical leader 8 years Internal security tooling
+P9 Engineer 2.5 years SAAS security
+P10 Engineer 13 years SAAS security
+P11 Technical leader 16 years Firmware
+P12 Technical leader 4 years Consultancy
+S13 Senior mgmt. 16 years Internal security tool
+P14 Research leader 13 years POC security software
+P15 Senior mgmt. 15 years Internal security tooling
+P16 Senior mgmt. 15 years SAAS
+P17 Manager 11 years Security tooling
+[…]
+IV-E Data Analysis
 We transcribed our interview recordings using www.rev.com’s human transcription service and anonymized all potentially identifiable information.
-To analyze our interviews, we employed a thematic analysis approach [9] alongside our modified usability framework (section 3, Figure 1) within an interpretivist paradigm [4].
-An interpretivist approach allows each participant’s response to be treated as a unique case, considering its specific context and surrounding variables.
-This is particularly valuable in our study, as participants belong to different teams with diverse implementation contexts and varying uses of Sigstore.
-Our analysis began with a data familiarization stage [80] and proceeded through three phases, carried out by two analysts  [^2] One of the analysts is the lead author..
+After transcription and anonymization, our analysis began with a data familiarization stage [77] and proceeded through four phases, carried out by two analysts.
+[^2] One of the analysts is the lead author.
 In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [78].
@@
 We describe these steps in detail next.
-Memoing & Codebook Creation: First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [56] and Campbell et al.
+We remind the reader that the transcript memoing and codebook was developed across the full interview transcripts, including Topics B and C.
+This manuscript focuses on the data and codes pertinent to Topic D.
+IV-E1 Memoing & Codebook Creation
+First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [79] and Campbell et al.
 [80].
@@
 [^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [79].
-We randomly selected approximately 33% (six transcripts) to develop our codebook.
-This slightly exceeded O’Connor & Joffe’s recommended percentage to allow the analysts a better familiarization with the nature of our collected data, and analysis format.
+We randomly chose 6 transcripts (33% of data) to develop our initial codebook.
 After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos.
@@
 Specifically, we randomly selected an unanalyzed transcript, which we coded independently.
-Following Feng et al.
-’s [20] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook.
-We obtained an 89% agreement score and resolved all disagreements through discussions and code refinements.
-This high agreement indicates the stability of our coding scheme, allowing us to reliably singly code the remaining transcripts [12, 56].
-Codebook Revision: Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round.
+Following Feng et al.’s [82] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook.
+We obtained an 89% agreement score, suggesting a stable coding scheme [80, 79].
+We resolved disagreements through discussion.
+IV-E2 Codebook Revision
+Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round.
 In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst.
@@
 By mutual agreement between both analysts, these were categorized as minor codes.
-To ensure our sample size and analysis was adequate, we calculated code saturation for each category in our codebook.
-Following the recommendations and methods of Guest et al.
-[26] and Maxam & Davis [51], we measured the cumulative number of new codes introduced in each interview.
+To assess the adequacy of our sample size and analysis, we calculated code saturation for each category in our codebook.
+Following the recommendations of Guest et al.
+[83], we measured the cumulative number of new codes introduced in each interview.
+We saw saturation after interview 12, with no new codes after that.
 The results are presented in Figure 7.
-As shown, saturation was achieved after interview 11 , with no new codes emerging beyond this point.
-Each interview had a median of 15 unique codes (for topic D alone), as illustrated in Figure 6.
-Thematic Analysis & Usability Framework Analysis: Next, following the method proposed by Braun & Clarke [9], we derived initial themes from our code categories.
-We focused on identifying themes related to the D part of our protocol, as highlighted earlier in Table 1.
-We then conducted another round of thematic analysis, incorporating the initial themes alongside our formative usability framework (section 3).
-In this phase, we mapped our initial themes onto the factors defined within our framework.
-Survey Attempt: In addition to our interviews, we designed and conducted a survey to gather insights from members of the Sigstore community who might not have had the time for an interview.
-However, our survey received only a few responses, highlighting the challenges of conducting surveys within highly specialized communities [78].
-We elaborate on our survey results and details in Appendix D.
-4.4 Study Limitation & Threats to Validity
-To discuss the limitations of our work, we follow Verdecchia et al.’s [87] recommendation to reflect on the greatest threats to our research.
-In addition, we discuss how our experiences and perspectives may have influenced our research.
-Construct Validity: Cresswell’s usability framework was developed for health information technology. We adapted it to software signing tooling, but may have introduced flaws.
-Internal Validity: Threats to internal validity reduce the reliability of conclusions of cause-and-effect relationships.
-An additional internal threat exists in the potential subjectivity of the development of our codebook.
-To mitigate this subjectivity, we utilized multiple raters and measured their agreement.
-External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore.
-We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature.
-A secondary threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects.
-A final threat to the validity of our research is that it reflects the state of Sigstore at only one point in time.
-Our interviews were conducted between November 1, 2023, and February 9, 2024 and reflect the experiences of Sigstore users at that time.
-Statement of Positionality: We acknowledge that one of our paper’s authors is a Sigstore tool contributor.
-This author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
-5 Results
-5.1 RQ1: Problems & Strengths of Sigstore
-Table 3: Summary of Practitioner-Reported Advantages of Using Sigstore
+IV-E3 Thematic Analysis & Usability Framework Analysis
+Each interview had a median of 15 codes for topic D.
+Next, following the method proposed by Braun & Clarke [84], we derived initial themes from our code categories.
+We focused on identifying themes related to Topic D of our protocol (Table I).
+We then conducted another round of analysis using a formative usability framework. We selected Cresswell’s framework (Figure 1) based on the nature of our data.
+V Results
+First, we present a Sankey diagram illustrating the changes in signing tool usage among our participants’ teams (Figure 4).
+Many participants (6) transitioned from using no tools to adopting Sigstore.
+Others transitioned from existing tools to Sigstore: 3 from GPG, 2 from Notary, 1 from Skopeo, and 2 from proprietary or internal tools.
+The final 4 subjects did not transition to Sigstore.
+Figure 4: Practitioners’ Changes in Software Signing Tools.
+To understand why they did (and did not) change tools, we structure the results by RQ using Cresswell’s framework.
+V-A RQ1: Strengths & Weaknesses of Sigstore
+The first aim of our research work was to ascertain from the the experiences of our subjects the strengths and weaknesses they have experienced while using Sigstore.
+V-A1 Sigstore’s Strength
+We present a summarized list of practitioner-reported advantages of using Sigstore in Table III.
+First, we extracted the benefits our participants experienced while using Sigstore from their responses.
+These reflections capture their firsthand experiences with Sigstore.
+TABLE III: Summary of Practitioner-Reported Advantages of Using Sigstore. Practitioners cited Sigstore’s technological and macroenvironmental factors as advantages.
 Topics & Associated Examples Subjects
@@
 Ease of Use 8
-Signing Workflow & Verification P1, P2, P7, P8, P10-13
+Signing Workflow & Verification P1, P2, P7, P9, P14-17
 Setting up with automated CI/CD actions P9
 No key distribution problems P1
-Use of Short-lived Keys & Certificate 3 –P2, P3, P11
+Use of Short-lived Keys & Certificate 3 –P2, P3, P15
 Signer ID Management 4
@@
 Compatibility with Several New Technologies 4
-Integrability with SLSA build P12
-Integrability with several container registries/technologies P9, P10
-Integrability with several cloud-native applications P11
+Integrability with SLSA build P16
+Integrability with several container registries/technologies P12, P14
+Integrability with several cloud-native applications P15
 Precence of a Transparency Log 3
-Transparency logs increase security P5, P10
-Evaluation of signing adoption using logs P8
+Transparency logs increase security P5, P14
+Evaluation of signing adoption using logs P9
 Bundling Signatures With Provenance Attestations 2 – P3, P4
@@
 Macroenvironmental Factors
-Free/Open-Source 2 – P7, P13
-The first aim of our research work was to ascertain from the the experiences of our subjects the problem and strengths they have experienced while using Sigstore.
-We categorized each identified theme based on our usability framework factors: Technological Factors – situations where observed conditions were linked to the behavior of Sigstore technology itself, Human/Social Factors – participants’ observations related to gaps in support resources, previous experiences (with other tools) Organizational Factors – aspects tied to users’ organizations or their organizational capacities, and Macroenvironmental Factors – external influences affecting adoption and use.
-5.1.1 Sigstore’s Strength
-We present a summarized list of practitioner-reported advantages of using Sigstore in Table 3.
-First, we extracted the benefits our participants experienced while using Sigstore from their responses.
-These reflections capture their firsthand experiences with Sigstore.
+Free/Open-Source 2 – P7, P17
 We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service.
-When mapped to our modified usability framework, these factors fell under two categories: Technological and Macroenvironmental factors from our framework.
+When mapped to the Cresswell framework, these factors fell under two categories: Technological and Macroenvironmental.
 Technological Factors
-Ease of use: Ease of Use was a predominant reason why many participants chose to use Sigstore, making it the most frequently cited advantage over other Software signing tools.
-Participants commonly highlighted the simplicity of following the Sigstore workflow for signing and verifying signatures.
+Ease of use: This was the most frequently discussed advantage of Sigstore over other Software signing tools.
+Participants commonly highlighted the simplicity of following the Sigstore workflow—particularly how key management is fully automated reduce configuration overhead.
 P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
-Identity management for Signers: Sigstore’s integration of OIDC identity management was a key strength reported by our participants.
-In traditional software signing, signer identity is often an ad-hoc feature that may or may not be required by the signing authority.
-In Sigstore, however, this feature—implemented via OIDC—is integrated with the Fulcio certificate authority as an identity management step before creating a signature.
+Identity management for Signers: Sigstore’s integration of OIDC identity management was a strength per our participants.
+In traditional software signing, signer identity is often an ad-hoc feature and may be optional for the signing authority.
+In Sigstore, however, this behavior is integrated with the Fulcio certificate authority as an identity management step before creating a signature.
 Participants greatly valued the elimination of the key generation and distribution phase, as it reduced the time spent distributing keys to key servers and searching for keys, streamlining the signing process.
-P12 (SAAS, SSC security tool) Highlights this point, “The other strength that it has is I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key … that’s super useful because I could say, "Oh, this is signed with Mike’s public GitHub identity." So unless my GitHub identity has been compromised, that’s much better.”
+This aligns with the intuition that identity management is easier than key management for developers [11].
+P16 (SAAS, SSC security tool) highlights this point: “I can associate my identity with an OIDC identity as opposed to generat[ing] a key and keep[ing] track of that key…I could say, ‘Oh, this is signed with [Jane]’s public GitHub identity’…that’s much better”
 Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage.
@@
 In most cases, the resources required to set up connections to these technologies were minimal or nonexistent.
-P9 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
+P12 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
 Use of Short-lived Keys/Certificates: Sigstore’s enforcement of short-lived keys and certificates as a security feature was recognized by our participants as a major strength of the tooling.
@@
 Because Sigstore uses short-lived keys that solves a lot of those issues around key storage and key distribution…”
-Presence of a Transparency Log: The inclusion of a transparency log in the Sigstore ecosystem, serving as a tamper-resistant secondary record of changes to an artifact’s metadata, was seen as highly beneficial by participants.
-This feature provided an additional layer of trust, allowing users to verify signed artifacts with greater confidence.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
-Other technological factors noted by participants are the reliability of Sigstore’s public instance was a valuable aspect of Sigstore.
+Presence of a Transparency Log: The ability to audit signing actions was crucial for adopters.
+Participants valued the transparency log in the Sigstore ecosystem because it provides a tamper-resistant record of all changes to an artifact’s metadata—functionality not offered by other signing systems.
+This auditability gave users greater confidence in the integrity of signed artifacts.
+{quoting} P9 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
+Other technological factors noted by participants are the reliability of Sigstore’s public instance.
 Macroenvironmental Factors
-Free/Open-Source:
-Sigstore’s open-source nature was also highlighted by participants as a key advantage.
-They noted that the no-cost solution reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
-5.1.2 Sigstore’s Weaknessses
-We summarize the difficulties reported by participants in using Sigstore in  Table 4.
-Most of the weaknesses reported often belong to multi-factors.
-e.g., In the problems associated with Enterprise applicabiliirty of Sigstore, Some participants reported this as a consequence of regulations place on organizations which are both macroenvironmental (regulations) and Organizational.
-Table 4: Reported Difficulties Encountered by Participants in Using Sigstore.
+Free/Open-Source: Sigstore’s open-source nature was also highlighted by participants as a key advantage.
+They noted that this reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments.
+{quoting} P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
+V-A2 Sigstore’s Weaknessses
+We summarize the difficulties reported by participants in using Sigstore in Table IV. Most of these weaknesses mapped to multiple usability factors.
+TABLE IV: Reported Difficulties Encountered by Participants in Using Sigstore.
 We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.
 Topics & Associated Examples Subjects
-Enterprise Adoption Limitations 7
-Rate Limiting Problems – T P3, P7, P10, P11
-Lack of dedicated Support & Maintenance – T & P P11, P2
+Enterprise Adoption Limitations 7 subjects
+Rate Limiting Problems – T P3, P7, P14, P15
+Lack of dedicated Support & Maintenance – T & P P15, P2
 Not Suited for Regulated Organizations – M & O P3
 Latency Concerns – T P6
-Transparency Log Issues 6
-Not Suitable for private Setup – T P2, P3, P6, P10, P13
-Use in Air Gap Conditions – T P2, P3, P8
+Transparency Log Issues 6 subjects
+Not Suitable for private Setup – T P2, P3, P6, P14, P17
+Use in Air Gap Conditions – T P2, P3, P9
 Efforts to Monitor Logs – P P2
-Private Sigstore Instance Setup 5
-Documentation – P P8, P9, P6
-Limited Community Support – M P11
+Private Sigstore Instance Setup 5 subjects
+Documentation – P & T P9, P12, P6
+Limited Community Support – M P15
 Infrastructure Requirements & Maintenance Costs – T P5, P6
-Other Documentations and Usage Information IssuesP 3 — P1, P10, P13
-Integration to Other Systems 3
+Other Documentations and Usage Information IssuesP & T 3 — P1, P14, P17
+Integration to Other Systems 3 subjects
 Attestation Storage – T P1
-Gitlab & Jenkins – T P8
-Other Unsupported technologies – T P12
+Gitlab & Jenkins – T P9
+Other Unsupported technologies – T P16
 Offline CapabilitiesT 2 – P3, P4
@@
 Software LibrariesT 1 – P7
+Macroenvironmental & Organizational
 Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications.
@@
 Additionally, participants highlighted the lack of dedicated support and maintenance teams due to Sigstore ’s open-source nature, as well as its unsuitability for organizations in regulated sectors for the same reason.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P8 (SAAS, SSC security): “I think the major challenge was that we were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise.
-We would constantly be getting rate limited by the service.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable.
+{quoting} P7 (security tools, cloud security): “We were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise [rate limit].” {quoting} P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable.
 And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
+Technological
 Transparency Log Issues: While the transparency log bundled with Sigstore was highlighted as a major advantage by participants, it was also a significant concern.
 Their concerns stemmed from the public nature of the log, where sensitive artifact metadata from company assets could be exposed in a publicly accessible record.
-Another common issue raised was the log’s usability in air-gapped environments—systems that are physically isolated from unsecured networks for security reasons.
+Another common issue raised was the log’s usability in air-gapped [85] environments.
 Additionally, participants noted the manual effort required to continuously monitor and manage the log.
 Most of these participants were in the process of experimenting with setting up the transparency log at the time of the interviews.
-Transparency log issues were mostly technological.
 {quoting}
-[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P6 (security tooling, digital technology): “We’ve got some teams piloting Rekor, for instance, for different traceability usages and so on…I’m all for the transparency log, but we just have to be careful who it’s transparent to at what point in time.”
-Setting up Private Sigstore Instance : Sigstore’s customizability was a feature frequently utilized and discussed by participants; however, it was also associated with certain challenges.
+P6 (security tooling, digital technology): “We’ve got some teams piloting Rekor, for instance, for different traceability usages…I’m all for the transparency log, but we…have to be careful who it’s transparent to at what [time].”
+Setting up Private Sigstore Instance: Sigstore’s customizability was a feature frequently utilized by participants.
+However, it was also associated with certain challenges.
 The primary issues identified included limited documentation on setup, scarce community usage information, and the infrastructure requirements for private deployments.
-Issues here were primarily related to participants’ need for support resources, which where macroenvironmental and social in nature.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P11 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to help people deploy Fulcio, like on prem, or in their own infrastructure.”
+Issues here were primarily related to participants’ need for support resources, which were macroenvironmental and social in nature.
+{quoting} P15 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to…deploy Fulcio, like on prem, or in their own infrastructure.”
 Other Documentation Issues: Overall, the documentation for Sigstore, along with practical usage information and support, was found to be insufficient.
 Participants noted that updates to the documentation lagged behind changes in the Sigstore product and that there was a lack of clarity regarding the use of different Sigstore components.
-P13 (security tooling, internet service) highlights this, “ I would say the documentation lacks a bit.
+P17 (security tooling, internet service) highlights this, “I would say the documentation lacks a bit.
 Just I think if you look at the documentation, it was updated quite a long time ago, like the ReadMe file.
@@
 Commonly mentioned integrations included other CI/CD platforms such as Jenkins and GitLab, as well as attestation storage databases.
-These limitations were often attributed to Sigstore being an emerging technology, as highlighted by P12 (SAAS, SSC security tool), “The weakness is not everything supports it.
-There’s a lot of weaknesses just around not just it’s being a newer technology”
+These limitations were often attributed to Sigstore being an emerging technology, as highlighted by P16 (SAAS, SSC security tool), “The weakness is not everything supports it.
+There’s a lot of weaknesses around…its being a newer technology”
 Fulcio Issues: Some Sigstore issues were also tied to its certificate authority’s (Fulcio) time stamping capabilities, and implementation of OIDC for keyless signing.
 Fulcio acts as an intermediary using OIDC identities to bind to a short-lived certificate, this was criticized by P3.
-P3 also commented on the initial lack of time stamping abilities for the initial versions of Fulcio.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “I think we do have some concerns about the way Fulcio operates as its own certificate authority.
+P3 also commented on the lack of timestamping support in early versions of Fulcio.
+{quoting} P3 (SAAS, SSC security tool): “We do have some concerns about the way Fulcio operates as its own certificate authority.
 So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider.
 And then I think the other big thing that we did, because Sigstore originally didn’t support it, was time-stamping.
-Because we’re using Fulcio, we want to ensure that the signature was created when the certificate was valid.
-And so we use external time-stamping authority in order to be able to do that.”
+We want to ensure that the signature was created when the certificate was valid.”
 Offline Capabilities: Another reported limitation of Sigstore is the absence of offline verification capabilities using offline keys.
@@
 Software Libraries: The programming libraries of Sigstore were also mentioned as a challenge, particularly in cases where users wanted to integrate Sigstore’s signing capabilities directly into an application.
-5.1.3 Effect of Identified Problems on Sigstore components used by Subjects
-Table 5: Sigstore Functionalities Used by Subjects
-Tool # of Subjects
-Cosign (Signing & Verification CLI) 9
-Certificate Authority (Fulcio) 9
-Keyless Signing & Identity Manager (OIDC) 9
-Transparency Log (Rekor) 8
-Gitsign (Commit Signing) 4
-Component Customization & Local Sigstore Deployment 2
-We assessed the effect of participants’ reported problems with Sigstore on their usage of each component. Table 5 summarizes our subjects’ reported usage of Sigstore components and functionalities.
-Key components like Cosign, Fulcio, and keyless signing saw high adoption, while local Sigstore deployment had low usage due to numerous reported issues.
-Several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table 4, section 5.1.2).
+V-A3 Effect of Identified Problems on Sigstore components used by Subjects
+We assessed the effect of participants’ reported problems with Sigstore on their usage of each component.
+By component, nine subjects use Cosign, nine use the Fulcio certificate authority, and nine use the OIDC keyless signing and identity manager.
+Eight use the Rekor transparency log.
+Only four use Gitsign and only two make use of customized components and local Sigstore deployments.
+Among these, our data shed the most light on local deployments and the Rekor log.
+For local deployment, several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table IV, section V-A2).
 The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues.
@@
 Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
-Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other crucial elements was notably absent.
+Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other elements was notably absent.
 In particular, there was very little mention of log witnessing [86], or monitoring [87] for either Fulcio or Rekor.
@@
 We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
-5.2 RQ2: How and Why Practitioners Change Software Signing Tools
-Table 6: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
-Topics & Associated Examples Subjects
-Human/Social Factors
-Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P8, P10
-GPG Issues 6
-Low adoption rates P1, P10
-Key management issues & Other usability concerns P1, P6, P9, P12, P13
-Steep learning curve P9, P12
-Compatibility with newer technology P12
-Notary Issues 2
-Non-demand from customers P9
-Compatibility with other tools P9
-Lack of regular updates P9
-Key & Identity Management P5
-Proprietary Tool Issues 1
-Difficult to setup P11
-Technological Factors
-Available Sigstore Functionalities 3 – P1, P5, P10
-A transparency log, etc P5, P10
-Compatibility to other Tools P1
-Macroenvironmental Factors
-Regulation & Standards 4 – P5, P6, P11, P13
-Large User Community 1 – P8
-Inherent Trust of Creators 1
-Trust of CNCF products P3
-Figure 3: Practitioners’ Evolution of Software Signing Tools.
-Most reported their organizations moved from no former tools, a form of GPG signing or notary to Sigstore.
-Most participants depicted with the flow from none to a sigstore implementation report that; they either initially used no tools, or have always used Sigstore.
-We also tackle the usability of software signing tool especially as it relates to decisions that prompts the switch of practitioners in using different tools.
-In the context of our study, we asked participants what factors prior to adopting Sigstore influence practitioner’s to switch to Sigstore.
-The factors here are mostly Macroenvironmental (relating to regulations, user communities), Human (Previous experiences), and Technological (available capabilities of sigstore)
-First, we present a Sankey flow diagram ( Figure 3) illustrating the evolution of signing tool usage among our participants’ teams up to the present.
-A significant portion of participants (6), as depicted transitioned from using no tools to adopting Sigstore —either they had always used Sigstore since the inception of their organization or continued using their initial signing tools alongside Sigstore.
-The most notable transition was from GPG-based signing implementations (3) to Sigstore, followed by a shift from Notary (2).
-Other reported transitions included Skopeo (1), unspecified internal tools (1), and a proprietary tool (1).
-The reasons most of our participants chose to switch to Sigstore stemmed from a combination of two key factors: issues related to a subpar user experience with previous tools and advantages associated with Sigstore.
-We summarize these factors in Table 6.
+V-B RQ2: How & Why Practitioners Change Tools
+V-B1 Drivers of Adoption Among Current Sigstore Users
+For participants who adopted Sigstore, we asked which factors—prior to adoption—influenced their decision (or their organization’s decision) to switch.
+Their responses fell into three main categories: macroenvironmental factors (e.g., regulations and user communities), human/social factors (e.g., prior experience with signing tools), and technological factors (e.g., Sigstore’s unique capabilities).
+Most of our participants chose to switch to Sigstore due to poor experiences with previous tools and due to perceived advantages associated with Sigstore. We summarize these factors in Table X.
 Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore.
@@
 Additionally, some participants mentioned that key members of their organizations were directly involved in the initial development of the Sigstore project.
-P10 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community.
+P14 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community.
 There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
-Available Sigstore Functionalities: Sigstore’s advertised functionalities also influenced its adoption.
-While several participants were drawn to the transparency log (as highlighted in section 5.1.1), its integrability with existing technologies they were using at the time also played a role in their decision.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “… Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, as well as they also maintain a transparency log as well …” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P1 (POC software, digital technology): “Factors I considered were, …, the existence of tools to use it.”
+Available Sigstore Functionalities: Sigstore’s functionalities also influenced its adoption.
+While several participants were drawn to the transparency log (as highlighted in section V-A1).
+{quoting} P5 (internal security tool, cloud security): “Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, [and] they also maintain a transparency log.”
+Integration with existing technologies also affected their decision. As P1 (POC software, digital technology) said, “Factors I considered were, …, the existence of tools to use it.”
 Regulations and Standards: The impact of regulations and standards on adopting Sigstore was significant among our participants.
@@
 {quoting}
-[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P5 (internal security tool, cloud security): “…standards actually really helped to convince everyone that we should be doing this … But once a standard is put in place, once the requirements are set, then it just sped up the process of adopting these new technologies like Sigstore …”
-User Community & Trust Of Sigstore Creators: The large user community of Sigstore was also a key reason for its adoption among participants.
-This aligns with feedback from some participants (P1, P10) who cited GPG’s smaller user base as a drawback.
-Additionally, some participants expressed inherent trust in the CNCF (Cloud Native Computing Foundation) security tooling community, which they had grown to rely on over time.
-{quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] P3 (SAAS, SSC security tool): “We do defer often at the higher level projects to different foundations.
-So we would be more likely to trust something from the CNCF, the Cloud Native Computing Foundation or Sigstore rather than small little independent developer projects or things like that.”
+P5 (internal security tool, cloud security): “…standards actually really helped to convince everyone that we should be doing this … But once a standard is put in place, once the requirements are set, then it just sped up the process.”
+User Community & Trust Of Sigstore Creators: The large user community of Sigstore also motivated its adoption among participants.
+Conversely, participants P1 and P14 cited GPG’s smaller user base as a drawback.
+Additionally, some participants expressed trust in industry consortia such as CNCF (Cloud Native Computing Foundation) {quoting} P3 (SAAS, SSC security tool): “We do defer often at the higher-level projects to different foundations.
+So we would be more likely to trust something from the CNCF…[than] independent developer projects.”
 GPG & Notary Issues: Issues associated with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore.
@@
 Our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies.
-Notary-related concerns primarily revolved around customer demands, compatibility with other technologies, and infrequent updates.
+Notary-related issues primarily concerned customer demands, compatibility with other technologies, and infrequent updates.
 Key and identity management challenges were common to both GPG and Notary.
 Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
-6 Discussion
-In this section, we highlight two reflections from our data.
-The role of community in software signing tool adoption: A substantial number of participants reported that they adopted Sigstore because of their involvement as contributors to the Sigstore tooling community (section 5.2).
-This reflects a key macroenvironmental factor from our usability frameworks that influenced users’ decisions to adopt these tools.
-Similarly, several reported challenges were linked to the lack of community support in setting up private Sigstore instances and accessing usage information (section 5.1.2).
-These findings highlight the critical role of community, as emphasized in prior research [72].
-New adopters often depend on communities they trust, contribute to, or receive support from, which significantly shapes their decision to adopt such tools.
-Toolmakers should prioritize fostering and maintaining a strong community around their tools, as this can greatly influence adoption and success.
-Privacy Concerns: Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern.
+V-B2 Barriers and Motivators Spurring Consideration of Sigstore Among Non-Sigstore Users
+TABLE V: Non Sigstore Users: Reasons for Not Adopting Sigstore.
+This table summarizes the reasons non-users gave for not adopting Sigstore, grouped by their current signing tool and mapped to our usability framework theme.
+Tool Subjects Reason (Topic)
+Internal tool P8, P11 (1) Third-party security risk from tool T – P11, P8
+(2) Organizational Use cases O – S8
+(3) Centralizing Organization’s code signing O – P8
+(4) Organization’s Open Source Policy O – P8
+Notary-V1 P10 (1) Tool owned by Organization O – S10
+PGP P13 (1) Sigstore Privacy concerns T – P13
+(2) Trust in Sigstore’s Maintainers M – P13
+In this section we highlight reasons mentioned by practitioners that have stopped them (or their organization) from switching to Next-gen tools like sigstore.
+We also report issues experienced by these non sigstore users that have prompted a consideration (or in some cases partial adoption of these next-gen tools).
+Barriers to Considering Sigstore: As highlighted in Table V, most reasons for non-adoption were organizational.
+{quoting} P8 (internal security tool & cloud APIs, internet services): “When it comes to open source projects and open source stuff, our Organization is very skeptical.
+Policy and everything drives very much teams away from that sort of stuff…There’s also the ongoing risk of having a product that’s built outside continually shipping code and being integrated into this product, where it deals with sensitive material.
+So we do evaluate these things, but the scrutiny against code written outside of the company is extreme, and with good reason given the context of all the supply chain stuff.” Other notable issues arose from technical concerns, including privacy risks associated with Sigstore’s public instance and the inherent risk of relying on a third-party tool maintained by open-source contributors.
+Factors Motivating Consideration for sigstore: When practitioners were asked about issues with their current tools and potential fixes, most pointed to existing tooling problems as the primary driver for considering alternatives—such as incorporating next-gen features into internal tools or partially adopting new solutions.
+Although some subjects reported drawing design inspiration from next-generation workflows, none specifically mentioned Sigstore.
+P11 (firmware & testing software, social technology) states, “.
+I’m sure at some point in time it was influenced by an open source project before they adopted it and made the right tweaks internally.” Lack of transparency and integration of these tools to other platforms were also common.
+We summarize these factors in Table XI.
+VI Discussion
+VI-A Lessons for Next-Generation Tools
+Next-generation signing tools offer unparalleled flexibility by integrating with external platforms, and services.
+However, our participants consistently highlighted integration challenges as a major obstacle.
+VI-A1 Automated Cross-Platform Integrations
+To fully leverage the flexibility of next-gen tools, practitioners need integration modules for popular CI/CD systems, registries, and orchestration platforms.
+Missing integration and sparse documentation forces users to implement custom connectors, increasing the risk of misconfiguration and potential security vulnerabilities that undermine the very guarantees signing is meant to provide.
+VI-A2 Need for Inter-Platform Standards
+Standardization of information exchange is a standard problem for new classes of software infrastructure e.g., , SBOM [88].
+These problems often harm integration capabilities.
+Next-generation signing tools would gain from common exchange standards—for example, uniform schemas for transparency logs or identity assertions.
+Projects like Diverfiy [89], which harmonize identity provisioning across multiple OIDC providers, are a promising start.
+Similar efforts are needed to streamline data flows between signing components, transparency services, and external tooling ecosystems.
+VI-A3 Verification Workflow
+A key security property in a software supply chain is the transparency of actors and artifacts, which requires verifying signer identities [4].
+Previous studies have highlighted challenges in signature verification across both organizational and open-source contexts [39, 5].
+Next-generation tools like Sigstore integrate multiple verification methods, such as the Cosign CLI and the transparency log, but most participants still reported usability issues with the transparency log, and documentation indicates a lack of automation for other verification methods.
+Further automating the verification workflow would enhance the security benefits of software signing with this class of tools.
+VI-B Usability Patterns of Sigstore vs Non-Sigstore users
+Our results reveal different drives in adoption motivations and shared usability concerns between Sigstore users and non-users.
+Sigstore adopters cited contributions to the Sigstore project (Human/Social factor) and macroenvironmental factors—alongside negative experiences with legacy tools—as their primary drivers.
+In contrast, non-users emphasized organizational considerations when choosing their signing solution.
+Despite these differences, both groups reported similar technological usability issues, particularly around integration with other systems and infrastructure transparency.
+Notably, non-users frequently highlighted signer identification challenges, whereas Sigstore users did not—underscoring the clear usability advantage of next-generation tooling.
+VI-C Privacy Concerns
+Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern.
+Privacy concerns were also major issues influencing users not to consider sigstore as suitable for enterprise.
 This issue also affects the use of OIDC for keyless signing.
@@
 While works like Speranza[71] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
-7 Conclusion
-Software signing allows software provenance to be guaranteed through cryptographic primitives.
-This property would address many software supply chain attacks.
-However, for decades, the widespread adoption of software signing has eluded the software industry.
-In this study, we examined a recent change in software signing practices by focusing on Sigstore, the leading next-generation software signing platform.
-We interviewed industry experts to understand Sigstore from a usability perspective.
-We identified the factors they weigh and the challenges they face.
-Our results shed light on usability issues for software signing in particular, and will inform more usable software tools writ broad.
-8 Acknowledgments
+VII Threats to Validity
+To discuss the limitations of our work, we follow Verdecchia et al.’s [90] recommendation to reflect on the greatest threats to our research.
+In addition, we discuss how our experiences and perspectives may have influenced our research (positionality).
+Construct Validity: Cresswell’s framework targets health IT, where “tooling” denotes integrated clinical systems, so applying it to software signing carries construct risks.
+We mitigated risks by inductively validating themes (from its core factors) against practitioner interviews, ensuring our adapted constructs reflect signing-tool usability.
+Internal Validity: Threats to internal validity reduce the reliability of conclusions of cause-and-effect relationships.
+An additional internal threat exists in the potential subjectivity of the development of our codebook.
+To mitigate this subjectivity, we utilized multiple raters and measured their agreement.
+External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore.
+We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature.
+A second threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects.
+We attempted a survey (details are in section -E) but received only a few responses, echoing the challenges of surveying specialized communities [91].
+A third threat is our study’s temporal scope.
+Interviews were conducted between November 1, 2023, and February 9, 2024, capturing Sigstore users’ experiences during that period.
+Since then, Sigstore has introduced some changes, but not the core workflows and components.
+This suggests our findings remain relevant.
+Statement of Positionality: The author team has expertise in software supply chain security and qualitative research methods, e.g., publishing top-tier papers in software engineering and cybersecurity venues.
+One of the authors is a Sigstore contributor.
+To avoid bias, this author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
+VIII Conclusion
+Software signing guarantees provenance via cryptographic primitives, reducing supply-chain attacks.
+Next-generation tools like Sigstore automate traditional challenges, yet their usability remains underexplored.
+We interviewed practitioners to examine Sigstore’s usability, identifying key adoption factors.
+Our findings show that Sigstore’s identity-based workflows improve usability, but integration complexities, privacy concerns, and organizational constraints hinder adoption.
+We recommend toolmakers provide official integration plugins and privacy controls for transparency logs.
+These insights inform automation improvements, promoting secure, usable next-generation signing in automated software engineering.
+IX Acknowledgments
 We thank the study participants for contributing their time to this work.
@@
 We also thank Hayden Blauzvern and all reviewers who provided feedback on this work.
-Appendix A Interview Protocol
-Table 1 gave a summary of the interview protocol.
-Here we describe the full protocol.
-We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style.
-Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
-We include Section D of the protocol for completeness, although it was not analyzed in this study.
-Table 7: Our full interview protocol.
-The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions.
-We only analyze parts A and D in this study
-Themes Sub-themes Questions
-A: Demographic A-1: What best describes your role in your team? (Security engineer, Infrastructure, software engineer, etc)
-A-2: What is your seniority level? How many years of experience?
-A-3: What is the team size?
-A-4: $\dagger$ What are the team’s major software products/artifacts?
-A-5: What type of Organization/company?
 B: Software Supply Chain
@@
 D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
-Appendix B Codebooks
-Table 8 describes our codebook, with code definition, and example quote tagged with that code.
-Table 8: Excerpts from the final codebook we used for our analysis.
+-C Codebooks
+Table VII describes our codebook, with code definition, and example quote tagged with that code.
+TABLE VII: Excerpts from the final codebook we used for our analysis.
 Code Definition Sample Quote
 Sigstore Criticisms Criticisms and weaknesses associated with implementing Sigstore for signing.
-S3: "We do have some concerns about the way Fulcio operates as its own certificate authority.
+P3: "We do have some concerns about the way Fulcio operates as its own certificate authority.
 So we’ve been looking at things like OpenPubkey as something that removes that intermediary and would allow you to do identity-based signing directly against the OIDC provider.
@@
 Sigstore Improvements Suggested improvements for Sigstore.
-S9: "I would say probably integrations within different platforms, so say you use GitLab or GitHub or Jenkins or something like that.
+P9: "I would say probably integrations within different platforms, so say you use GitLab or GitHub or Jenkins or something like that.
 Try to leverage some type of integration to make that, I guess, less friction for the developers to implement, something like that."
-Sigstore Strengths Strengths of Sigstore and reasons why subjects prefer using Sigstore. S7: "The convenience and the good user experience is probably what keeps us using Sigstore "
+Sigstore Strengths Strengths of Sigstore and reasons why subjects prefer using Sigstore. P7: "The convenience and the good user experience is probably what keeps us using Sigstore "
 Sigstore Alternatives Considered Alternatives of Sigstore that were or are being considered by the subject (for subjects whose primary signing implementation is Sigstore).
-S5: "We are also looking at other new alternatives that are coming up, like OpenPubkey.
+P5: "We are also looking at other new alternatives that are coming up, like OpenPubkey.
 Docker just released a new standard called OpenPubKey, which is trying to solve this issue where they will not require you to have infrastructures and stuff."
-Sigstore Components Used Components of Sigstore (Cosign, Fulcio, etc.) being used by the subject. S2: "Mostly fulcio is what’s integrated into our product."
+Sigstore Components Used Components of Sigstore (Cosign, Fulcio, etc.) being used by the subject. P2: "Mostly fulcio is what’s integrated into our product."
 Notary Criticism Criticisms and weaknesses Associated with Notary – promoting users to stop using them.
-"S5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
-Appendix C Traditional vs Next-Gen Software Signing Workflow
-We show a typical traditional package signing workflow in Figure 4 compared to a next-generation workflow in Figure 5.
-Figure 4: Typical workflow for software signing and verifying signatures.
-The component author packages (A) and signs (B) their software.
-The signed package (C) and public key (D) are published.
-To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
-Figure 5: Typical workflow for next-generation software signing.
-The component author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider.
-The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
-Traditional Signing: The signature creation process begins when the maintainer’s artifact is ready for submission.
-The signer generates a key pair that provides them with a private key for signing their software artifact and a public key for a verifier to validate the artifact’s signature.
-Once the artifact is signed, it must be submitted to a package registry (such as PyPi, npm, or Maven) or the intended party.
-During this phase, the signer is responsible for making their public key available and discoverable so future users can verify the signer’s identity.
-In the verification phase, the verifier retrieves the signer’s public key, uses it to decrypt the signature file, and checks for equivalency with the software artifact’s hashed digest.
-Next-generation Signing:: Keyless/Ephemeral software signing mitigates the security risk of private key compromise.
-Keyless software signing differs from traditional public-key methodology in the fact that the signer does not rely on a long-term private key to sign their messages.
-Instead of the signer being responsible for the secure maintenance of a private key, in keyless signing, the signer uses a short-lived key that only lasts long enough to sign their artifact [2].
-Appendix D Survey Attempt
+"P5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
+-D Code Saturation
+Figure 7: Saturation curve.
+Interviews are plotted in the order in which they were conducted.
+Each interview covered many topics in detail (orange line, blue line).
+However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
+We present a saturation analysis of our studies in  Figure 7.
+The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview.
+We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 12.
+Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4.
+This is also reflected in the orange trend line (unique number of codes).
+-E Survey Attempt
 In addition to our interviews, we developed a survey distributed to Sigstore users at the 2024 SigstoreCon conference and through post-conference communications with the Sigstore user community.
@@
 2/3 selected that they check Rekor logs.
-Appendix E Code Saturation
-Figure 6: Saturation curve.
-Interviews are plotted in the order in which they were conducted.
-Each interview covered many topics in detail (orange line, blue line).
-However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
-We present a saturation analysis of our studies in  Figure 6.
-The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview.
-We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 11.
-Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4.
-This is also reflected in the orange trendline (unique number of codes).
-Appendix F Code Evolution Diagram
-Figure 7 describes the evolution of our codebook over six rounds of revisions.
-Figure 7: Codebook Evolution: The codebook underwent six rounds of revisions, with changes highlighted by different colors: round 2 (pink), round 3 (blue), round 4 (green), round 5 (yellow), and round 6 (purple).
-The final codes from round 6 were used in this paper, and categorized thematically.
-We combined framework and thematic analyses to develop the codebook.
-First, we induced codes from each transcript and analyzed them thematically.
-During the framework analysis, we mapped codes to the four stages/blocks of the software supply chain factory model.
-Round 1: The codebook was defined based on emerging themes.
-Round 2: No changes.
-Round 3: Blue-highlighted codes (5) were added after memoing and coding the second set of transcripts.
-Round 4: Green-highlighted codes (2) were added.
-Round 5: No changes.
-Round 6: Analysts A1 and A2 determined some codes should be further divided.
-Purple-highlighted codes (4) were added and the codebook was finalized.
-We focused on the topics of relating to sigstore(signing tools) including, Notary, Vault, Spiffe, GPG, and others.
+-F Sigstore Components Most Commonly Used by Our Sample Population
+TABLE VIII: Sigstore Functionalities Used by Subjects.
+Tool # of Subjects
+Cosign (Signing & Verification CLI) 9
+Certificate Authority (Fulcio) 9
+Keyless Signing & Identity Manager (OIDC) 9
+Transparency Log (Rekor) 8
+Gitsign (Commit Signing) 4
+Component Customization & Local Sigstore Deployment 2
+-G Expanded Demographics Table
+TABLE IX: Subject Demographics.
+For anonymity, we used generic job roles, not specific titles.
+“Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff.
+We highlight the type of software subject’s immediate team produces.
+* refers to cases where subject’s team built other types of software in addition to the main stated.
+O refers to cases where subject belonged to multiple teams with other products than the one stated.
+ID Role Experience Software Type
+P1 Research leader 5 years Internal POC software
+P2 Senior mgmt. 15 years SAAS security tool*O
+P3 Senior mgmt. 13 years SAAS security tool*O
+P4 Technical leader 20 years Open-source tooling
+P5 Engineer 2 years Internal security tooling
+P6 Technical leader 27 years Internal security tooling*O
+P7 Manager 6 years Security tooling*O
+P8 Technical leader 8 years Internal security tooling*
+P9 Engineer 2.5 years SAAS security*
+P10 Engineer 13 years SAAS security*
+P11 Technical leader 16 years Firmware*
+P12 Technical leader 4 years Consultancy
+S13 Senior mgmt. 16 years Internal security tool
+P14 Research leader 13 years POC security Software*
+P15 Senior mgmt. 15 years Internal security tooling
+P16 Senior mgmt. 15 years SAAS
+P17 Manager 11 years Security tooling
+-H Factors Influencing Sigstore Adoption (Sigstore Users)
+TABLE X: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
+Practitioners’ decisions are driven primarily by human/social factors—such as community contribution, frustration with legacy tools, and usability issues—supplemented by Sigstore’s technological capabilities and reinforced by macroenvironmental pressures like regulations and trust in the CNCF ecosystem.
+Topics & Associated Examples Subjects
+Human/Social Factors
+Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P9, P14
+GPG Issues 6
+Low adoption rates P1, P14
+Key management issues & Other usability concerns P1, P6, P12, P16, P17
+Steep learning curve P12, P16
+Compatibility with newer technology P16
+Notary Issues 2
+Non-demand from customers P12
+Compatibility with other tools P12
+Lack of regular updates P12
+Key & Identity Management P5
+Proprietary Tool Issues 1
+Difficult to setup P15
+Technological Factors
+Available Sigstore Functionalities 3 – P1, P5, P14
+A transparency log, etc P5, P14
+Compatibility to other Tools P1
+Macroenvironmental Factors
+Regulation & Standards 4 – P5, P6, P15, P17
+Large User Community 1 – P9
+Inherent Trust of Creators 1
+Trust of CNCF products P3
+-I Factors Influencing Sigstore Consideration (Non-Sigstore Users).
+TABLE XI: Non Sigstore Users: Tooling Issues causing a consideration of Sigstore and other Next-gen tools.
+Tool Subjects Topic
+Internal tool P8, P11 Tooling Issues
+(1) Integration with other tools/platformsT – P8, P11
+(2) Non unified tool implementation across teams O – P8, P11
+(3) Signer ID Management T – P11
+(4) Usage info/documentation unclear T – P11
+(5) Verification not built in with tool O – P8
+(6) Non transparency in signing infrastructure T – P11
+(7) Security of Signing Infrastructure T – P11
+(8) Engineers don’t understand the current tool P – P11
+Considered (or Adopted) Fix
+(1) Organization building new tools with Next-gen features – P8, P11
+Notary-V1 P10 Tooling Issues
+(1) Key ManagementT
+(2) Lack of an inbuilt signing policy T
+(3) Engineers don’t understand the current tool H – P11
+(4) Integration with other tools/platformsT
+Considered (or Adopted) Fix
+(1) Organization has adopted Openpubkey for key management
+(1) Organization considering Sigstore, and Notary V2
+PGP P13 (1) Tooling Issues
+(1) Key ManagementT
+(2) Non transparency in signing infrastructure T
+(3) Security of Signing Infrastructure T
+Considered (or Adopted) Fix
+(1) Use of Key discovery and delivery systems like TUF [^4] The Update Framework[93]
+-J Additional Discussions & Implications of Results
+In addition to the implications discussed in section VI, we outline further implications of our findings below.
+-J1 Generalizability of Findings to Other Classes of Tools in Software Engineering
+Our findings reinforce several adoption factors noted in prior literature on open source tools—such as ease of use, compatibility with existing technologies, and integration flexibility—which were especially salient for next-generation signing tools like Sigstore.
+They also highlight additional considerations that may generalize to other classes of secure software engineering tools example the role played by macroenvironmental factors like the tooling community, and regulations.
+Furthermore, our results suggest adoption patterns not widely discussed in existing research.
+In particular, we observe the emergence of a migration pathway model, where teams incrementally adopt next-generation tools.
+Several non-Sigstore users reported experimenting with or partially integrating smaller solutions—such as OpenPubKey and TUF—to meet specific needs like key delivery and management.
+This suggests a stepwise transition process in tooling adoption that may be common across emerging tool ecosystems.
+This model can also be important to understand and further propose a streamlined organizational adoption pathway for software engineering tools.
+-J2 The Surprising Role of Community in Software Signing Tool Adoption
+Many participants cited their involvement in the Sigstore community as a key reason for adopting the tool (Table X).
+While this may partly reflect our sampling, half of the non-Sigstore users who adopted external tools also emphasized community affiliation as a motivator.
+Challenges like setting up private Sigstore instances or accessing usage knowledge were frequently linked to the absence of sufficient community support (section V-A2).
+Other community-related influences, such as the size of the user community and trust in the tool creators, also shaped adoption decisions.
+These findings echo prior work on the importance of community in developer tool adoption [94].
+Toolmakers should actively foster supportive, trustworthy communities around their tools, as these networks can be pivotal to both adoption and long-term sustainability.
+-K Legible Images
+We have added a more legible version of our Sigstore workflow imageFigure 3 here Figure 8.
+Figure 8: Sigstore Signing Workflow.
+The software author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider.
+The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.

### v5

@@
 Software signing is the most robust method for ensuring the integrity and authenticity of components in a software supply chain.
-However, traditional signing tools place key management and signer identification burdens on practitioners, leading to both security vulnerabilities and usability challenges.
-Next-generation signing tools such as Sigstore have automated some of these concerns, but little is known about their usability and adoption dynamics.
-This knowledge gap hampers the integration of signing into the software engineering process.
-To fill this gap, we conducted a usability study of Sigstore, a pioneering and widely adopted exemplar in this space.
-Through 18 interviews, we explored (1) the factors practitioners consider when selecting a signing tool, (2) the problems and advantages associated with the tooling choices of practitioners, and (3) practitioners’ signing-tool usage has evolved over time.
-Our findings illuminate the usability factors of next-generation signing tools and yield recommendations for toolmakers, including: (1) enhance integration flexibility through officially supported plugins and APIs, and (2) balance transparency with privacy by offering configurable logging options for enterprise use.
-I Introduction
-The reuse of software components in modern products creates complex supply chains and expands the attack surface for these systems [1, 2, 3].
-A core strategy for mitigating this risk is to establish provenance—verifying actor authenticity and artifact integrity—through software signing [4, 5].
-Software signing is the strongest guarantee of provenance [6] and is widely advocated by academia, industry consortia, and government regulations [7, 8, 9].
-High-profile supply-chain attacks—such as the SolarWinds breach—have further underscored the need for stronger, more automated signing infrastructure.
-In response, tools have evolved into two classes: traditional solutions (e.g., OpenPGP [10]), which delegate key management and signer identification to practitioners, and next-generation, identity-based tools (e.g., Sigstore [11]), which further automate the process via short-lived keys and external identity providers.
-Tooling has significantly advanced various technological fields [12, 13, 14, 15] and has often played a key role in the adoption of these technologies [16, 17].
-There is a continual need to evaluate the adoption practices and usability of such tools to inform practitioners, potential users, researchers, and toolmakers of their effectiveness in different contexts [18].
-Traditional software signing tools have been studied in this way, most notably in several “Why Johnny Can’t Encrypt” inspired studies[19, 20, 21].
-Although next-generation signing tools promise greater automation, there remains a lack of comprehensive usability and adoption analyses for these solutions.
-Such studies could assess how automation affects adoption suitability, elicit practitioners’ desired improvements, and guide future tool design and research.
-Our work offers the first examination of next-generation software signing tools’ usability as a critical factor influencing their adoption and practice.
-Building on and improving previous methodologies — which relied on novice users and summative usability assessments criticized for their lack of actionable feedback[18, 22] — we conducted interviews with 18 experienced security practitioners to study current software signing tools, focusing on Sigstore as a case study.
-We analyzed our data with a formative usability framework [23], focusing on two usability dimensions: practitioners’ experiences with Sigstore and its usability relative to other signing tools.
-From these interviews, we distilled three key themes: the evolution and motivations behind tool adoption, the limitations and challenges of currently adopted tools, and the specific implementation practices shaping real-world signing workflows.
-Our results highlight the usability factors practitioners consider when adopting next-generation software signing tools.
-This serves as formative feedback for tool designers and researchers, offering insights into the drivers of tool usage.
-We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools.
-Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of next-generation Software signing tools.
+Traditional signing tools burdened practitioners with key management and signer identification, creating both usability challenges and security risks.
+A new class of next-generation signing tools have automated many of these concerns, but little is known about their usability and its effect on adoption and effectiveness in practice.
+A usability evaluation can clarify the extent to which next-generation designs succeed and highlight priorities for improvement.
+To fill this gap, we conducted a usability study of Sigstore, a pioneering and widely adopted exemplar of next-generation signing.
+Through interviews with 17 industry experts, we examined (1) the problems and advantages associated with practitioners’ tooling choices, (2) how and why their signing-tool usage has evolved over time, and (3) the contexts that cause usability concerns.
+Our findings illuminate the usability factors of next-generation signing tools and yield recommendations for toolmakers, adopting organizations, and the research community.
+Notably, components of next-generation tooling exhibit different levels of maturity and readiness for adoption, and integration flexibility is a common pain point but potentially mitigable through plugins and APIs.
+Our results will help next-generation signing toolmakers further strengthen software supply chain security.
+1 Introduction
+The reuse of software components in modern software production creates complex supply chains that are susceptible to the unauthorized introduction of code [100, 84, 11, 48].
+To mitigate this risk, engineers need provenance—evidence of actor authenticity and artifact integrity [15].
+Software signing provides the strongest guarantee of provenance currently available [70] and has thus been widely advocated by academia, industry consortia, and government regulators [92, 22, 21, 19].
+Like other security practices, the effectiveness of software signing in practice depends on the availability and usability of tools that support it.
+Traditional signing tools such as OpenPGP are famously challenging to use [98], with several “Why Johnny Can’t Encrypt”–inspired studies consistently revealing usability challenges that limit adoption [99, 67, 66].
+In recent years, next-generation, identity-based systems (e.g., Sigstore [104]) have been introduced that automate much of the process using short-lived keys and external identity providers [26].
+However, despite growing adoption driven by recent supply chain attacks and efforts to lower usability barriers [104, 69, 55], we lack evidence about the usability of these next-generation tools and need to understand how that affects adoption in practice [51].
+The innovations in next-generation approaches to signing creates a pressing need to evaluate how usability concerns affect adoption and effectiveness.
+A usability evaluation can clarify the extent to which next-generation designs succeed, highlight how usability compares to traditional approaches, and surface priorities for improvement.
+To this end, we offer the first empirical study of the usability of next-generation software signing tools.
+Our work examines Sigstore—a pioneer next-generation signing tool that has seen rapid uptake across major open-source ecosystems[81, 36, 49], cloud-native projects, and corporations[96, 45, 76, 75, 77], involving a large number of engineers worldwide.
+We interviewed 17 experienced security practitioners to investigate the usability concerns that drive or discourage Sigstore’s adoption.
+We adopted an exploratory data analysis approach, because this study addresses a new class of tools that has not previously been analyzed.
+Our analysis centered on two aspects: practitioners’ direct experiences with Sigstore and their perceptions of its usability relative to other signing tools.
+Our study highlights the usability factors practitioners weigh when adopting next-generation software signing tools.
+We show that while next-generation signing tools ease traditional challenges, adoption is still constrained by integration hurdles, privacy concerns, and organizational factors.
+Second, our results provide adopting organizations with insight into the readiness levels of these tools, helping them manage adoption risks.
+Finally, we offer toolmakers and designers concrete directions for improving next-generation signing tools.
+Together, these findings provide formative feedback situated in real-world usage contexts.
 In summary, we make the following contributions:
@@
     We discuss how practitioners’ choices of software signing tools change over time and the factors influencing those decisions.
-Significance of Study: Our work targets Sigstore—a pioneer next-generation signing tool that has seen rapid uptake across major open-source ecosystems[24, 25, 26], cloud-native projects, and corporations[27, 28, 29, 30, 31], involving a large number of engineers worldwide.
-By unpacking the usability factors that drive or hinder Sigstore’s adoption, we not only inform the ongoing refinement of its automation workflows (e.g., certificate issuance and key management) but also provide a template for evaluating and improving similar Next-generation(identity-based) signing solutions.
-Also, since Sigstore exemplifies a broader class of tools that embed security into automated build and deployment processes, our findings can be potentially generalized to guide the design and implementation of other automated software-supply-chain technologies, ultimately advancing the usability—and thus the security—of automated software engineering practices.
-II Background & Related Works
-In this section, we review software signing technologies (section II-A) and usability assessments (section II-B).
-II-A Software Signing
+  3. 3.
+    We discuss how organizational deployment contexts influence the usability concerns raised by practitioners.
+Significance: Our work contributes to the security of software signing tools and, more broadly, to software supply chains.
+By unpacking the usability factors that promote and hinder Sigstore’s adoption, we inform the ongoing refinement of its workflows and provide a template for evaluating and improving similar next-generation (identity-based) signing solutions.
+Because Sigstore exemplifies other software signing tools, our findings can be generalized to guide the design and implementation of this class of software, ultimately advancing usability (and thus security) across the software supply chain.
+2 Background & Related Works
+In this section, we review software signing technologies (section˜2.1) and usability assessments (section˜2.2).
+2.1 Software Signing
 Software signing ensures the authorship and integrity of a software artifact by linking a maintainer’s cryptographic signature to their software artifact.
-The implementation of software signing is shaped by several factors: the software artifacts to be signed (e.g., driver signing[32]), the root-of-trust design (e.g., public key infrastructure/PKI [33]), and ecosystem policies (e.g., PyPi[34] and Maven[35]).
-II-A1 Software Signing Tools
+From literature, the implementation of software signing is shaped by several factors: the software artifacts to be signed (e.g., driver signing[24]), the root-of-trust design (e.g., public key infrastructure/PKI [105]), and ecosystem policies (e.g., PyPi[87] and Maven[80]).
+Software Signing Tools
 Software signing tools can be classified into two groups: traditional and next-generation (identity-based) [71].
@@
 Next-generation (identity-based) signing tools automate key management by issuing short-lived, ephemeral certificates tied to a signer’s identity via external OpenID Connect (OIDC) or OAuth providers.
-This “keyless signing” approach — exemplified by Sigstore [11], OpenPubKey [37], and SignServer [38] — eliminates long-term key handling and embeds transparency through append-only logs, improving usability and security.
-We discuss the architecture and workflow for next-generation signing tools in section IV-B using the case of Sigstore.
-II-A2 Effects of Signing Tools on Software Signing Adoption
-Previous studies recognize that better tooling influences software-signing adoption.
+This “keyless signing” approach, exemplified by Sigstore [104], OpenPubKey [26], and SignServer [73], eliminates long-term key handling and embeds transparency and auditability through append-only logs.
+We discuss the architecture and workflow for next-generation signing tools in section˜4.2 using the case of Sigstore.
+Effects of Signing Tools on Software Signing Adoption
+Previous studies[55, 70, 42] recognize that better tooling influences software-signing adoption.
 For example, Schorlemmer et al.
@@
 While these works emphasize broad usability factors, they offer little detailed guidance on what makes a signing tool usable, overlooking the fine-grained technical and workflow requirements that drive real-world tool adoption.
-II-B Usability Studies in Software Signing
-II-B1 Usability – Definition, Theory & Evaluation Frameworks
+2.2 Usability Studies in Software Signing
+2.2.1 Usability — Definition and Evaluation Approaches
 Tools enable engineers to adopt new software development practices [85].
-A usable tool is one whose functionality effectively supports its intended purpose—achieving product goals [18, 41].
+A usable tool is one whose functionality effectively supports its intended purpose, achieving product goals [51, 1].
 Studies assessing implementation strategies of tools are typically framed as usability evaluations [4].
 To evaluate a tool’s usability, studies typically follow a usability evaluation framework [4].
-There are two kinds: summative, which measures a tool’s effectiveness [43, 44], and formative, which provides iterative feedback for improvement [18, 45].
-Given the aims of this work, we took a formative approach.
-We considered several formative frameworks [46, 47, 48, 49] and found that Cresswell et al.’s four-factor framework [23] helped explain our data the best.
-That framework is summarized in Figure 1.
-Figure 1: Cresswell’s Usability Framework [23].
-It assesses usability based on four factors.
-Technological (T) concerns the technical characteristics of the system (e.g., performance, ease of use) and data management (availability, integrity).
-Social/Human (P) relates to users and their interactions with the tool, e.g., user satisfaction, engagement, prior experience, and correct use of features.
-Organizational Context (O) are factors internal to participants’ organization, e.g., training and leadership support.
-Macroenvironmental (M) are external influences, e.g., regulations, economics, and vendor relations.
-II-B2 Empirical Studies
-The usability of traditional signing tools—often studied in the context of encryption—remains a significant challenge.
+There are two kinds of approaches.
+Summative work measures a tool’s effectiveness [39, 16], e.g., by task success rates, completion time, and user satisfaction.
+Formative assessments provide feedback to guide ongoing improvement [51, 93].
+Given our aims, we took a formative approach.
+We considered several formative frameworks [86, 95, 89, 47] and found that Cresswell et al.’s four-factor framework [41] best explained our data.
+That framework is summarized in Figure˜1.
+Usability depends on the context in which a tool is used.
+Many factors, such as organizational objectives, personnel, policies, and compliance requirements, can make a tool usable in one setting and impractical in another [99, 50].
+This contextual aspect informs our study design.
+Figure 1: Study Methodology and Usability Framework.
+We developed our interview protocol by reviewing academic and grey literature on software signing, then conducted 17 practitioner interviews.
+Data were analyzed through thematic analysis and mapped onto Cresswell’s formative usability evaluation framework [41].
+The framework organizes usability concerns into four factors: Technological (T) — technical characteristics such as performance, ease of use, and data integrity; Social/Human (P) — user experience, and correct use of features; Organizational (O) — internal factors such as training, leadership, and resource support; and Macroenvironmental (M) — external influences including regulation, economics, and ecosystem community.
+2.2.2 Empirical Studies of Signing Usability
+The usability of traditional signing tools, often studied in the context of encryption, remains a significant challenge.
 Whitten and Tygar’s seminal study, “Why Johnny Can’t Encrypt” [99], found PGP 5.0’s interface too complex for non-experts, a problem that persisted in PGP 9 [67], especially around key verification, transparency, and signing.
-Subsequent work [50, 51, 52] has underscored the public-key model’s inherent complexity.
-Research comparing manual versus automated signing tools has yielded mixed results: some works observed no clear usability advantage [53, 54], whereas Atwater et al.
-showed a user preference for automated PGP-based solutions [55].
-Although informative, these studies focus on traditional encryption scenarios (e.g., email and messaging), limiting their applicability to broader software-signing workflows.
-These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing.
-First, they primarily focus on evaluating the usability of signing tooling in the context of signing communication (e.g., email, messages), limiting their application to signing software (i.e., securing the software supply chain).
-Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences.
-Their methodology predominantly employs summative usability assessments that rely on novice volunteers, which limits the applicability of their findings for guiding tool enhancements [56, 57].
-III Knowledge Gaps & Research Questions
-Existing literature on software signing (see section II) has examined practice and adoption across private and open-source ecosystems, highlighting the importance of tool choice in signature quality.
-Prior usability studies have focused exclusively on traditional signing tools and have been limited to evaluating user-interface design.
-However, next-generation signing tools remain understudied, creating a need for:
-  • A holistic evaluation of how software signing tools influences software-signing adoption and practice, and how usability affects this process.
-  • Practitioner perspectives on the nuanced role of tooling in shaping software-signing use and practice.
-Current research (see section II-B2) is limited because it primarily relies on experiments with novice users, while studies in section II-A2 use quasi-experiments or general practitioner interviews without a specific focus on signing tools.
-In light of these gaps, a comprehensive usability evaluation of next-generation signing tools would offer valuable insights into their role in adoption, improve their design, and enhance their adaptability.
-Research Questions: Building on identified gaps in next-generation tooling usability and its impact on adoption, we frame our research around Sigstore—our chosen next-generation signing tool case study—and ask:
+Subsequent work [14, 65, 5] has underscored the public-key model’s inherent complexity, Several works have subsequently examined the use and automation of signing in specific domains, such as social media conversations [28, 64, 8].
+Most pertinent to software signing, a recent measurement study by Schorlemmer et al.
+offers a summative assessment — that traditional software signing tools are unusable [88].
+These prior usability studies of signing tools are limited in applicability to software signing.
+Much of these works have focused on communication contexts such as email and message encryption, where users have less technical expertise [46, 43] than software engineers securing their software supply chains.
+These studies also examine only traditional signing technologies.
+We contribute the first usability study of modern “next-generation” software signing tools.
+3 Knowledge Gaps & Research Questions
+Our literature review shows a gap in formative assessments of software signing tools, Moreover, no studies have examined next-generation signing tools, which are structurally and functionally different from traditional signing tools.
+We ask:
   • RQ1:
-    What strengths and weaknesses of Sigstore influence its adoption in practice?
+    What usability strengths and weaknesses of next-generation software signing tools shape their adoption in practice?
   • RQ2:
-    What factors influence practitioners’ decisions to change the software signing tools they use?
-IV Methodology
-We selected a case study method to answer our RQs.
-Our methodology is illustrated in Figure 2.
-This section provides our design rationale (section IV-A), defines the case study context (section IV-B), discusses instrument design (section IV-C), data collection (section IV-D), and analysis (section IV-E).
+    What factors drive practitioners to adopt, retain, or replace software signing tools over time?
+4 Methodology
+To investigate our research questions, we adopted a case study approach.
+Our overall methodology is summarized in Figure˜1.
+This section details our research design, including the study rationale (section˜4.1), case study context (section˜4.2), instrument design (section˜4.3), data collection (section˜4.4), and analysis procedures (section˜4.5).
 Our Institutional Review Board (IRB) approved this work.
-IV-A Rationale for Case Study
-The case study is a research method used to investigate a phenomenon within its real-life context  [58, 59, 60, 61].
-We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice, rather than a broad but superficial cross-analysis [62].
+4.1 Rationale for Case Study
+We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice [63, 103, 23, 34], rather than a broad but superficial cross-analysis [60].
 In selecting Sigstore, we adhered to the criteria outlined by Yin [103] and Crowe et al.
@@
 Specifically, we considered the intrinsic value and uniqueness of the case study unit (high), access to data (adequate), and risks of participation (low).
-In section VI we discuss how our findings may generalize.
-Figure 2: Study Methodology.
-We developed our interview protocol by reviewing Sigstore grey literature and the academic literature.
-Finally, we analyzed our results using a thematic analysis informed by Cresswell’s formative usability evaluation framework (Figure 1).
-IV-B Case Study Context – Sigstore
+In section˜6 we discuss generalizability.
+4.2 Case Study Context – Sigstore
 Following guidelines outlined by Runeson & Höst [63], we describe the importance of our selected case study context and the technical details of the system.
-Use in Practice: Sigstore launched in 2021.
-It provides keyless, ephemeral keys and signer authentication mechanisms to simplify the signing process used by traditional tools.
-Within 13 months of its public release, Sigstore recorded 46 million artifact signatures [25].
-It has significant adoption and support from open-source ecosystems (e.g., PYPI [26], Maven [24], NPM [63]) and many companies (e.g., Autodesk [31], Verizon [29], Yahoo[64]).
-Thus, Sigstore is a major next-generation software signing platform, making it a suitable context for examining usability in modern signing tools.
-Components & Workflow: We detail Sigstore’s architecture and signing workflow, to familiarize the reader with this example of next-generation software signing tools, and to interpret participants’ statements referring to these components.
-Sigstore’s primary components are [11]:
-  • Cosign: CLI and Go library to create, verify, and bundle signatures and attestations.
-  • OIDC Identity Provider: OpenID Connect services (e.g., GitHub, Google, Microsoft) that authenticate signers and issue short-lived identity tokens.
+Use in Practice: Sigstore provides keyless signing, where developers authenticate using existing OIDC identities (e.g., GitHub, Google) instead of generating and distributing long-lived keys.
+It issues short-lived (ephemeral) certificates that simplify key management and reduce long-term security risks compared to traditional signing tools.
+Within 13 months of its public release in 2021, Sigstore recorded 46 million artifact signatures [36].
+It has since gained broad adoption across open-source ecosystems (e.g., PyPI [49], Maven [81], NPM [78]) and major companies (e.g., Autodesk [77], Verizon [76], Yahoo [102]).
+Given its rapid uptake, integration into critical ecosystems, and pioneering design features, Sigstore provides a compelling case study for examining the usability of next-generation software signing tools.
+Components & Workflow: Because our results directly reference specific Sigstore components and workflow steps, we outline Sigstore’s architecture and workflow to orient the reader.
+Following Newman et al.
+[104], the components are:
+  • Cosign: Command-line and library support (e.g., cosign, sigstore-java) for creating, verifying, and bundling signatures and attestations.
+Integrations may differ across ecosystems (e.g., GitHub Actions vs.
+Maven plugins).
+  • Identity Provider: Services such as GitHub or Google that authenticate signers and issue short-lived identity tokens (e.g., via OpenID Connect).
   • Fulcio (Certificate Authority): Issues ephemeral signing certificates after verifying the signer’s OIDC token.
   • Rekor (Transparency Log): Append-only ledger that records each signing event and certificate issuance.
-  • Monitors and Verifiers: Services or tools that validate the signer’s identity, certificate status, and signature integrity.
-Figure 3 shows the Sigstore software signing workflow.
+  • Monitors and Verifiers: Tools that validate the signer’s identity, certificate status, and signature integrity.
+Figure˜2 summarizes the Sigstore software signing workflow.
 First, the signer authenticates with an OIDC provider to obtain a short-lived token.
-Next, Fulcio verifies this token and issues a one-time signing certificate.
+Next, the certificate authority Fulcio verifies this token and issues a one-time signing certificate.
 The signer then uses Cosign to apply that certificate to the software artifact.
@@
 The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
-See section -K for legible figure
 Commonality with other Next-Gen Tools: Sigstore, like other next-generation signing tools such as OpenPubKey [26], AWS Signer [7], and SignServer [73], shares core functionality in signer identity management and automated key handling.
@@
 Instead, such services must be added via external integrations.
-While our focus on Sigstore allows for an in-depth case analysis, many of its design principles—such as identity-based authentication and ephemeral keys—are shared by its peers, suggesting potential generalizability across next-generation signing tools
-IV-C Instrument Design & Development
-Following guidance from Saldaña [66], we used semi-structured interviews to examine various aspects of tooling usability, e.g., implementation challenges, perceived benefits, and comparison to other signing tools.
-This approach let us pose consistent questions across all subjects, with the ability to probe about unique aspects of each subject’s circumstances.
-To develop our interview instrument, we did not bind our questions to any prior usability framework.
-Sigstore is a nascent technology and in such contexts an exploratory design is recommended [59, 58].
-Only after this exploratory phase did we map our findings onto a formal usability framework, ensuring it reflected practitioners’ lived experiences rather than constraining data collection to preconceived categories.
-Therefore, to construct the interview protocol, we:
+While our focus on Sigstore allows for an in-depth case analysis, many of its design principles—such as identity-based authentication and ephemeral keys—are shared by its peers, suggesting potential generalizability across next-generation signing tools.
+4.3 Instrument Design & Development
+Per section˜2.2, our study is a formative usability study.
+This type of evaluation offers methodological flexibility and prioritizes input from experienced practitioners to better articulate a system’s usability concerns and produce actionable guidance [51].
+To elicit long-form, context-rich opinions and detailed observations, we employed semi-structured interviews to examine various aspects of tooling usability (e.g., implementation challenges, perceived benefits, and comparisons to other signing tools), following guidelines from Saldaña [68].
+This approach allowed us to pose a consistent core of questions across all participants while probing unique aspects of each participant’s circumstances.
+Sigstore is an emerging (though maturing) technology, and case study research recommends exploratory designs in such contexts [103, 63].
+To develop our interview protocol, we therefore pursued open, inductive inquiry to capture the full range of practitioner usage experiences.
+After this exploratory phase, we mapped our findings onto a formal usability framework, ensuring it reflected practitioners’ experiences rather than constraining data collection to preconceived categories (section˜4.5).
+Therefore, in constructing the protocol, we:
   1. 1.
@@
   2. 2.
-    Used a snowball review of top security–conference works on signing primitives [11, 71] and empirical studies (see section II), to identify gaps and avoid redundant questions.
-Next, we refined our interview protocols through a series of practice and pilot interviews[72].
-We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
-For example, we moved the question about each participant’s team’s major product to the demographics section to provide context for their responses.
-Our full interview instrument consisted of three main topics (B-D).
-Topics B and C did not fit into the same analysis framework as Topic D.
-Thus, following the guidance of Voils et al.
-[73], we analyzed those topics in another work [BLINDED].
-[^1] That work is blinded for review.
-We attest no duplication of material.
-In this work, we focus on the data for Topic D.
-A summary of the questions in our interview protocol is shown in Table I.
-Topic D consisted of five main question categories.
-Questions D3–D4 probe current tooling experiences.
-Questions D1–D2 explore previous tools, and D5 asks about transition factors.
-See section -B for the full protocol.
-TABLE I: Summary of interview protocol. Data from Topics B and C are analyzed in [BLINDED]. We analyze Topic D in this paper. We present our full protocol in section -B
-Topic(# Questions) Sample Questions
-A. Demographics (4) What is your role in your team?
+    Used a snowball review of top papers on signing primitives [104, 55] and relevant empirical studies (see section˜2) to identify gaps and avoid redundant questions.
+We seeded the review with works from prominent cybersecurity and software engineering venues (e.g., USENIX Security, IEEE S&P, ICSE, FSE) published in the past 10 years.
+Next, we refined our interview protocols through a series of initial interviews [20].
+We first conducted practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
+Table˜1 summarizes our final interview protocol.
+After the demographic section (Topic A), our interview instrument consisted of two main topics (B & C).
+Topic B comprised contextual questions about the supply chain risks that prompt signing and how signing is implemented across the software production process (e.g., where signing is required vs.
+optional, which artifacts are signed, and relevant workflow/policy constraints).
+Topic C presented questions to assess the usability of the signing tool(s) used by the participant’s organization, including perceived strengths and weaknesses, adoption rationale, implementation details, challenges, and any considerations about switching tools.
+Table 1: Summary of interview protocol. We present our full protocol in Appendix˜B
+Topic Sample Questions
+A. Demographics What is your role in your team?
 B.
-Perceived software supply chain risks (4) Can you describe any specific software supply chain attacks (e.g., , incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
-C. Mitigating risks with signing (8) How does the team use software signing to protect its source code (what parts of the process is signing required)?
-D. Signing tool adoption (5)
-Question Categories in Topic D
-D1. What factors did the team consider before adopting this tool/method (Sigstore) over others?
-D2. What was the team’s previous signing practice/tool before the introduction of Sigstore?
-D3. How does your team implement Sigstore (which components of Sigstore does the team mostly use)?
-D4. Have you encountered any challenge(s) using this tool of choice?
-D5. Have you/your team considered switching Sigstore for another tool?
-IV-D Data Collection
-Population: Our study investigates signing tools using a case study of the Sigstore ecosystem.
-As such, our target participant pool is centered on Sigstore users.
-Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is used in signing, and the challenges and benefits of Sigstore.
-To effectively answer these questions related to decision-making, we required a target population of high-ranking or experienced practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
+Contextual Information (Risks & Use of Signing) Can you describe any specific software supply chain attacks (e.g., incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
+How does the team use software signing to protect its source code (what parts of the process is signing required)?
+C. Signing Tool Usability Questions
+C1. What factors did the team consider before adopting this tool/method (Sigstore) over others?
+C2. What was the team’s previous signing practice/tool before the introduction of Sigstore?
+C3. How does your team implement Sigstore (which components of Sigstore does the team mostly use)?
+C4. Have you encountered any challenge(s) using this tool of choice?
+C5. Have you/your team considered switching Sigstore for another tool?
+4.4 Data Collection
+Population: Our study investigates signing tools using a case study of the Sigstore tool ecosystem.
+Accordingly, our participant pool focused on current and prospective users of this tool.
+Furthermore, since our research questions pertain to organizational decision-making, we targeted expert practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
 To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [9].
@@
 This initial invitation yielded 6 participants.
-We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (5) and authors’ contacts (7) to recruit an additional 12 participants.
-We interviewed 18 subjects, but only 17 participants could share enough details for analysis.
+We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (4) and authors’ contacts (7) to recruit an additional 11 participants.
+We report on data from 17 participants.
+[^1] We conducted 18 interviews.
+Participant 18 was a senior security expert.
+While aware of the organization’s overall use of signing, they reported limited knowledge of the signing tool’s specific operation and deployment.
 Participant Demographics: Our subjects were experienced security practitioners responsible for initiating or implementing their organization’s security controls or compliance.
@@
 Subjects came from 13 distinct organizations, all companies.
-Relevant demographics are in Table II.
+Relevant demographics are in Table˜2.
 Thirteen participants reported their organizations use Sigstore.
 Two use internally developed tools and two rely on Notary v1 and PGP signing.
-Those four participants enrich our findings, as their experiences shed light on usability factors that hinder adoption.
-TABLE II: Demographics.
+Those four participants enrich our findings, as their experiences shed light on usability factors that hinder adoption of sigstore by non-users.
+Table 2: Demographics.
 For Role, “Senior management” are senior managers/directors/executives; “Technical leader” are senior/lead/partner/principal engineers; and “Engineer” and “Manager” are junior staff.
+Table˜10 gives organizational info.
 ID Role Experience Software Type
@@
 P12 Technical leader 4 years Consultancy
-S13 Senior mgmt. 16 years Internal security tool
+P13 Senior mgmt. 16 years Internal security tool
 P14 Research leader 13 years POC security software
@@
 P17 Manager 11 years Security tooling
-[…]
-IV-E Data Analysis
-We transcribed our interview recordings using www.rev.com’s human transcription service and anonymized all potentially identifiable information.
-After transcription and anonymization, our analysis began with a data familiarization stage [77] and proceeded through four phases, carried out by two analysts.
-[^2] One of the analysts is the lead author.
-In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [78].
-In the final phase, we derived themes from our inductively generated codes and then applied our usability framework to further refine the extracted themes.
-We describe these steps in detail next.
-We remind the reader that the transcript memoing and codebook was developed across the full interview transcripts, including Topics B and C.
-This manuscript focuses on the data and codes pertinent to Topic D.
-IV-E1 Memoing & Codebook Creation
-First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [79] and Campbell et al.
+Interviews: We conducted our interviews over Zoom.
+Each interview lasted $\sim$50 minutes.
+The lead author conducted these interviews.
+We offered each subject a \$100 gift card as an incentive in recognition of their expertise.
+Survey Attempt: We recognize the importance of data triangulation in case studies and therefore fielded a follow-on survey to complement our interviews.
+We deployed a survey to validate and model the prevalence and stability of our interview themes over time and across contexts.
+Despite five months of repeated solicitation in the Sigstore community, we received only 13 completed responses, of which 7 were sufficiently complete for analysis (see Appendix˜E).
+This regrettably echoes the well-documented difficulty of recruiting participants from highly specialized populations [83].
+Data triangulation is encouraged but not mandatory in case study research.
+As Stake notes, combining multiple data sources strengthens the substantiation of constructs and hypotheses, yet practical constraints necessarily shape feasible designs [82].
+We encountered such constraints.
+4.5 Data Analysis
+After transcription and anonymization, we began with data familiarization [90].
+To analyze the data, first, we used an interpretive [27] and exploratory approach for initial data analysis, as recommended for studies characterized by uncertainty [62].
+Second, because our work is a usability study, we situated the exploratory findings in a usability context by applying a formative usability framework retrospectively to our inductively generated results.
+This retrospective mapping of emergent themes onto existing constructs is an established interpretive strategy in qualitative analysis.
+This method uses prior constructs to structure and extend inductively derived findings while preserving empirical grounding (e.g., [30, 33, 18]).
+Memoing & Codebook Creation
+We began with the initial 17 transcripts.
+Two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations of O’Connor & Joffe [59] and Campbell et al.
 [40].
-This approach ensured the reliability of our analysis process while optimizing data resources.
-[^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [79].
-We randomly chose 6 transcripts (33% of data) to develop our initial codebook.
-After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos.
-We then reviewed the coded memos and finalized an initial codebook with 36 code categories.
+This process improved the reliability of the analysis while making efficient use of data resources.
+[^2] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe note that selecting 10–25% is typical [59].
+We randomly selected six transcripts ( $\sim$35% of our dataset) to develop our initial codebook.
+After each round, we discussed emerging codes in meetings, which produced 506 coded memos in total.
+We reviewed these memos to obtain a codebook of 36 categories.
 To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [54] and Campbell et al.
@@
 We obtained an 89% agreement score, suggesting a stable coding scheme [40, 59].
-We resolved disagreements through discussion.
-IV-E2 Codebook Revision
-Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round.
+We resolved disagreements by discussion.
+Codebook Revision
+Following the reliability assessment of our codebook and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six and five transcripts per round, respectively.
 In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst.
@@
 Following the recommendations of Guest et al.
-[83], we measured the cumulative number of new codes introduced in each interview.
-We saw saturation after interview 12, with no new codes after that.
-The results are presented in Figure 7.
-IV-E3 Thematic Analysis & Usability Framework Analysis
-Each interview had a median of 15 codes for topic D.
-Next, following the method proposed by Braun & Clarke [84], we derived initial themes from our code categories.
-We focused on identifying themes related to Topic D of our protocol (Table I).
-We then conducted another round of analysis using a formative usability framework. We selected Cresswell’s framework (Figure 1) based on the nature of our data.
-V Results
-First, we present a Sankey diagram illustrating the changes in signing tool usage among our participants’ teams (Figure 4).
+[32], we measured the cumulative number of new codes introduced in each interview (Figure˜8).
+We saw saturation after interview #12.
+Thematic Analysis & Usability Analysis
+Thematic Analysis: Each interview yielded a median of 15 codes.
+Next, following Braun & Clarke [13], we derived initial themes from our code categories (see Table˜1 for protocol focus).
+This reflexive thematic analysis provided an inductive account of participants’ experiences and practices.
+Usability Framework Analysis: From the emergent themes, we interpreted our findings using a formative usability framework.
+Lacking formative usability studies set in similar contexts in the Software Engineering literature, we followed advice [52, 31, 79] to seek theory from adjacent disciplines.
+Surveying candidate frameworks beyond SE, we found three groups: formative usability assessments of software systems [86]; applications of technology systems in other disciplines [47, 95, 41]; and general usability frameworks [89].
+Among these, we selected Cresswell et al.’s four-factor TPOM framework (Technology, People, Organization, Macro-environment[41] because it best aligned with the shape of our inductively generated themes (see Figure˜1).
+Our intent was to synthesize those themes into higher-level conceptual groupings that correspond to established usability determinants of adoption.
+The TPOM categories also suggest interventions, e.g., People issues may not resolve via Technology.
+We then conducted a framework-informed analysis, mapping the emergent themes to the TPOM factors to refine and structure our interpretations.
+5 Results
+We begin by presenting a Sankey diagram (Figure˜3) that illustrates the distribution of our participants between Sigstore and non-Sigstore users, and how their teams changed signing tool usage over time.
 Many participants (6) transitioned from using no tools to adopting Sigstore.
-Others transitioned from existing tools to Sigstore: 3 from GPG, 2 from Notary, 1 from Skopeo, and 2 from proprietary or internal tools.
-The final 4 subjects did not transition to Sigstore.
+Others transitioned from existing tools to Sigstore: 3 from implementations that had some combinations of GPG/PGP (of which one practitioner reported combining PGP with Skopeo[61]), 2 from Notary, and 2 from proprietary or internal tools.
+4 subjects did not adopt Sigstore.
 Figure 3: Practitioners’ Changes in Software Signing Tools.
-To understand why they did (and did not) change tools, we structure the results by RQ using Cresswell’s framework.
-V-A RQ1: Strengths & Weaknesses of Sigstore
+To understand the usability concerns that Sigstore adopters face; and why practitioners did (and did not) change tools, we structure the results by RQ using Cresswell’s framework.
+5.1 RQ1: Strengths & Weaknesses of Sigstore
 The first aim of our research work was to ascertain from the the experiences of our subjects the strengths and weaknesses they have experienced while using Sigstore.
-V-A1 Sigstore’s Strength
-We present a summarized list of practitioner-reported advantages of using Sigstore in Table III.
-First, we extracted the benefits our participants experienced while using Sigstore from their responses.
+5.1.1 Sigstore’s Strength
+We extracted the benefits our participants experienced while using Sigstore from their responses.
 These reflections capture their firsthand experiences with Sigstore.
-TABLE III: Summary of Practitioner-Reported Advantages of Using Sigstore. Practitioners cited Sigstore’s technological and macroenvironmental factors as advantages.
+We grouped their responses into eight categories as summarized in Table˜3.
+When mapped to the Cresswell framework, these factors fell into two categories: Technological and Macroenvironmental.
+Table 3: Summary of Practitioner-Reported Advantages of Using Sigstore. Practitioners cited Sigstore’s Technological and Macroenvironmental factors as advantages.
 Topics & Associated Examples Subjects
 Technological Factors
-Ease of Use 8
-Signing Workflow & Verification P1, P2, P7, P9, P14-17
-Setting up with automated CI/CD actions P9
-No key distribution problems P1
-Use of Short-lived Keys & Certificate 3 –P2, P3, P15
-Signer ID Management 4
-Use of OIDC(Keyless) to authenticate signers P3, P5, P10, P12
-Compatibility with Several New Technologies 4
-Integrability with SLSA build P16
-Integrability with several container registries/technologies P12, P14
-Integrability with several cloud-native applications P15
-Precence of a Transparency Log 3
-Transparency logs increase security P5, P14
-Evaluation of signing adoption using logs P9
-Bundling Signatures With Provenance Attestations 2 – P3, P4
-Reliability of Service 1 – P7
+Ease of Use 8 subjects
+1. Signing Workflow & Verification P1, P2, P7, P9, P14-17
+2. Setting up with automated CI/CD actions P9
+3. No key distribution problems P1
+Use of Short-lived Keys & Certificate 3 subjects P2, P3, P15
+Signer ID Management 4 subjects
+1. Use of OIDC(Keyless) to authenticate signers P3, P5, P10, P12
+Compatibility with Several New Technologies 4 subjects
+1. Integrability with SLSA build P16
+2. Integrability with several container registries/technologies P12, P14
+3. Integrability with several cloud-native applications P15
+Precence of a Transparency Log 3 subjects
+1. Transparency logs increase security P5, P14
+2. Evaluation of signing adoption using logs P9
+Bundling Signatures With Provenance Attestations 2 subjects P3, P4
+Reliability of Service 1 subject P7
 Macroenvironmental Factors
-Free/Open-Source 2 – P7, P17
-We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service.
-When mapped to the Cresswell framework, these factors fell under two categories: Technological and Macroenvironmental.
+Free/Open-Source 2 subjects P7, P17
 Technological Factors
-Ease of use: This was the most frequently discussed advantage of Sigstore over other Software signing tools.
-Participants commonly highlighted the simplicity of following the Sigstore workflow—particularly how key management is fully automated reduce configuration overhead.
-P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
+We identified five factors:
+Ease of use: This was the most frequently discussed advantage of Sigstore over other tools.
+Participants commonly highlighted the simplicity of the Sigstore workflow.
+As P7 (security tools, cloud security) put it, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other ease-of-use advantages mentioned were the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
 Identity management for Signers: Sigstore’s integration of OIDC identity management was a strength per our participants.
@@
 In most cases, the resources required to set up connections to these technologies were minimal or nonexistent.
-P12 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
-Use of Short-lived Keys/Certificates: Sigstore’s enforcement of short-lived keys and certificates as a security feature was recognized by our participants as a major strength of the tooling.
-Traditional software signing tools require users to manually set expiration times for cryptographic elements such as signatures, keys, and certificates at the time of generation.
-While this approach provides customization flexibility, it also introduces the risk of long-lived cryptographic materials remaining valid for extended periods, increasing the likelihood of compromise.
-Participants appreciated Sigstore’s automated management of key and certificate lifetimes, which mitigates this risk.
-P2 (SAAS, SSC security tool) states, “So (with) PGP, you still have to figure out the key distribution problem.
-Because Sigstore uses short-lived keys that solves a lot of those issues around key storage and key distribution…”
-Presence of a Transparency Log: The ability to audit signing actions was crucial for adopters.
-Participants valued the transparency log in the Sigstore ecosystem because it provides a tamper-resistant record of all changes to an artifact’s metadata—functionality not offered by other signing systems.
-This auditability gave users greater confidence in the integrity of signed artifacts.
+P12 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers… Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is easy.”
+Use of Short-lived Keys/Certificates: Participants appreciated Sigstore’s enforcement of short-lived keys and certificates.
+Some signing tools require users to manually set expiration times for cryptographic elements such as keys and certificates at the time of generation.
+That customizability also introduces the risk of long-lived cryptographic materials, increasing the likelihood of compromise.
+Participants preferred Sigstore’s automated management of key and certificate lifetimes.
+As P2 (SAAS, SSC security tool) stated, “PGP, you still have to figure out the key distribution problem….Sigstore uses short-lived keys, that solves a lot of those issues.”
+Presence of a Transparency Log: Adopters valued the ability to audit signing actions.
+Sigstore’s transparency log provides a tamper-resistant record of all changes to an artifact’s metadata, a functionality not offered by other signing systems.
 {quoting} P9 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
-Other technological factors noted by participants are the reliability of Sigstore’s public instance.
 Macroenvironmental Factors
@@
 They noted that this reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments.
-{quoting} P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
-V-A2 Sigstore’s Weaknessses
-We summarize the difficulties reported by participants in using Sigstore in Table IV. Most of these weaknesses mapped to multiple usability factors.
-TABLE IV: Reported Difficulties Encountered by Participants in Using Sigstore.
-We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.
+{quoting} P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public” This enthusiasm reflects both the appeal of free/open-source (FOSS) software and the practical benefits of public accessibility.
+However, note that open-source $\neq$ free; enterprises may pay for private deployments, support, and integration.
+5.1.2 Sigstore’s Weaknessses
+We summarize the difficulties reported by participants in using Sigstore in Table˜4. Most of these weaknesses mapped to multiple usability factors.
+Table 4: Summary of Practitioner-Reported Difficulties Using Sigstore. Difficulties are grouped by Cresswell factors.
 Topics & Associated Examples Subjects
-Enterprise Adoption Limitations 7 subjects
-Rate Limiting Problems – T P3, P7, P14, P15
-Lack of dedicated Support & Maintenance – T & P P15, P2
-Not Suited for Regulated Organizations – M & O P3
-Latency Concerns – T P6
-Transparency Log Issues 6 subjects
-Not Suitable for private Setup – T P2, P3, P6, P14, P17
-Use in Air Gap Conditions – T P2, P3, P9
-Efforts to Monitor Logs – P P2
-Private Sigstore Instance Setup 5 subjects
-Documentation – P & T P9, P12, P6
-Limited Community Support – M P15
-Infrastructure Requirements & Maintenance Costs – T P5, P6
-Other Documentations and Usage Information IssuesP & T 3 — P1, P14, P17
-Integration to Other Systems 3 subjects
-Attestation Storage – T P1
-Gitlab & Jenkins – T P9
-Other Unsupported technologies – T P16
-Offline CapabilitiesT 2 – P3, P4
-Fulcio Issues 1 – P3
-Timestamping Issues – T
-Fulcio-OIDC Workflow – T
-Software LibrariesT 1 – P7
+Technological Factors
+Transparency Log Constraints 6 subjects
+1. Not suitable for private setup P2, P3, P6, P14, P17
+2. Use in air-gapped/offline conditions P2, P3, P9
+Performance & Rate Limits 6 subjects
+1. Rate limiting problems P3, P7, P14, P15
+2. Latency concerns P6
+Integrations & Tooling 3 subjects
+1. GitLab/Jenkins P9
+2. Other unsupported technologies P16
+3. Attestation storage P1
+Documentation (technical aspects) 6 subjects
+1. Private instance setup documentation P9, P12, P15
+2. Other documentation / usage issues P1, P14, P17
+Private Instance: Infra & Cost 2 subjects
+1. Infra requirements & maintenance cost P5, P6
+Fulcio / Timestamping Workflow 1 subject
+1. Timestamping issues P3
+2. Fulcio–OIDC workflow P3
+Software Libraries 1 subject
+1. Unsupported software libraries P7
+Social / Human Factors
+Log Monitoring Burden 1 subject
+1. Effort to monitor logs P2
+Support & Maintenance 3 subjects
+1. Lack of dedicated support & maintenance P15, P2, P6
+Organizational Factors
+Regulatory Suitability (Cross-factor) 1 subject
+1. Not suited for regulated organizations — M & O P3
+Macroenvironmental Factors
+Community Maturity/Support 1 subject
+1. Limited community support P15
 Macroenvironmental & Organizational
+The primary weakness along this dimension related to Sigstore’s use in large enterprise contexts.
 Enterprise Adoption Limitations: Many of our participants mentioned that Sigstore is not suited for large enterprise applications.
@@
 {quoting} P7 (security tools, cloud security): “We were relying on the public Sigstore instance, and I don’t think it could meet our signing needs in terms of just the capacity of signatures we needed as an enterprise [rate limit].” {quoting} P3 (SAAS, SSC security tool): “For customers that are very large in scale, the biggest of the Fortune 100 or customers operating in very highly regulated environments, I think operating your own Sigstore instance within that air gap in a private environment could be valuable.
-And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.” The issues here were a mixture of all usability factors.
+And I think it depends on the amount of software you’re producing and the frequency with which you’re producing it.”
 Technological
+We identified seven technological issues.
 Transparency Log Issues: While the transparency log bundled with Sigstore was highlighted as a major advantage by participants, it was also a significant concern.
@@
 Issues here were primarily related to participants’ need for support resources, which were macroenvironmental and social in nature.
-{quoting} P15 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to…deploy Fulcio, like on prem, or in their own infrastructure.”
+{quoting} P15 (internal security tooling, aerospace security): “There’s not a big enough community of practitioners, or help to…deploy Fulcio, like on premises.”
 Other Documentation Issues: Overall, the documentation for Sigstore, along with practical usage information and support, was found to be insufficient.
@@
 However, in this case, participants specifically desired the ability to use Sigstore’s signing and verification features in offline scenarios.
-Software Libraries: The programming libraries of Sigstore were also mentioned as a challenge, particularly in cases where users wanted to integrate Sigstore’s signing capabilities directly into an application.
-V-A3 Effect of Identified Problems on Sigstore components used by Subjects
+Software Libraries: The programming libraries of Sigstore were also reported as a challenge, particularly when participants sought to embed signing capabilities directly into applications.
+They noted the absence of well-supported, easy-to-use client libraries, which forced them to rely on workarounds such as shelling out to the CLI.
+5.1.3 Impact of Issues on Sigstore Component Use
 We assessed the effect of participants’ reported problems with Sigstore on their usage of each component.
@@
 Among these, our data shed the most light on local deployments and the Rekor log.
-For local deployment, several participants cited difficulties with documentation and setup, and as a result, only two successfully deployed private instances, likely due to these challenges (Table IV, section V-A2).
+For local deployment, several participants cited difficulties with documentation and setup.
+Only two successfully deployed private instances, likely due to these challenges (Table˜4, section˜5.1.2).
 The Rekor transparency log presents a mixed case, with both high adoption and significant reported issues.
@@
 Although Rekor was recognized as a key factor driving adoption due to its strengths, it also faced a high number of reported issues, indicating that, despite its benefits, challenges persist for some participants.
-Other Components Without Mention Though the participants in the survey shared experiences with user-facing components, discussion about other elements was notably absent.
-In particular, there was very little mention of log witnessing [86], or monitoring [87] for either Fulcio or Rekor.
+Other Components Without Mention
+Though the participants in the survey shared experiences with user-facing components, discussion about other elements was notably absent.
+In particular, there was very little mention of log witnessing [35] or monitoring [3] for either Fulcio or Rekor.
 This is remarkable, as transparency solutions require the existence of these parties to ensure log integrity and identify log misbehavior.
+The lack of discussion may suggest that practitioners are either not yet engaging with these components or may not fully recognize their importance in correct use of the ecosystem.
 We believe that targeted future work should identify possible gaps in user perception, as well as identify user stories so users verify log witnesses and monitor information.
-V-B RQ2: How & Why Practitioners Change Tools
-V-B1 Drivers of Adoption Among Current Sigstore Users
+5.1.4 Effects of Organization and Signing Tools on Identified Usability Concerns
+        (a) Subjects from small organizations (6 subjects).         (b) Subjects from large organizations (7 subjects).
+Figure 4: Tools vs.
+Evaluation Criteria.
+Reported usability issues across organizational contexts.
+Heatmaps show the number of unique participants raising each issue; darker cells indicate higher concentrations of reports.
+Abbreviations: DOC = Documentation (technical); TIME = Fulcio / Timestamping Workflow; INTEG = Integrations & Tooling; LOG = Log Monitoring Burden; PERF = Performance & Rate Limits; COST = Private Instance: Infrastructure & Cost; REG = Regulatory Suitability (Cross-factor); SUPP = Support & Maintenance; TLOG = Transparency Log Constraints; CMAT = Community Maturity/Support.
+To provide additional context for the usability issues identified above, we highlight three observations from our analysis of organizational size and participant roles.
+See Figures˜4 and 9 for visualizations.
+Large and small organizations share similar pain points
+Across both large and small organizations, the densest cells occur for topics related to private/air-gapped deployments, transparency-log constraints, and documentation for customized/ private instances.
+In large organizations, technical leaders frequently surface additional concerns around support and maintenance expectations.
+In contrast, smaller organizations are more willing to combine Sigstore with other tooling to compensate for gaps.
+Notably, large organizations in our sample were more likely to retain internal/proprietary solutions or delay full Sigstore adoption.
+A usability lesson here is to prioritize private-deployment guidance (including monitoring playbooks), clarify support boundaries, and provide migration patterns and user stories for mixed stacks.
+Scale effects: rate limits and latency concentrate in larger orgs
+Performance and rate-limit issues are most prominent in medium/large organizations, consistent with higher CI/CD concurrency.
+This suggests the need for prescriptive CI recipes (e.g., batching, caching, retries/backoff, and concurrency guidance) alongside clearer Fulcio/OIDC and timestamping failure-mode documentation.
+Medium organizations tend to “work around” with plain Sigstore
+Participants from medium-sized organizations more often stick with plain Sigstore and work around gaps rather than self-host.
+This group benefits most from concise integration blueprints (GitHub/GitLab/Jenkins), ready-to-use attestation storage patterns, and targeted documentation updates, rather than heavy private-infrastructure guidance.
+5.2 RQ2: Why Practitioners Change Tools
+5.2.1 Drivers of Adoption Among Sigstore Users
 For participants who adopted Sigstore, we asked which factors—prior to adoption—influenced their decision (or their organization’s decision) to switch.
 Their responses fell into three main categories: macroenvironmental factors (e.g., regulations and user communities), human/social factors (e.g., prior experience with signing tools), and technological factors (e.g., Sigstore’s unique capabilities).
-Most of our participants chose to switch to Sigstore due to poor experiences with previous tools and due to perceived advantages associated with Sigstore. We summarize these factors in Table X.
+Most of our participants chose to switch to Sigstore due to poor experiences with previous tools and due to perceived advantages associated with Sigstore. We summarize these factors in Table˜11.
 Contributions to Sigstore: Being part of the Sigstore community as contributors was a major reason for adopting Sigstore.
@@
 P14 (POC software & security tool, cloud security) reflects this, “I’d say many of the early staff, the first 10 to 20 staff were involved in the Sigstore community.
-There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.”
+There’s not just our interest as a company in a Sigstore, but there’s a lot of personal ties to Sigstore.” While this pattern may partly reflect our interview sample, it highlights how prior involvement and community ties can strongly shape adoption decisions—a dynamic that could also apply more broadly in open-source ecosystems.
 Available Sigstore Functionalities: Sigstore’s functionalities also influenced its adoption.
-While several participants were drawn to the transparency log (as highlighted in section V-A1).
+While several participants were drawn to the transparency log (as highlighted in section˜5.1.1).
 {quoting} P5 (internal security tool, cloud security): “Sigstore has other features, like you can use your OIDC identities, like your GitHub identities as well, to do the signing, [and] they also maintain a transparency log.”
@@
 Some participants who previously used other tools (2) switched to Sigstore due to security frameworks and standards recommending it as a best practice.
-Additionally, the remaining participants (2) noted that regulations played a partial role in their decision to adopt Sigstore.
+Additionally, the remaining participants (2) noted that regulations (e.g., E0-14028) played a role in their decision to adopt Sigstore.
 {quoting}
@@
 Conversely, participants P1 and P14 cited GPG’s smaller user base as a drawback.
-Additionally, some participants expressed trust in industry consortia such as CNCF (Cloud Native Computing Foundation) {quoting} P3 (SAAS, SSC security tool): “We do defer often at the higher-level projects to different foundations.
+In addition, some participants expressed trust in industry consortia such as the CNCF (Cloud Native Computing Foundation), noting that Sigstore’s development and governance under CNCF gave it additional legitimacy and credibility compared to tools maintained by smaller, independent groups.
+{quoting} P3 (SAAS, SSC security tool): “We do defer often at the higher-level projects to different foundations.
 So we would be more likely to trust something from the CNCF…[than] independent developer projects.”
-GPG & Notary Issues: Issues associated with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore.
-GPG-based signing implementations have historically been criticized for key management challenges and usability concerns.
-Our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies.
+GPG & Notary Issues: Issues with previously existing tools played a major role in practitioners’ decisions to switch to Sigstore.
+Both GPG- and Notary-based signing implementations were reported as problematic.
+GPG has long been criticized for key management challenges and usability concerns, and our participants echoed these issues, citing additional factors such as low adoption rates, steep learning curves, and limited integration with modern technologies.
 Notary-related issues primarily concerned customer demands, compatibility with other technologies, and infrequent updates.
@@
 Meanwhile, users of proprietary signing tools faced difficulties with complex setup processes.
-V-B2 Barriers and Motivators Spurring Consideration of Sigstore Among Non-Sigstore Users
-TABLE V: Non Sigstore Users: Reasons for Not Adopting Sigstore.
+5.2.2 Considerations from Non-Sigstore Users
+Table 5: Non Sigstore Users: Reasons for Not Adopting Sigstore.
 This table summarizes the reasons non-users gave for not adopting Sigstore, grouped by their current signing tool and mapped to our usability framework theme.
@@
 In this section we highlight reasons mentioned by practitioners that have stopped them (or their organization) from switching to Next-gen tools like Sigstore.
-We also report issues experienced by these non sigstore users that have prompted a consideration (or in some cases partial adoption of these next-gen tools).
-Barriers to Considering Sigstore: As highlighted in Table V, most reasons for non-adoption were organizational.
-{quoting} P8 (internal security tool & cloud APIs, internet services): “When it comes to open source projects and open source stuff, our Organization is very skeptical.
-Policy and everything drives very much teams away from that sort of stuff…There’s also the ongoing risk of having a product that’s built outside continually shipping code and being integrated into this product, where it deals with sensitive material.
-So we do evaluate these things, but the scrutiny against code written outside of the company is extreme, and with good reason given the context of all the supply chain stuff.” Other notable issues arose from technical concerns, including privacy risks associated with Sigstore’s public instance and the inherent risk of relying on a third-party tool maintained by open-source contributors.
-Factors Motivating Consideration for sigstore: When practitioners were asked about issues with their current tools and potential fixes, most pointed to existing tooling problems as the primary driver for considering alternatives—such as incorporating next-gen features into internal tools or partially adopting new solutions.
+We also report issues experienced by these non-Sigstore users that have prompted a consideration (or in some cases partial adoption) of these next-gen tools.
+Barriers to Considering Sigstore: As highlighted in Table˜5, most reasons for non-adoption were organizational.
+{quoting} P8 (internal security tool & cloud APIs, internet services): “When it comes to open source projects and open source stuff, our organization is skeptical.
+Policy and everything drives teams away…There’s also the ongoing risk of having a product that’s built outside continually shipping code and being integrated into this product, where it deals with sensitive material.
+We do evaluate these things, but the scrutiny against code written outside of the company is extreme, and with good reason given the [supply chain] context.” Other notable issues arose from technical concerns, like privacy risks associated with Sigstore’s public instance and Macro environmental concern of the inherent risk of relying on a third-party tool maintained by open-source contributors.
+Factors Motivating Consideration for Sigstore: When practitioners were asked about issues with their current tools and potential fixes, most pointed to problems with existing tools as their main driver to alternatives, e.g., incorporating next-gen features into internal tools or partially adopting new solutions.
 Although some subjects reported drawing design inspiration from next-generation workflows, none specifically mentioned Sigstore.
-P11 (firmware & testing software, social technology) states, “.
-I’m sure at some point in time it was influenced by an open source project before they adopted it and made the right tweaks internally.” Lack of transparency and integration of these tools to other platforms were also common.
-We summarize these factors in Table XI.
-VI Discussion
-VI-A Lessons for Next-Generation Tools
-Next-generation signing tools offer unparalleled flexibility by integrating with external platforms, and services.
-However, our participants consistently highlighted integration challenges as a major obstacle.
-VI-A1 Automated Cross-Platform Integrations
-To fully leverage the flexibility of next-gen tools, practitioners need integration modules for popular CI/CD systems, registries, and orchestration platforms.
-Missing integration and sparse documentation forces users to implement custom connectors, increasing the risk of misconfiguration and potential security vulnerabilities that undermine the very guarantees signing is meant to provide.
-VI-A2 Need for Inter-Platform Standards
-Standardization of information exchange is a standard problem for new classes of software infrastructure e.g., , SBOM [88].
-These problems often harm integration capabilities.
-Next-generation signing tools would gain from common exchange standards—for example, uniform schemas for transparency logs or identity assertions.
-Projects like Diverfiy [89], which harmonize identity provisioning across multiple OIDC providers, are a promising start.
-Similar efforts are needed to streamline data flows between signing components, transparency services, and external tooling ecosystems.
-VI-A3 Verification Workflow
-A key security property in a software supply chain is the transparency of actors and artifacts, which requires verifying signer identities [4].
-Previous studies have highlighted challenges in signature verification across both organizational and open-source contexts [39, 5].
-Next-generation tools like Sigstore integrate multiple verification methods, such as the Cosign CLI and the transparency log, but most participants still reported usability issues with the transparency log, and documentation indicates a lack of automation for other verification methods.
-Further automating the verification workflow would enhance the security benefits of software signing with this class of tools.
-VI-B Usability Patterns of Sigstore vs Non-Sigstore users
-Our results reveal different drives in adoption motivations and shared usability concerns between Sigstore users and non-users.
-Sigstore adopters cited contributions to the Sigstore project (Human/Social factor) and macroenvironmental factors—alongside negative experiences with legacy tools—as their primary drivers.
-In contrast, non-users emphasized organizational considerations when choosing their signing solution.
-Despite these differences, both groups reported similar technological usability issues, particularly around integration with other systems and infrastructure transparency.
-Notably, non-users frequently highlighted signer identification challenges, whereas Sigstore users did not—underscoring the clear usability advantage of next-generation tooling.
-VI-C Privacy Concerns
-Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern.
-Privacy concerns were also major issues influencing users not to consider sigstore as suitable for enterprise.
-This issue also affects the use of OIDC for keyless signing.
-While this challenge is not unique to Sigstore [71] and is a recurring issue in software signing and modern cryptography, it underscores the need for continued innovation to address privacy-preserving mechanisms that balance transparency with user privacy, ensuring broader adoption and trust in cryptographic tools like Sigstore.
-While works like Speranza[71] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
-VII Threats to Validity
+P11 (firmware & testing software, social technology) states, “I’m sure … [the internal signing tool] was influenced by [an open-source next-gen tool]…they adopted it and made the right tweaks internally.” Lack of transparency and integration of these tools to other platforms were also common.
+We summarize these factors in Table˜12.
+5.2.3 Effects of Organizational Context
+Figure 5: Push-pull-barrier heatmap.
+Combining the results in section˜5.2.1 (Table˜11) and section˜5.2.2 (Tables˜12 and 5) with organizational size, former signing tools, and current signing tools reveals several patterns.
+We summarize these in the Push-pull-Barrier heatmap (Figure˜5)
+Push, Pull, and Barrier factors influencing adoption
+We grouped issues raised by participants into Push, Pull, and Barrier factors.
+We define these factors drawing on Bansal et al.’s Push-Pull-Mooring framework for consumer switching behavior [10].
+Push factors are problems with prior tools that drove current users toward Sigstore.
+Pull factors are attractions of the new tool (and ecosystem).
+Barriers are reasons that stalled or prevented adoption.
+While large organizations exhibit the greatest number of barriers, they report push and pull factors in roughly similar measure.
+By current-tool status, many push factors come from teams now using Sigstore alongside other mechanisms (e.g., attestations, internal tooling, PGP), which often reflects phased introduction of Sigstore into existing processes (and vice versa).
+A summary of the coded factors appears in Table˜13.
+A usability takeaway is that switching is primarily driven by human/social push (e.g., frustration with legacy workflows) and only secondarily by pull from Sigstore features and regulatory pressure (Sigstore being referenced by some standards).
+Thus, even as technical innovations “pull” adoption, community-building and other social supports also materially “push” change.
+Non-adoption reasons are often organizational and non-technological
+Non-adoption reasons are often organizational and non-technological.
+Although non-adopters are both pushed by shortcomings in internal tools and pulled by next-generation capabilities, adoption frequently stalls due to organizational constraints such as risk posture, support expectations, or regulatory requirements.
+This pattern is not unique to software signing—similar dynamics have been observed in other infrastructure domains (e.g., cloud technologies[6])—but it is particularly salient here because identity-based signing tools must be trusted across organizational boundaries.
+Our interpretation is therefore not that next-generation tools have failed, but rather that their technical attractiveness has created demand for more mature deployment models: private instances, enterprise-grade support, and stronger service guarantees.
+Tool migration patterns
+In our sample we observe recurring migration paths: PGP $\rightarrow$ Sigstore (sometimes retaining PGP during transition); Notary v1 $\rightarrow$ Sigstore or Notary v2; and internal/proprietary $\rightarrow$ Sigstore (or hybrid).
+Organizations mostly transition between tools slowly.
+Larger organizations are more likely to prefer internal/hybrid arrangements (reflecting policy, privacy, and support concerns), whereas small/medium organizations are more likely to adopt “pure” Sigstore sooner, using hybrids as transitional steps or to meet offline/private constraints.
+6 Discussion
+We highlight the main takeaways from our results, noting that this research is a case study of next-generation software signing tools (section˜2.1) using Sigstore.
+We discuss implications for this class of tools along three axes: a technological readiness assessment; their transformative potential; and the implications of the observed usability patterns
+6.1 TRLs and Next-Generation Signing Tools
+Because Sigstore, an exemplar of next-generation (identity-based) signing, is relatively recent, the maturity of its core capabilities is uneven.
+Understanding this maturity helps organizations assess implementation risk and prioritize investment [53].
+Our usability study provides a preliminary signal.
+Drawing on participant reports and interpreting them through the Technology Readiness Level (TRL) lens [53], we note that TRL is a technology-maturation framework (not a usability scale).
+In practice, lower TRLs (concepts, prototypes) tend to surface mostly technical concerns, while higher TRLs (operational environments) bring non-technical constraints (organizational fit, governance, support/SLAs, regulation) to the foreground.
+Our results fit this progression: technological factors were most frequently discussed.
+Several capabilities, e.g., short-lived (ephemeral) keys, OIDC-based signer identity, and GitHub Actions workflows, received consistently positive mentions across different organizational contexts, suggesting higher maturity for these components (and analogous features in similar next-gen tools).
+In contrast, capabilities such as private/self-hosted deployments and operation in air-gapped or regulated environments surfaced moderate friction, indicating mid-level readiness and a need for stronger documentation, reference architectures, and support pathways.
+For organizations that have not yet adopted next-generation tools, these findings suggest comparatively lower risk for the first set of capabilities, while private/regulated deployments warrant careful evaluation and piloting.
+These assessments are preliminary and should be read in light of our case-study scope.
+From Sigstore to the broader class: Sigstore demonstrates that identity-based signing can work at scale (across multiple organization sizes and use contexts in our sample), serving as a proof-of-concept for the class.
+Where we observed Sigstore-specific friction (e.g., private Fulcio/Rekor deployments, public-log privacy trade-offs), the underlying class implications are clear: enterprise deployment models, integrations, and operational guardrails are the next maturity hurdles for identity-based signing generally.
+6.2 Transformative Potential
+Sigstore, as the most widely deployed exemplar of next-generation (identity-based) signing, goes well beyond the minimal OIDC–key binding model offered by alternatives like OpenPubKey.
+In addition to identity-based signing and short-lived keys, Sigstore integrates a transparency log (Rekor) and broad ecosystem compatibility with attestations, container registries, and CI/CD workflows.
+Participants repeatedly emphasized these additional capabilities as unique differentiators, even while reporting usability frictions.
+This dual reality—strong adoption coupled with persistent friction—underscores Sigstore’s role as a living testbed for both practice and research.
+Sigstore is already being leveraged in production ecosystems (e.g., Maven, PyPI, Kubernetes) and has catalyzed new research trajectories, such as privacy-preserving transparency [44] and diversified identity provisioning [57].
+The broader implication is that Sigstore is not merely a tool to be adopted, but an infrastructure on which future signing solutions and empirical studies can (and are) being built.
+Widespread adoption makes its usability challenges consequential — and resolving them has cascading benefits across the supply chain.
+Its design decisions and shortcomings provide researchers with concrete leverage points for advancing the next generation of signing tools.
+6.3 Lessons for Next-Gen Signing Usability
+While our results reveal different adoption drivers between sigstore users and non users, there exists some shared usability concerns—between Sigstore users and non-users (section˜5.2).
+Both groups reported similar technological issues, particularly around integration with other systems and the transparency of signing infrastructures.
+This suggests cross-cutting concerns that affect not only Sigstore users but also non-users.
+A preliminary look at enhancement/issue trackers for other next-generation signing tools (e.g., OpenPubkey) shows similar themes [58].
+Building on these observations, we distill actionable lessons for next-generation tools and outline opportunities for future work.
+Automated Cross-Platform Integrations: To fully leverage flexibility, practitioners need maintained integration modules for popular CI/CD systems, registries, and orchestration platforms.
+Missing integrations and sparse documentation force teams to build custom connectors, increasing misconfiguration risk and undermining security guarantees.
+Need for Inter-Platform Standards: Standardization of information exchange is a typical problem for new infrastructure (e.g., SBOM [101]).
+Next-generation signing would benefit from uniform schemas for transparency events and identity assertions.
+Recent efforts like DiVerify [57], which harmonizes identity provisioning across multiple OIDC providers, take a step toward smoother data flows between signing components, transparency services, and external toolchains.
+Verification Workflow: A core supply-chain property is transparency of actors and artifacts, which requires reliable, usable verification of signer identities [15].
+Prior work reports verification pain in both organizational and open-source settings [42, 71].
+Although Sigstore provides multiple verification paths (Cosign CLI, transparency log), participants reported log-related usability issues and limited automation elsewhere.
+Further automating verification would better realize the security benefits of this tool class.
+Privacy Concerns: Privacy concerns emerged prominently around Sigstore’s Rekor transparency log: five of the six participants who reported concerns with Rekor, cited privacy as a significant issue.
+Similar concerns also surfaced for keyless signing via OIDC.
+These are not unique to Sigstore [55] but reflect broader tensions in cryptography between transparency and confidentiality.
+This underscores the need for privacy-preserving designs (e.g., selective disclosure, redaction regimes, witness policies) that balance auditability with confidentiality.
+In Creswell’s terms, these concerns span macroenvironmental (policy/regulation), organizational (data governance and risk posture), and human/social (developer risk perception) factors, reinforcing that privacy is not solely a technical hurdle but a cross-factor adoption barrier.
+Population Scope & Future Populations: Our study primarily sampled experienced practitioners operating within organizations, which shapes both the opportunities and frictions we observed.
+Usability needs likely differ for other communities and skill levels—e.g., individual open-source maintainers, small/startup teams, and security novices.
+In light of the biases from our particular sample, we recommend further research that (1) uses organizational contexts and user personas as a primary selection criterion; and (2) considers distinct signing use cases for a more fine-grained understanding.
+7 Threats to Validity
 To discuss the limitations of our work, we follow Verdecchia et al.’s [97] recommendation to reflect on the greatest threats to our research.
-In addition, we discuss how our experiences and perspectives may have influenced our research (positionality).
+In addition, we discuss how our perspectives may have influenced our research (positionality).
 Construct Validity: Cresswell’s framework targets health IT, where “tooling” denotes integrated clinical systems, so applying it to software signing carries construct risks.
@@
 To mitigate this subjectivity, we utilized multiple raters and measured their agreement.
-External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore.
-We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature.
+External Validity: The primary generalizability threat is that we recruited from a population biased toward the use of Sigstore.
+We viewed this as necessary in order to recruit sufficient subjects.
 A second threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects.
-We attempted a survey (details are in section -E) but received only a few responses, echoing the challenges of surveying specialized communities [91].
 A third threat is our study’s temporal scope.
 Interviews were conducted between November 1, 2023, and February 9, 2024, capturing Sigstore users’ experiences during that period.
-Since then, Sigstore has introduced some changes, but not the core workflows and components.
-This suggests our findings remain relevant.
+Sigstore has changed since then, but not the core workflows and components.
+We believe our findings remain relevant.
 Statement of Positionality: The author team has expertise in software supply chain security and qualitative research methods, e.g., publishing top-tier papers in software engineering and cybersecurity venues.
@@
 To avoid bias, this author was not involved in developing the research design nor the initial analysis of the results, but provided insight in later analysis.
-VIII Conclusion
-Software signing guarantees provenance via cryptographic primitives, reducing supply-chain attacks.
-Next-generation tools like Sigstore automate traditional challenges, yet their usability remains underexplored.
-We interviewed practitioners to examine Sigstore’s usability, identifying key adoption factors.
-Our findings show that Sigstore’s identity-based workflows improve usability, but integration complexities, privacy concerns, and organizational constraints hinder adoption.
-We recommend toolmakers provide official integration plugins and privacy controls for transparency logs.
-These insights inform automation improvements, promoting secure, usable next-generation signing in automated software engineering.
-IX Acknowledgments
+8 Conclusion
+Software signing underpins supply chain security by guaranteeing provenance and integrity, yet its effectiveness depends on usable tooling.
+Next-generation systems like Sigstore reduce traditional burdens of key management and signer identification, but our study shows that adoption is shaped by both strengths and frictions.
+Identity-based workflows, ephemeral keys, and CI/CD compatibility ease use, while integration gaps, transparency log privacy concerns, and organizational constraints remain barriers—especially in large enterprises.
+By surfacing these factors, we show that usability is central to supply-chain security.
+Without attention to adoption barriers, even the best technical mechanisms will fail to achieve their security potential.
+Our findings reveal uneven maturity across Sigstore’s components, and suggest lessons for other signing toolmakers, e.g., OpenPubKey and SignServer which share similar design patterns.
+We recommend that toolmakers prioritize official integrations and configurable privacy controls to improve usability and trust.
+By addressing these challenges, next-generation signing tools can expand the coverage of verifiable artifacts and attestations, shrinking supply chain attack surfaces to improve software security.
+9 Acknowledgments
 We thank the study participants for contributing their time to this work.
@@
 We also thank Hayden Blauzvern and all reviewers who provided feedback on this work.
-B: Software Supply Chain
-Failure Experienced
-B-1: Describe briefly the team’s process from project conception to product release and maintenance – This is to understand the unique context of each practitioner’s software production case.
-B-2: $\star$ What do you consider a Software Supply chain attack/incident to be?
-B-3: Can you describe any specific software supply chain risks (or incidences with third-party dependencies, code contributors, open source, etc.) that your team has encountered during your software development process?
-How were these addressed?
-(Alternatively – How did this affect the decision to implement software signing?
-[…]
-B-5: What is your team’s greatest source of software supply chain security risk?
-Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
-C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from Section B): why did the team choose to implement software signing?
-Is software signing the team’s major strategy to secure its supply chain?
-Any other complimentary security efforts and methods?
-Are these strategies (software signing and complementary security techniques) influenced by regulations and standards?
-What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices?
-$\star$ How do different team members contribute to the implementation of Software Signing?
-What roles and responsibilities are involved?
-(After Practise).
-Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain?
-(Added after Pilot).
-Third-party dependencies C-2: What are the team’s peculiar selection strategies for Third-party dependencies?
-How does the presence of a signature influence this decision?
-How is the authenticity of this signature verified?
-Any other methods/practices to ensure the trustworthiness of the third-party dependencies before integrating them into your projects?
-Possible Follow-up – Can you describe these methods/practices?
-C-3: What influences the team to discontinue the use of a package? How does the Signature on the dependency influence this decision?
-C-4: How does the team manage its third-party dependencies’ security? How do you maintain awareness of potential vulnerabilities or threats related to third-party components?
-First-party source code C-5: Is software signing a requirement for team developers (insider threats) and open-source contributors (if any)?
-how does the team use software signing to protect its source code (what parts of the process is signing required e.g.
-Commit signing)?
-Build Process C-6: (How) Does the team utilize signing in its build process?
-Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
-C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism? (This question is derived from  [92]).
-D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
-D-2: What factors did the team consider before adopting this tooling over others?
-D-3: What was the team’s previous signing practice before the introduction of current tooling?
-D-4: How does your team implement this tooling (which components of tool does the team use)?
-D-5: Have you encountered any challenge(s) using this tool of choice?
-How have you coped with this limitation(s)?
-Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
-D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
--C Codebooks
-Table VII describes our codebook, with code definition, and example quote tagged with that code.
-TABLE VII: Excerpts from the final codebook we used for our analysis.
+Ethical Considerations
+Our study investigates the usability of next-generation software signing tools, focusing on Sigstore as an exemplar.
+We conducted semi-structured interviews with 17 security practitioners across 13 organizations, analyzing their perspectives on usability strengths, weaknesses, and adoption dynamics.
+Stakeholders & Potential Harms.
+Following the guidance of Davis et al. [25], we identify the following stakeholders.
+  • Direct stakeholders: (i) Interview participants, whose responses could unintentionally expose their organization’s practices; (ii) The organizations employing these participants, which may face reputational, regulatory, or security risks if weaknesses are revealed; (iii) The research team, responsible for balancing transparency and confidentiality; and (iv) Adversaries, who could misuse disclosed weaknesses in signing tools or deployment practices.
+  • Indirect stakeholders: (i) Other software-producing organizations, who may be affected if the study highlights industry-wide shortcomings in signing tool used by such organization; (ii) Standards bodies, whose guidelines may be challenged; (iii) Software consumers, whose trust in signed artifacts could be affected; and (iv) The broader public, for whom systemic weaknesses in signing infrastructure may undermine confidence in software supply chains.
+Mitigations Implemented.
+Recruitment Policies:Data were collected through voluntary interviews, with written consent obtained prior to participation.
+Recruitment was conducted through established professional networks; no unsolicited contact or spamming was used.
+Participants were free to decline or withdraw at any stage, and recruitment avoided any form of coercion.
+All organizations were treated equally in data collection and analysis.
+At no point were sensitive keys, cryptographic material, or detailed system configurations collected.
+Content safeguards: All transcripts were anonymized; neither participant names nor organizational identifiers appear in published results.
+Findings are reported in aggregate at the thematic level (e.g., categories of usability issues) rather than specific configurations or vulnerabilities.
+Privacy & data minimization: Only de-identified transcripts were retained; analysis outputs emphasize themes across participants, not individual accounts.
+No identifiable deployment details are included in the publication.
+Operational safety: We deliberately excluded or generalized any descriptions of signing deployments that could aid adversaries, including specific weaknesses in Sigstore’s components.
+Where risks were mentioned (e.g., transparency log privacy concerns, private instance deployment issues), results are presented at the level of usability factors rather than detailed attack vectors.
+Participant Welfare.
+As mentioned earlier, all participants gave informed consent.
+Potential distress was minimal, though participants might experience discomfort when reflecting on organizational challenges.
+To mitigate this, questions emphasized practices and perceptions rather than failures, and responses were anonymized.
+The study did not involve deception.
+Oversight and Consent Context.
+The academic institution’s IRB reviewed and approved the protocol. Prior to consenting, all participants were informed of the study’s aims and the confidentiality safeguards in place.
+Decision to Proceed and Publish.
+We judged that the expected benefits—independent evidence on the usability of next-generation signing tools, insights into adoption barriers, and guidance for improving software supply chain security—outweigh residual risks under the mitigations described above.
+To minimize adversarial risk, we publish only aggregated themes, not sensitive technical details, ensuring that our findings strengthen rather than weaken trust in signing infrastructures.
+Open Science
+We acknowledge that USENIX Security has an open science policy: that authors are expected to openly share their research artifacts by default. The research artifacts associated with this study are:
+  • Raw transcripts of interviews
+  • Anonymized transcripts of interviews
+  • Interview protocol
+  • Codebook
+Things we have shared: The interview protocol is a crucial part of any interview study, since it allows for the critical review of a study design as well as its replication.
+Our full interview protocol is included in Appendix˜B.
+Since it is semi-structured, we include all of the questions asked of all subjects, as well as examples of the follow-up questions we asked.
+We also share the codebook, with codes, definitions, and example quotes that we coded for each code — see Appendix˜C.
+Things we cannot share: For subject privacy reasons, and for IRB compliance, we cannot share the raw transcripts.
+We are also unwilling to share the anonymized transcripts of the interviews.
+Given the high organizational ranks of many of our subjects, and the small size of the subject pool resulting from our recruiting strategy, we believe there is a high risk of de-anonymization even of anonymized transcripts.
+Therefore, we cannot share the anonymized transcripts.
+Artifact Repository: An anonymized artifact containing our interview protocol, codebook, etc. is available at: https://github.com/nextgenusability/next-gen-usabilty
+Appendix A Traditional Signing Tools Workflow
+The signature creation process begins when the maintainer’s artifact is ready for submission.
+The signer generates a key pair that provides them with a private key for signing their software artifact and a public key for a verifier to validate the artifact’s signature.
+Once the artifact is signed, it must be submitted to a package registry (such as PyPi, npm, or Maven) or the intended party.
+During this phase, the signer is responsible for making their public key available and discoverable so future users can verify the signer’s identity.
+In the verification phase, the verifier retrieves the signer’s public key, uses it to decrypt the signature file, and checks for equivalency with the software artifact’s hashed digest.
+We show a typical traditional package signing workflow in Figure˜6 compared to a next-generation workflow in Figure˜7.
+Figure 6: Typical workflow for software signing and verifying signatures.
+The component author packages (A) and signs (B) their software.
+The signed package (C) and public key (D) are published.
+To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
+Figure 7: Typical workflow for next-generation software signing.
+The component author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider.
+The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
+Appendix B Interview Protocol
+Table˜1 gave a summary of the interview protocol.
+Here we describe the full protocol.
+We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style.
+Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
+We include Section B & C of the protocol for completeness, although they were not analyzed in this study.
+Table 6: Interview Protocol. Sections A–C reflect demographics, signing use/motivation context, and signing tool usability. We analyze part C in this study and use the other themes as context.
+Theme Description / Questions
+A:
+Demographics
+We ask basic demographic questions to establish participant experience and context (role, years of experience, team size, and major products).
+A-1: What best describes your role in your team? (Security engineer, Infrastructure, Software engineer, etc.)
+A-2: What is your seniority level? How many years of experience?
+A-3: What is the team size?
+A-4: What are the team’s major software products/artifacts?
+B:
+Signing Use &
+Context
+We ask about the organizational context of signing: why and how signing is used, what motivates adoption, what infrastructure supports it, and how it fits into broader supply chain security practices.
+Focus is on general practices, not tooling specifics.
+B-1: Why did the team choose to implement software signing?
+B-2: Is software signing the team’s major strategy to secure its supply chain? Any complementary efforts?
+B-3: Are these strategies influenced by regulations or standards?
+B-4: What challenges did your team face while integrating signing into supply chain practices?
+B-5: How do team members contribute to the implementation of signing? Roles/responsibilities?
+B-6: Do you believe signing (if implemented correctly) is sufficient to secure a supply chain?
+B-7: How does the presence of a signature influence dependency selection?
+B-8: How is authenticity of a signature verified in your process?
+B-9: What influences your team to discontinue a dependency? Does signature status play a role?
+B-10: How does your team manage third-party dependency security and vulnerabilities?
+B-11: Is signing a requirement for developers and contributors? If so, in which parts of the workflow?
+B-12: How is signing applied in your build process?
+B-13: How is signing applied to artifacts and deployment (binaries, repositories, pipelines)?
+C:
+Signing Tool
+Usability
+We ask about the signing tools themselves: what tools are in use, factors influencing adoption, challenges, and considerations of alternatives. Each question includes a brief explanation.
+C-1: What software signing tool does the team use? (Understanding current tool choices and components in practice.)
+C-2: What factors did the team consider before adopting this tool over others? (Identifying decision-making drivers.)
+C-3: What was the team’s previous signing practice before the current tool? (Capturing historical evolution of practices.)
+C-4: How does the team implement this tool (which components are in use)? (Understanding deployment in workflows.)
+C-5: What challenges or limitations has the team faced with this tool? How did you cope with these? (Surfacing usability pain points and coping strategies.)
+C-6: Has the team considered switching tools? If yes, which alternatives? (Exploring sustainability of current tool choices.)
+Appendix C Codebooks
+Table˜7 describes our codebook, with code definition, and example quote tagged with that code.
+Table 7: Excerpts from the final codebook we used for our analysis.
 Code Definition Sample Quote
@@
 "P5: "We were using Notary, and Notary only verifies that the image was signed using a specific key, but you cannot verify that who is the owner of the key."
--D Code Saturation
+Appendix D Code Saturation
 Figure 8: Saturation curve.
@@
 However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
-We present a saturation analysis of our studies in  Figure 7.
+We present a saturation analysis of our studies in  Figure˜8.
 The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview.
@@
 This is also reflected in the orange trend line (unique number of codes).
--E Survey Attempt
+Appendix E Survey Attempt
 In addition to our interviews, we developed a survey distributed to Sigstore users at the 2024 SigstoreCon conference and through post-conference communications with the Sigstore user community.
@@
 Here, we summarize the answers to an excerpt of the questions responded to in our survey:
-What signing tool did your team use prior to the introduction of Sigstore?: Among 7 participants who utilize Sigstore, 3/7 responded that they use PGP/GPG-based signing and 4/7 did not use any toolings prior to using Sigstore.
+What signing tool did your team use prior to the introduction of Sigstore?
+Among 7 participants who utilize Sigstore, 3/7 responded that they use PGP/GPG-based signing and 4/7 did not use any toolings prior to using Sigstore.
 Which of the following factors have influenced your team’s use of signing?
@@
 Which Sigstore components does your team use? (select all): Among 7 participants who utilize Sigstore, 3 responded to this question, and all 3 participants selected Cosign, Fulcio, and Rekor.
-In which of the following ways have you implemented Sigstore?: Among 7 participants who utilize Sigstore, 3 responded to this question, with 2 responding that they use a private deployment of Sigstore and 1 responding that they use a public instance of Sigstore.
-Do you use or require Sigstore (signing) at any of the following stages of your product?: Among 7 participants who utilize Sigstore, 3 responded to this question.
+In which of the following ways have you implemented Sigstore?
+Among 7 participants who utilize Sigstore, 3 responded to this question, with 2 responding that they use a private deployment of Sigstore and 1 responding that they use a public instance of Sigstore.
+Do you use or require Sigstore (signing) at any of the following stages of your product?
+Among 7 participants who utilize Sigstore, 3 responded to this question.
 All 3 participants shared the "Build/Deployment/final product" stage as part of their answer.
@@
 2/3 selected that they check Rekor logs.
--F Sigstore Components Most Commonly Used by Our Sample Population
-TABLE VIII: Sigstore Functionalities Used by Subjects.
+E.1 Sigstore Components Most Commonly Used by Our Sample Population
+Table 8: Sigstore Functionalities Used by Subjects.
 Tool # of Subjects
@@
 Component Customization & Local Sigstore Deployment 2
--G Expanded Demographics Table
-TABLE IX: Subject Demographics.
+E.2 Other Results
+In this appendix we prresent a continuation of the results presented and shown in Figure˜4 to show the Organizational and usage context report from participants in medium-scale organizations.
+Figure 9: Heatmap showing the organizational usability context for subjects from medium scale organizations (4 subjects).
+Similar to the one showed for small and large-scale organizations in Figure˜4.
+In comparing to Figure˜4, note that some columns are omitted in this figure because no subjects mentioned those aspects.
+E.3 Expanded Demographics Table
+Table 9: Subject Demographics.
 For anonymity, we used generic job roles, not specific titles.
@@
 P17 Manager 11 years Security tooling
--H Factors Influencing Sigstore Adoption (Sigstore Users)
-TABLE X: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
+E.3.1 Organizational Demographics
+We summarize additional organizational demographic information of our subjects in Table˜10.
+This table provides further context for the subject’s organizational size, type of software product, and how many of our subjects belonged to each organization.
+Table 10: Organizational Demographics. To enhance anonymity, We only highlight the number of subjects in each organizational category. letters A-L is used to depict each organization.
+Type Breakdown (#Organizations|#Subjects)
+Organizational Size (Employee Size) Small (<100) (4/6), Medium (<1500) (3/4), Large (>1500) (6/7)
+Product Area Digital technology (1/3), SSC Security (2/4), Social technology (1/1), Dev tools (1/1), Telecommunications (1/1), Cloud security (2/3), Aerospace Security (1/1), Internet services (2/2), Cloud + OSS Security (1/1), Cloud + Dev toools (1/1)
+Subject Distribution A (2), B (3), C (1), D (1), E (2), F (1), G (1), H (1), I (1), J (1), J (1), K (1), L (1), M(1)
+E.4 Factors Influencing Sigstore Adoption (Sigstore Users)
+In this appendix, we present a tabular organization of the reasons given by Sigstore users for their teams’ and organizations’ adoption of Sigstore, structured according to Cresswell et al.
+’s four-factor framework.
+Table 11: Reasons Practitioners Choose or Switch to Sigstore Before Adoption.
 Practitioners’ decisions are driven primarily by human/social factors—such as community contribution, frustration with legacy tools, and usability issues—supplemented by Sigstore’s technological capabilities and reinforced by macroenvironmental pressures like regulations and trust in the CNCF ecosystem.
@@
 Practitioners Contribute to Sigstore 6 – P2, P4, P6, P7, P9, P14
-GPG Issues 6
+GPG Issues 4
 Low adoption rates P1, P14
+Steep learning curve P12, P16
+Notary Issues 1
+Non-demand from customers P12
+Technological Factors
+Available Sigstore Functionalities 3 – P1, P5, P14
+A transparency log, etc P5, P14
+Compatibility to other Tools P1
+GPG Issues 5
 Key management issues & Other usability concerns P1, P6, P12, P16, P17
-Steep learning curve P12, P16
 Compatibility with newer technology P16
 Notary Issues 2
-Non-demand from customers P12
 Compatibility with other tools P12
@@
 Difficult to setup P15
-Technological Factors
-Available Sigstore Functionalities 3 – P1, P5, P14
-A transparency log, etc P5, P14
-Compatibility to other Tools P1
 Macroenvironmental Factors
-Regulation & Standards 4 – P5, P6, P15, P17
-Large User Community 1 – P9
+Regulation & Standards (Cross-factor)- M & O 4 – P5, P6, P15, P17
+Large Sigstore User Community 1 – P9
 Inherent Trust of Creators 1
 Trust of CNCF products P3
--I Factors Influencing Sigstore Consideration (Non-Sigstore Users).
-TABLE XI: Non Sigstore Users: Tooling Issues causing a consideration of Sigstore and other Next-gen tools.
+E.5 Factors Influencing Sigstore Consideration (Non-Sigstore Users).
+Table 12: Non Sigstore Users: Tooling Issues causing a consideration of Sigstore and other Next-gen tools.
 Tool Subjects Topic
@@
 (4) Usage info/documentation unclear T – P11
-(5) Verification not built in with tool O – P8
+(5) Verification not built in with tool T – P8
 (6) Non transparency in signing infrastructure T – P11
@@
 (2) Lack of an inbuilt signing policy T
-(3) Engineers don’t understand the current tool H – P11
+(3) Engineers don’t understand the current tool P – P11
 (4) Integration with other tools/platformsT
@@
 (1) Use of Key discovery and delivery systems like TUF [^3] The Update Framework[17]
--J Additional Discussions & Implications of Results
-In addition to the implications discussed in section VI, we outline further implications of our findings below.
--J1 Generalizability of Findings to Other Classes of Tools in Software Engineering
+E.6 Additional Discussions & Implications of Results
+In addition to the implications discussed in section˜6, we outline further implications of our findings below.
+E.6.1 Generalizability of Findings to Other Classes of Tools in Software Engineering
 Our findings reinforce several adoption factors noted in prior literature on open source tools—such as ease of use, compatibility with existing technologies, and integration flexibility—which were especially salient for next-generation signing tools like Sigstore.
@@
 This model can also be important to understand and further propose a streamlined organizational adoption pathway for software engineering tools.
--J2 The Surprising Role of Community in Software Signing Tool Adoption
-Many participants cited their involvement in the Sigstore community as a key reason for adopting the tool (Table X).
+E.6.2 The Role of Community in Software Signing Tool Adoption
+Many participants cited their involvement in the Sigstore community as a key reason for adopting the tool (Table˜11).
 While this may partly reflect our sampling, half of the non-Sigstore users who adopted external tools also emphasized community affiliation as a motivator.
-Challenges like setting up private Sigstore instances or accessing usage knowledge were frequently linked to the absence of sufficient community support (section V-A2).
+Challenges like setting up private Sigstore instances or accessing usage knowledge were frequently linked to the absence of sufficient community support (section˜5.1.2).
 Other community-related influences, such as the size of the user community and trust in the tool creators, also shaped adoption decisions.
@@
 Toolmakers should actively foster supportive, trustworthy communities around their tools, as these networks can be pivotal to both adoption and long-term sustainability.
--K Legible Images
-We have added a more legible version of our Sigstore workflow imageFigure 3 here Figure 8.
-Figure 8: Sigstore Signing Workflow.
-The software author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider.
-The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
+E.6.3 Push-Pull-Barrier Definition
+In this appendix, we map our generated themes and subthemes to the Push, Pull, and Barrier factors from the TRL framework by Bansal et al.
+[10], as described in the results section.
+The table defines the driver type, corresponding theme, and examples.
+As this analysis was highly objective, a single rater performed this mapping.
+The table is here: Table˜13.
+Table 13: Driver labels (Push/Pull/Barrier) mapped to concrete codes with factor tags (T/P/O/M) and example participant IDs.
+Driver Theme Code Factor Examples
+Barrier Policy/ownership Centralize org code signing internally O P8
+Barrier Policy/ownership Organization’s open-source policy constraints O P8
+Barrier Policy/ownership Organizational use cases don’t fit Sigstore O P8
+Barrier Policy/ownership Tool owned by organization O P10
+Barrier Risk/privacy/trust Low trust in Sigstore maintainers M P13
+Barrier Risk/privacy/trust Sigstore privacy concerns T P13
+Barrier Risk/privacy/trust Third-party security risk concerns T P11, P8
+Pull Community/engagement Practitioners contribute to Sigstore H P2, P4, P6, P7, P9, P14
+Pull Macro pressures Large Sigstore user community M/H P9
+Pull Macro pressures Regulation & standards alignment M/O P5, P6, P15, P17
+Pull Macro pressures Trust in CNCF/creators M P3
+Pull Sigstore capabilities Compatibility with other tools/CI T P1
+Pull Sigstore capabilities Transparency log & related features T P5, P14
+Push General Key & identity management pain T P5
+Push GPG issues (human) Low adoption rates H P1, P14
+Push GPG issues (human) Steep learning curve H P12, P16
+Push GPG issues (tech) Compatibility with newer technology T P16
+Push GPG issues (tech) Key management & other usability pain T P1, P6, P12, P16, P17
+Push Notary v1 issues Engineers don’t understand tool P P11
+Push Notary v1 issues Key management T P10
+Push Notary v1 issues Lack of in-built signing policy T P10
+Push Notary v1 issues Lack of regular updates T P12
+Push Notary v1 issues Compatibility problems T P12
+Push Proprietary/Internal tool issues Difficult to set up T P15
+Push Proprietary/Internal tool issues Engineers don’t understand current tool P P11
+Push Proprietary/Internal tool issues Integration with other tools/platforms T P8, P11
+Push Proprietary/Internal tool issues Non-transparent signing infrastructure T P11
+Push Proprietary/Internal tool issues Non-unified implementations across teams O P8, P11
+Push Proprietary/Internal tool issues Security concerns about signing infrastructure T P11
+Push Proprietary/Internal tool issues Signer ID management problems T P11
+Push Proprietary/Internal tool issues Unclear usage info/docs T P11